### v1

# 111* Corresponding author

Abstract.
A novel approach is introduced for computing the multi-graded Betti numbers of normal edge rings. This method is employed to delve into the edge rings of three distinct classes of simple graphs that adhere to the odd-cycle condition. These classes include compact graphs, which are devoid of even cycles and satisfy the odd-cycle condition; graphs comprised of multiple paths converging at two shared vertices; and graphs introduced in [10] that exhibit both even and odd cycles. Explicit formulas are provided for the multi-graded Betti numbers pertaining to the edge rings of these graphs.
Key words and phrases:
Compact graph, Canonical module, Graded Betti number, Toric algebra
2010 Mathematics Subject Classification:
Primary 13A02, 05E40 Secondary 06D50.
1. Introduction
Let $G=(V(G),E(G))$ be a simple graph with vertex set $V(G)=\{x_{1},\ldots,x_{n}\}$ and edge set $E(G)=\{e_{1},\ldots,e_{m}\}$ and let $\mathbb{K}$ be a field. The edge ring $\mathbb{K}[G]$ is the toric ring $\mathbb{K}[x_{e}\mid e\in E(G)]\subseteq\mathbb{K}[x_{1},\ldots,x_{n}]$. Here, $x_{e}=\prod_{x_{i}\in e}x_{i}$ for all $e\in E(G)$. Let $\mathbb{K}[E(G)]$ be the polynomial ring $\mathbb{K}[e_{1},\ldots,e_{m}]$. Then there is exactly one ring homomorphism
$\phi:\mathbb{K}[E(G)]\rightarrow\mathbb{K}[V(G)]\quad{\mbox{ such that }}~{}e_{i}\mapsto x_{e_{i}}~{}\mbox{~{}for~{}}{i=1,\ldots,m.}$
The kernel of the homomorphism map $\phi$ is called the toric ideal or the defining ideal of $\mathbb{K}[G]$ or $G$, which is denoted by $I_{G}$. It follows that $\mathbb{K}[G]\cong\mathbb{K}[E(G)]/I_{G}$. In general, calculating the Betti numbers or other invariants of a toric ideal for a specific type of graphs is challenging. For recent advancements in this area, please refer to [2], [6], [7], and [8] and the references therein.
In 1999, Ohsugi and Hibi established in [15] that the edge ring $\mathbb{K}[G]$ is a normal monoid algebra precisely when the graph $G$ fulfills the odd-cycle condition. Recall that a simple graph satisfies the odd-cycle condition whenever any two cycles $C_{1}$ and $C_{2}$ either share at least one vertex or are connected by an edge bridging a vertex from $C_{1}$ to a vertex from $C_{2}$. Assuming the normality of $\mathbb{K}[G]$, Theorem 6.31 in [1] reveals that the multi-graded canonical module of the multi-graded $\mathbb{K}[E(G)]$-algebra $\mathbb{K}[G]$, denoted as $\omega_{\mathbb{K}[G]}$, constitutes an ideal of $\mathbb{K}[G]$ with a generating set that possesses a geometric interpretation, thereby simplifying its computation. Leveraging this understanding, we can determine the top multi-graded Betti numbers of $\mathbb{K}[G]$ by invoking Formula 6.6 from [1]. Note that the top multi-graded Betti numbers of a finitely generated $\mathbb{Z}^{m}$-graded module $M$ refer to the numbers $\beta_{p,h}(M)$ for $h\in\mathbb{Z}^{m}$ if $M$ has a projective dimension $p$. In the following text, we will also utilize the terminology second top multi-graded Betti numbers of $M$ to refer to the numbers $\beta_{p-1,h}(M)$ for $h\in\mathbb{Z}^{m}$.
In this paper, motivated by the aforementioned facts, we introduce a novel approach for computing the graded Betti numbers of normal edge rings. This methodology is further founded on the following result.
Lemma 1.1.
Let $H$ be an induced subgraph of a simple graph $G$. Then for $h\in\mathbb{Z}_{\geq 0}^{|V(G)|}$ with $\mathrm{supp}(h)\subseteq V(H)$, we have
$\beta_{i,h}(\mathbb{K}[H])=\beta_{i,h}(\mathbb{K}[G])\qquad\forall i\geq 0.$
Let $G$ be a simple graph that satisfies the odd-cycle condition. Assuming that the total Betti numbers $\beta_{i}(\mathbb{K}[E(G)]/\mathrm{in}_{<}(I_{G}))$ for some monomial order $<$ and the projective dimension $p$ of $\mathbb{K}[G]$ have been computed, our approach proceeds as follows.
  1. (1)
    Initially, we compute the top multi-graded Betti numbers $\beta_{p,h}(\mathbb{K}[G])$ of $\mathbb{K}[G]$ by determining the minimal generators of $\omega_{\mathbb{K}[G]}$.
  2. (2)
    For values of $i$ less than $p$, we identify all induced subgraphs of $G$ (which also adhere to the odd-cycle condition) with edge rings possessing a projective dimension of $i$. Subsequently, we calculate the top Betti-numbers of the edge rings of these induced subgraphs and add them up.
  3. (3)
    If the sum matches $\beta_{i}(\mathbb{K}[E(G)]/\mathrm{in}_{<}(I_{G}))$, we have successfully computed multi-graded Betti numbers $\beta_{i,h}(\mathbb{K}[G])$ for $h\in\mathbb{Z}^{|V(G)|}$.
  4. (4)
    Otherwise, we may need to consider induced subgraphs with edge rings of projective dimension $i+1$ and to compute their second top Betti numbers.
In this paper, we successfully utilize this procedure to compute the multi-graded Betti numbers of edge rings associated with three distinct classes of simple graphs that adhere to the odd-cycle condition.
According to [20], a simple graph is calleded compact if it lacks even cycles and satisfies the odd-cycle condition. We computed the type, regularity and the top graded Betti numbers of their edge rings in [20]. Herein, we present an explicit formula for the multi-graded Betti numbers of the edge ring $\mathbb{K}[G]$ associated with a compact graph $G$. We call a vertex of a compact graph to be a big vertex if it has a degree greater than 2. In view of [20, Proposition 2.3], there are at most three big vertices in a compact graph. In the scenario where $G$ possesses three big vertices, i.e., $G$ is a compact graph of type 3, the fourth step of the aforementioned procedure becomes necessary. The multi-graded Betti numbers of a compact graph are computed in the following theorem.
Theorem 1.2.
(Theorem 4.11) Let $G$ be a compact graph, and denote by $p$ the projective dimension of $\mathbb{K}[G]$. Consider an integer $i$ satisfying $1\leq i\leq p$. Let $H_{j}$ for $j=1,\ldots,s$ represent all the compact induced subgraphs of $G$ having exactly $i+1$ minimal cycles, and let $L_{j}$ for $j=1,\ldots,t$ be all the compact induced subgraphs of $G$ belonging to type 3 with $i+2$ minimal cycles. Define the following sets:
$\mathcal{N}_{i}(G)=\bigcup_{j=1}^{s}\mathcal{N}_{H_{j}},\quad\mathcal{M}_{i}^{1}(G)=\bigcup_{j=1}^{t}\mathcal{M}_{L_{j}}^{1},\quad\text{and}\quad\mathcal{M}_{i}^{2}(G)=\bigcup_{j=1}^{t}M_{L_{j}}^{2}.$
Then, the $i$-th multi-graded Betti numbers of $\mathbb{K}[G]$ can be expressed as follows:
$\beta_{i,\alpha}(\mathbb{K}[G])=\begin{cases}2,&\text{if }\alpha\in\mathcal{M}_{i}^{2}(G),\\ 1,&\text{if }\alpha\in\mathcal{M}_{i}^{1}(G)\cup\mathcal{N}_{i}(G),\\ 0,&\text{otherwise}.\end{cases}$
Here, $\mathcal{N}_{H},\mathcal{M}_{H}^{1}$ and $\mathcal{M}_{H}^{2}$ are sets of monomials that could be read from the graph $H$ directly if the compact graph $H$ is labeled.
In our pursuit of obtaining this result, we positively answer the following question posed in [20]:
Does $\mathbb{K}[G]$ and $\mathbb{K}[E(G)]/\mathrm{in}_{<}(I_{G})$ always share the same graded Betti numbers? Here, $<$ is a monomial order discussed in [20].
Furthermore, we investigate two additional classes of simple graphs that designated as $\mathbf{G}_{\underline{\ell}}$ and $\mathbf{G}_{m}$, respectively. Let ¯$\underline{\ell}=(\ell_{1},\ldots,\ell_{m})$ be a given integral vector, where $\ell_{i}\geq 1$ for all $i=1,\ldots,m$, with the condition that there is at most one index $i$ where $\ell_{i}=1$. By definition, the graph $\mathbf{G}_{\underline{\ell}}$ comprises $m$ paths, each having lengths $\ell_{1},\ldots,\ell_{m}$ respectively. These paths all originate from a common vertex and converge at another shared vertex. We call such graphs multi-path graphs. This class of graphs includes many graphs that have been investigated in literatures. Specifically, in the paper [7] (resp. in [14]), a particular class of graphs designated as $G_{r,d}$ (and $G_{r,s,d}$ respectively) was introduced. It was then demonstrated that there is an initial ideal of the toric ideal of $G_{r,d}$ (or $G_{r,s,d}$) which possesses the same graded Betti number as the toric ideal itself. Furthermore, the formula for computing their graded Betti numbers were provided. Notably, the graph $G_{r,d}$ (and $G_{r,s,d}$ accordingly) is represented as $\mathbf{G}_{\underline{\ell}}$, where $\underline{\ell}$ is defined as $(2r-2,\underbrace{2,\ldots,2}_{d})$ for $G_{r,d}$ and $(2r-2,2s-2,\underbrace{2,\ldots,2}_{d})$ for $G_{r,s,d}$. In another paper [11], the graphs $B_{\underline{\ell},h}$ was introduced and the graded Betti number of its toric ideal were computed in specific scenarios. It is noteworthy that the graph $B_{\underline{\ell},h}$ is also represented as $\mathbf{G}_{\underline{\ell}}$, where each component $\ell_{i}$ of $\underline{\ell}$ is an even number.
The class of graphs $\mathbf{G}_{m}$ for $m\geq 1$ was introduced in [10], where an intriguing property of the toric ideals associated with their edge rings was revealed. We refer to graphs $\mathbf{G}_{m}$ as two-ear graphs based on their shape, as illustrated in Figure 2. It was shown in [10] that the initial ideal of $I_{\mathbf{G}_{m}}$ has a depth of 7 with respect to the $<_{rev}$ order, whereas it exhibits the Cohen-Macaulay property when considered under the $<_{lex}$ order. In this paper, we will present theorems like Theorem 4.11 to compute multi-graded Betti numbers for the edge rings pertaining to multi-path graphs and two-ear graphs.
Theorem 1.3.
(Theorem 5.13 and Theorem 6.10) Let $G$ be either a two-ear graph $\mathbf{G}_{m}$ or a multi-path graph $\mathbf{G}_{\underline{\ell}}$ with each $\ell_{i}>1$. Let $\mathcal{N}_{i}(G)$ denote the union of all $\mathcal{N}_{H}$ such that $H$ is an induced subgraph of $G$ satisfying $\mathrm{pdim}(\mathbb{K}[H])=i$. Then for any monomial $\alpha$ in $\mathbb{K}[G]$ and for any integer $i\geq 1$, we have
$\beta_{i,\alpha}(\mathbb{K}[G])=\left\{\begin{array}[]{ll}1,&\hbox{$\alpha\in\mathcal{N}_{i}(G)$;}\\ 0,&\hbox{otherwise.}\end{array}\right.$
These results and their corollaries recover many results mentioned before. For example, the result that $\mathrm{in}_{<_{lex}}(I_{\mathbf{G}_{m}})$ is Cohen-Macaulay is a direct consequence of our result that $\mathrm{in}_{<_{lex}}(I_{\mathbf{G}_{m}})$ and $I_{\mathbf{G}_{m}}$ share the same Betti numbers, see Proposition 7.3.
The structure of this paper is outlined as follows. Initially, in the second section, we introduce some fundamental concepts and outcomes that will serve as the foundation for our subsequent discussions. Subsequently, in Sections 3 and 4, we delve into the intricate study of the edge rings of compact graphs, culminating in the presentation of Theorem 4.11. Then, in Sections 5 and 6, we turn our attention to the examination of multi-path graphs and two-ear graphs. Within these sections, we present theorems analogous to Theorem 4.11, tailored to address these two distinct categories of graphs. Finally, in the concluding section, we offer several closing remarks to summarize our findings.
2. Preliminaries
In this section, we provide a brief review of the notation and fundamental facts that will be utilized later on.
2.1. Betti numbers and Monomial ideals with regular quotients
Let $\mathbb{K}$ be a field, and let $R:=\mathbb{K}[X_{1},\ldots,X_{n}]$ be the polynomial ring in variables $X_{1},\ldots,X_{n}$ that are $\mathbb{Z}^{m}$-graded. Here, we do not require that $m=n$. For a finitely generated $\mathbb{Z}^{m}$-graded $R-$module $M$, there exists the minimal multi-graded free resolution of $M$ that has the form:
$0\rightarrow\underset{h\in\mathbb{Z}^{m}}{\bigoplus}R[-h]^{\beta_{p,h}(M)}\rightarrow\cdots\rightarrow\underset{h\in\mathbb{Z}^{m}}{\bigoplus}R[-h]^{\beta_{1,h}(M)}\rightarrow\underset{h\in\mathbb{Z}^{m}}{\bigoplus}R[-h]^{\beta_{0,h}(M)}\rightarrow M\rightarrow 0$
Here, $R[-h]$ is the cyclic free $R$-module generated in degree $h$. The projective dimension of $M$ are defined to be $\mathrm{pdim}\,(M):={\mbox{max}}\,\{i\mid\beta_{i,\,h}(M)\neq 0\mbox{ for some}h\}.$ The number $\beta_{i,h}(M):={\rm{dim}}_{\mathbb{K}}\mathrm{Tor}_{i}^{R}(M,\mathbb{K})_{h}$ is called the $(i,h)$-th multi-graded Betti number of $M$ and $\beta_{i}(M):=\sum_{h\in\mathbb{Z}^{m}}\beta_{i,h}(M)$ is called the $i$-th total Betti number of $M$.
Denote by $p$ the projective dimension of the module $M$. Subsequently, $\beta_{p}(M)$ is referred to as the top total Betti number of $M$, while $\beta_{p,h}(M)$ for $h\in\mathbb{Z}^{m}$ are designated as the top multi-graded Betti numbers of $M$. On the other hand, $\beta_{p-1}(M)$ is labeled as the second top total Betti number of $M$, and $\beta_{p-1,h}(M)$ for $h\in\mathbb{Z}^{m}$ are designated as the second top multi-graded Betti numbers of $M$.
To derive upper bounds for Betti numbers of toric ideals under our investigation, we need a notion that was originally introduced in [12] and subsequently explored in [11]. Recall that a monomial ideal $I$ in $R$ is said to possess regular quotients if there exists a linear ordering $\alpha_{1}\prec\alpha_{2}\prec\ldots\prec\alpha_{r}$ among the minimal generators $\{\alpha_{1},\alpha_{2},\ldots,\alpha_{r}\}$ of $I$ such that, for each $s=2,\ldots,r$, the quotient ideal $(\alpha_{1},\ldots,\alpha_{s-1}):\alpha_{s}$ is generated by a regular sequence. Notably, a sequence of monomials $\beta_{1},\ldots,\beta_{r}$ is regular precisely when the supports of distinct monomials are disjoint, i.e., $\mathrm{supp}(\beta_{i})\cap\mathrm{supp}(\beta_{j})=\emptyset$ for all $1\leq s<t\leq r$. Here, the support of a monomial $\alpha$, denoted by $\mathrm{supp}(\alpha)$, refers to the set of variables that divide $\alpha$.
Assuming that the quotient ideal $(\alpha_{1},\ldots,\alpha_{s-1}):\alpha_{s}$, denoted as $L_{s}$, is generated by a regular sequence of length $k_{s}$ for each $s$ ranging from $2$ to $r$, we can apply Koszul theory to deduce that the total Betti numbers $\beta_{i}(R/L_{s})$ are given by $\beta_{i}(R/L_{s})=\binom{k_{s}}{i}$ for all non-negative integers $i$. Let $I_{s}$ represent the monomial ideal generated by $\alpha_{1},\ldots,\alpha_{s}$, and let $d_{s}$ represent the degree of $\alpha_{s}$ for each $s$ from $1$ to $r$.
Now, consider the following short exact sequences:
$0\rightarrow R/L_{s}[-d_{s}]\rightarrow R/I_{s-1}\rightarrow R/I_{s}\rightarrow 0,\quad\text{for all}\ s=2,\ldots,r.$
Utilizing the mapping cone construction (see e.g. [17, section 27] for the details), we can deduce that
$\beta_{i}(R/I_{s})\leq\beta_{i-1}(R/L_{s})+\beta_{i}(R/I_{s-1}).$
Repeatedly applying the above inequality, we arrive at
$\beta_{i}(R/I)\leq\beta_{i-1}(R/L_{r})+\cdots+\beta_{i-1}(R/L_{2})+\beta_{i}(R/(\alpha_{1}))$
and
$\beta_{i}(I)\leq\beta_{i}(R/L_{r})+\cdots+\beta_{i}(R/L_{2})+\beta_{i}(R).$
This implies that, for all non-negative integers $i$, we have the following inequality, which will be crucial in our subsequent discussion:
(1) $\beta_{i}(I)\leq\sum_{j=2}^{r}\binom{k_{j}}{i}+\binom{0}{i}.$
Apart from this formula, the following result, as mentioned, for instance, in [9, Corollary 3.3.3], is necessary for us to derive the upper bound for Betti numbers of edge rings.
Lemma 2.1.
Let $I$ be a homogeneous ideal of the polynomial ring $R=\mathbb{K}[X_{1},\ldots,X_{n}]$, and let $<$ be any monomial order on $R$. Then for all $i,j\geq 0$, we have
$\beta_{i,j}(R/I)\leq\beta_{i,j}(R/\mathrm{in}_{<}(I)).$
In particularly,
$\beta_{i}(R/I)\leq\beta_{i}(R/\mathrm{in}_{<}(I))\mbox{ and }\mathrm{pdim}(R/I)\leq\mathrm{pdim}(R/\mathrm{in}_{<}(I)).$
2.2. Betti numbers of monomial ideals in a monomial algebras
Let $T$ represent the subalgebra of the polynomial ring $\mathbb{K}[Y_{1},\ldots,Y_{m}]$ generated by monomials $y_{1},\ldots,y_{n}$. In [18], $T$ is called a monomial algebra. Within $T$, all the monomials form a monoid under multiplication, denoted by $\mathcal{U}$, and $T$ is isomorphic to the positive affine monoid ring $\mathbb{K}[\mathcal{U}]$. We define $R$ as the polynomial ring $\mathbb{K}[X_{1},\ldots,X_{n}]$ and endow it with a multi-grading in $\mathbb{Z}^{m}$ by assigning the multi-degree of $y_{i}$ to $X_{i}$. Consider an ideal $I$ of $T$ generated by certain monomials from $\mathcal{U}$.
By the existence of a surjective homomorphism from $R$ to $T$ mapping $X_{i}$ to $y_{i}$, it follows that $T$, $I$ and $T/I$ are all finitely generated $\mathbb{Z}^{m}$-graded $R$-modules.
A formula for computing of the multi-graded Betti numbers of $T/I$ was previously established in [3] using relative simplicial homology. However, in this paper, we require to calculate the multi-graded Betti numbers of $I$. For this, we associate every degree $h\in\mathbb{Z}_{\geq 0}^{m}$ with the following simplicial complex:
$\Gamma_{h}(I)=\{F\subseteq[n]\mid Y^{h}=y^{F}\alpha\mbox{ for some monomial }\alpha\in I\}.$
Here, $Y^{h}$ is defined as $Y_{1}^{h_{1}}\cdots Y_{m}^{h_{m}}$ for $h=(h_{1},\ldots,h_{m})\in\mathbb{Z}_{\geq 0}^{m}$, and $y^{F}$ represents the product of all monomials $y_{i}$ with $i\in F$.
Lemma 2.2.
Let $I$ be a monomial ideal of $T$ and $h$ a degree $h\in\mathbb{Z}_{\geq 0}^{m}$. Then
$\beta_{i,h}^{R}(I)=\dim_{\mathbb{K}}\widetilde{H}_{i-1}(\Gamma_{h}(I),\mathbb{K}).$
Proof.
Let $K_{\bullet}$ denote the Koszul complex of $X_{1},\ldots,X_{n}$. Namely, $K_{\bullet}$ is the following complex
$0\rightarrow K_{n}\rightarrow\cdots\rightarrow K_{i}\rightarrow\cdots\rightarrow K_{0}\rightarrow 0$
Here, $K_{i}$ is a multi-graded free $R$-module that possesses a base composed of elements $e_{F}$ such that $|F|=i$ and $F\subseteq[n]$. Moreover,
$\deg(e_{F})=\sum_{i\in F}\deg(X_{i})=\sum_{i\in F}\deg(y_{i})\in\mathbb{Z}_{\geq 0}^{m}.$
In particular, $K_{0}=Re_{\emptyset}=R$. Since $K_{\bullet}$ is a multi-graded minimal free resolution of $\mathbb{K}$, we have
$\beta_{i,h}^{R}(I)=\dim_{\mathbb{K}}\mathrm{Tor}_{i}^{R}(\mathbb{K},I)_{h}=\dim_{\mathbb{K}}H_{i}(K_{\bullet}\otimes_{R}I)_{h}.$
On the other hand,
$(K_{i}\otimes I)_{h}=\bigoplus_{|F|=i}(Re_{F}\otimes_{R}I)_{h}=\bigoplus_{|F|=i}(I(\deg(e_{F}))_{h}.$
Note that
$(I(\deg(e_{F}))_{h}=I_{h-\deg(e_{F})}=\left\{\begin{array}[]{ll}\mathbb{K},&\hbox{$F\in\Gamma_{h}(I)$;}\\ 0,&\hbox{otherwise,}\end{array}\right.$
it follows that there is an isomorphism between the complex $K_{\bullet}\otimes_{R}I$ and the complex $\widetilde{C}_{\bullet}(\Gamma_{h}(I),\mathbb{K})[-1]$, yielding the desired formula.
We may look up this lemma as a generalization of [13, Theorem 1.34], which address the same question on monomial ideals in a polynomial ring.
2.3. Simple graphs and minimal cycles
Let $G$ be a simple graph, i.e., a finite graph without loops and multiple edges, with vertex set $V(G)$ and edge set $E(G)$. Recall that a walk of $G$ of length $q$ is a subgraph $W$ of $G$ such that $E(W)=\{\{i_{0},i_{1}\},\{i_{1},i_{2}\},\ldots,\{i_{q-1},i_{q}\}\}$, where $i_{0},i_{1},\ldots,i_{q}$ are vertices of $G$. Graph $G$ is connected if, for any two vertices $i_{j}$ and $i_{k}$ of $G$, there is a walk between $i_{j}$ and $i_{k}$.
A cycle of $G$ of length $q$ is a subgraph $C$ of $G$ such that
$E(C)=\{\{i_{1},i_{2}\},\{i_{2},i_{3}\},\ldots,\{i_{q-1},i_{q}\},\{i_{q},i_{1}\}\}$
where $i_{1},i_{2},\ldots,i_{q}$ are vertices of $G$ and where $i_{j}\neq i_{k}$ if $j\neq k$. A cycle is called even (resp. odd) if $q$ is even (resp. odd). For a subset $W$ of $V(G)$, the induced subgraph $G_{W}$ is the graph with vertex set $W$ and for every pair $j,k\in W$, they are adjacent in $G_{W}$ if and only if they are adjacent in $G$. For a vertex $j\in V(G)$, let $G\setminus j$ be the induced subgraph $G_{V(G)\setminus\{j\}}$. A cycle of $G$ is called an induced cycle if it is an induced subgraph of $G$. We also call an induced cycle of $G$ to be a minimal cycle of $G$, and use $\mathfrak{t}(G)$ to denote the number of the minimal cycles of $G$.
2.4. Edge cones and canonical modules
For any $e=\{u,v\}\in E(G)$ denote $v_{e}=\mathbf{u}+\mathbf{v}$. The edge cone of $G$, denoted by ${\mathbb{R}}_{+}(G)$, is defined to be the cone of ${\mathbb{R}}^{|V(G)|}$ generated by $\{v_{e}\mid e\in E(G)\}$. In other words,
${\mathbb{R}}_{+}(G)=\{\sum\limits_{e\in E(G)}{a_{e}v_{e}}\mid a_{e}\in{\mathbb{R}}_{+}\mbox{ for all }e\in E(G)\}.$
If $G$ satisfies the odd-cycle condition, then the edge ring $\mathbb{K}[G]$ is normal, see [15]. In [1, Chapter 6], the canonical module of multi-graded version were introduced. By [1, Theorem 6.31], we have the ideal of $\mathbb{K}[G]$ generated all the monomials $x^{h}$ with $h\in\mbox{relint}({\mathbb{R}}_{+}(G))\cap{\mathbb{Z}}^{|V(G)|}$ is the canonical module of the $\mathbb{Z}^{|V(G)|}$-graded local algebra $\mathbb{K}[G]$, which we denote by $\omega_{\mathbb{K}[G]}$. Let $D_{G}$ denote the multi-degree $\sum_{e\in E(G)}v_{e}$. Then we have the following result in view of [1, Formula 6.6].
Lemma 2.3.
Let $G$ be a simple graph satisfying the odd-cycle condition and denote by $p$ the projective dimension of $\mathbb{K}[G]$. Then
$\beta_{p-i,h}(\mathbb{K}[G])=\beta_{i,D_{G}-h}(\omega_{\mathbb{K}[G]})\qquad\mbox{ for all }\quad 0\leq i\leq p\mbox{ and for all }h\in\mathbb{Z}_{\geq 0}^{|V(G)|}.$
Let us express the edge cone ${\mathbb{R}}_{+}(G)$ in terms of linear inequalities, or equivalently, through its facets. To achieve this, additional notation is required. A non-empty subset $T\subseteq V(G)$ is called an independent set of $G$ if $\{j,k\}\not\in E(G)$ for any $j,k\in T$. For an independent set $T\subseteq V(G)$, we use $N_{G}(T)$ (or simply $N(T)$ for brevity) to denote the neighborhood set of $T$, namely,
$N_{G}(T):=\{j\in V(G)\mid\{j,k\}\in E(G)\mbox{ for some }k\in T\}.$
Given a vertex $i$ of $G$, let $H_{i},H_{i}^{+}$ and $H_{i}^{>}$ denote the following equalities or inequalities respectively:
$H_{i}:x_{i}=0,\qquad H_{i}^{+}:x_{i}\geq 0,\qquad H_{i}^{>}:x_{i}>0.$
Given an independent set $T$ of $G$, let $H_{T},H_{T}^{-}$ and $H_{T}^{<}$ denote the following equalities or inequalities respectively:
$H_{T}:\sum_{i\in T}x_{i}=\sum_{i\in N(T)}x_{i},\qquad H_{T}^{-}:\sum_{i\in T}x_{i}\leq\sum_{i\in N(T)}x_{i},\qquad H_{T}^{<}:\sum_{i\in T}x_{i}<\sum_{i\in N(T)}x_{i}.$
Note that $H_{i}\neq H_{\{i\}}$, since the latter represents the equality $x_{i}=\sum_{j\in N(i)}x_{j}$. If $G$ is bipartite, the facets of ${\mathbb{R}}_{+}(G)$ are given in [19, Proposition 3.6 ], that we record in the following result for convenience.
Lemma 2.4.
If $G$ is a bipartite connected graph with bipartition $(V_{1},V_{2})$, then $F$ is a facet of ${\mathbb{R}}_{+}(G)$ if and only if either
  1. $(a)$
    $F=H_{i}\cap{\mathbb{R}}_{+}(G)$ for some $i\in V(G)$ with $G\setminus i$ connected ;
  2. $(b)$
    $F=H_{T}\cap{\mathbb{R}}_{+}(G)$ for some $T\varsubsetneq V_{1}$, such that $G_{T\cup N(T)}$ and $G_{V(G)\setminus(T\cup N(T))}$ connected.
Let $\mathcal{I}$ denote the set of all vertices satisfying condition (a) and $\mathcal{T}$ denote the set of all independent sets satisfying (b). Then, since $\mathrm{aff}(\mathbb{R}_{+}(G))=H_{V_{1}}=H_{V_{2}}$, we have
($\Delta_{1}$) $\mathbb{R}_{+}(G)=H_{V_{1}}\cap\bigcap_{i\in\mathcal{I}}H_{i}^{+}\cap\bigcap_{T\in\mathcal{T}}H_{T}^{-}$
and
$\mathrm{relint}(\mathbb{R}_{+}(G))=H_{V_{1}}\cap\bigcap_{i\in\mathcal{I}}H_{i}^{>}\cap\bigcap_{T\in\mathcal{T}}H_{T}^{<}.$
For the case when $G$ is non-bipartite, we need to introduce the follwoing notions on graphs.
  • We call a vertex $j$ of $G$ regular if each connected component of $G\setminus j$ contains an odd cycle.
  • We say that an independent set $T$ of $V(G)$ is a fundamental set if
    • the bipartite graph on the vertex set $T\cup N(T)$ with the edge set $E(G)\cap\{\{j,k\}\mid j\in T,k\in N(T)\}$ is connected, and
    • either $T\cup N(T)=V(G)$ or each of the connected components of the graph $G_{V(G)\setminus(T\cup N(T))}$ contains an odd cycle.
It follows from [18, Theorem 3.2] or ([15, Theorem 1.7 (a)]) that
Lemma 2.5.
Let $G$ be a non-bipartite connected simple graph on $V(G)=\{1,\ldots,n\}$. Then ${\mathbb{R}}_{+}(G)$ consists of the elements $(x_{1},\ldots,x_{n})\in{\mathbb{R}}^{n}$ satisfying all the following inequalities:
($\Delta_{2}$) $\begin{split}x_{i}&\geq 0\;\;\text{ for any regular vertex }i;\\ \sum\limits_{i\in N(T)}x_{i}&\geq\sum\limits_{i\in T}x_{i}\;\;\text{ for any fundamental set }T.\end{split}$
2.5. Conventions and notions
We conclude this section by fixing notions that plays an important role in the following.
Conventions 2.6.
Let $G$ be a simple graph, with its vertices labelled by letters or subscripted letters, such as $u,u_{i,j},v,v_{i,j},w,w_{i,j}$, and so forth. We establish a correspondence between these labels and unit vectors in ${\mathbb{R}}^{|V(G)|}$ as follows:
$u,u_{i,j},v,v_{i,j},w,w_{i,j},x,y$
correspond to
${\mathbf{u}},{\mathbf{u}}_{i,j},{\mathbf{v}},{\mathbf{v}}_{i,j},{\mathbf{w}},{\mathbf{w}}_{i,j},{\mathbf{x}},{\mathbf{y}}$
respectively. With this notation, we can express ${\mathbb{R}}^{|V(G)|}$ as the set
$\left\{\sum_{i,j}a_{i,j}{\mathbf{u}}_{i,j}+a{\mathbf{u}}+\sum_{i,j}b_{i,j}{\mathbf{v}}_{i,j}+b{\mathbf{v}}+\sum_{i,j}c_{i,j}{\mathbf{w}}_{i,j}+c{\mathbf{w}}\mid a_{i,j},a,b_{i,j},b,c_{i,j},c\in{\mathbb{R}}\right\}.$
By way of example, $\deg(uv)={\mathbf{u}}_{i}+{\mathbf{v}}_{j}$, and $\deg(u^{2}wu_{1,2})=2{\mathbf{u}}+{\mathbf{w}}+{\mathbf{u}}_{1,2}$, and so on.
We identify a monomial with its multi-degree. For example, we will write $\beta_{i,uv}(\mathbb{K}[G])$ for $\beta_{i,{\mathbf{u}}+{\mathbf{v}}}(\mathbb{K}[G])$. Under this identification, we look up $\mathbb{Z}^{|V(H)|}$ as a subset of $\mathbb{Z}^{|V(G)|}$ if $H$ is an induced subgraph of $G$.
Adopting these conventions, for a simple graph $G$ with vertex set $V(G)$, both the edge ring $\mathbb{K}[G]$ and its toric ideal $I_{G}$ possess an $\mathbb{Z}_{\geq 0}^{|V(G)|}$-grading. Consequently, it is permissible to consider the $\mathbb{Z}_{\geq 0}^{|V(G)|}$-graded Betti numbers of $\mathbb{K}[G]$ and $I_{G}$. Additionally, by assigning a degree of 1 to each edge $e$ in $E(G)$, both $\mathbb{K}[G]$ and $I_{G}$ acquire a $\mathbb{Z}$-grading. The relationship between the $\mathbb{Z}_{\geq 0}^{|V(G)|}$-graded Betti numbers and the $\mathbb{Z}_{\geq 0}$-graded Betti numbers is as follows:
$\beta_{i,j}(\mathbb{K}[G])=\sum_{\begin{subarray}{c}h\in\mathbb{Z}_{\geq 0}^{|V(G)|}\end{subarray}}^{|h|=2j}\beta_{i,h}(\mathbb{K}[G]).$
The following key lemma is well-known. We provide a proof for the completeness.
Lemma 2.7.
Let $H$ be an induced subgraph of a simple graph $G$. Then for $h\in\mathbb{Z}_{\geq 0}^{|V(G)|}$ with $\mathrm{supp}(h)\subseteq V(H)$, we have
$\beta_{i,h}(\mathbb{K}[H])=\beta_{i,h}(\mathbb{K}[G])\qquad\mbox{ for all }i\geq 0.$
Proof.
Since $\mathrm{supp}(h)\subseteq V(H)$ we have $\Gamma_{h}(\mathbb{K}[H])=\Gamma_{h}(\mathbb{K}[G])$ by definition. From this fact, the result follows directly in view of Lemma 2.2.
Notation 2.8.
Let $H$ be a simple graph satisfying the odd-cycle condition. We have defined $D_{H}$ to be the multi-degree $\sum_{e\in E(H)}v_{e}$. Since we identify a monomial with its multi-degree, $D_{H}$ can alternatively be interpreted as the product of all edges of $H$. Here, an edge is considered as the product of its two vertices. Furthermore, we employ $\Theta_{H}$ to represent the product of all the vertices that have a degree of two.
It is straightforward to observe that $\Theta_{H}^{2}$ divides $D_{H}$.
Notation 2.9.
Let $H$ be a simple graph satisfying the odd-cycle condition, and let $p$ denote the projective dimension of $\mathbb{K}[H]$. Denote by $\mathrm{Mon}(\mathbb{K}[H])$ the set of all monomials in $\mathbb{K}[H]$. Then $\mathcal{N}_{H}$ is defined to be the set:
$\mathcal{N}_{H}=\{\alpha\in\mathrm{Mon}(\mathbb{K}[H])\mid\beta_{p,\alpha}(\mathbb{K}[H])\neq 0\}.$
Note that $\beta_{p,\alpha}(\mathbb{K}[H])\neq 0$ if and only if $\beta_{p,\alpha}(\mathbb{K}[H])=1$. According to Lemma 2.3, we have
(2) $\mathcal{N}_{H}=\{D_{H}/\alpha\mid\alpha\mbox{ is a minimal generator of }\omega_{\mathbb{K}[H]}\}).$
3. Basics for compact graphs
This section aims to identify the members of $\mathcal{N}_{H}$ for each compact graph $H$. Basic properties of $\mathcal{N}_{H}$ are also presented. To achieve this, we revisit the precise definition and classification of compact graphs as presented in [20].
Definition 3.1.
Let $G$ be a connected simple graph where every vertex has a degree greater than one. We call $G$ to be a compact graph if it does not contain any even cycles and satisfies the odd-cycle condition.
We call a vertex of a compact graph to be a big vertex if it has a degree greater than 2. According to [20, Proposition 2.3], there are at most three big vertices in a compact graph. Therefore, compact graphs containing more than one cycle can be categorized into three classes, each determined by the number of big vertices. More specifically, we say a compact graph falls into type $i$ if it possesses $i$ big vertices for $i=1,2,3.$
A compact graph of type 1 is a finite collection of odd cycles that share a vertex.
A compact graph of type 2 consists of two disjoint compact graphs of type 1, where the two big vertices are connected either by an edge or by an edge as well as a path of even length.
A compact graph of type 3 consists of three disjoint compact graphs of type 1, where every pair of big vertices is connected by an edge.
Suppose $\underline{p}=(p_{1},\ldots,p_{m})$, $\underline{q}=(q_{1},\ldots,q_{n})$ and $\underline{r}=(r_{1},\ldots,r_{k})$ are positive integral vectors with dimensions $m,n$ and $k$ respectively. We denote a compact graph of type 1, where the odd cycles have lengths $12p_{1}+1,\ldots,2p_{m}+$ 1 respectively, as $\mathbf{A}_{\underline{p}}$ or $\mathbf{A}_{(p_{1},\ldots,p_{m})}$.
By $\mathbf{B}_{\underline{p}:\underline{q}}^{0}$ we mean a compact graph of type 2 where the two disjoint compact graphs of type 1 that compose it are $\mathbf{A}_{\underline{p}}$ and $\mathbf{A}_{\underline{q}}$ and where the two big vertices are connected by an edge. Furthermore, if $s>0$ is an even number, then $\mathbf{B}_{\underline{p};\underline{q}}^{s}$ represents the graph obtained by appending a path of length $s$ connecting two big vertices to $\mathbf{B}_{\underline{p}:\underline{q}}^{0}$.
A compact graph of type 3 is denoted by $\mathbf{C}_{\underline{p}:\underline{q}:\underline{r}}$ if the three disjoint compact graphs of type 1 that make up it are $\mathbf{A}_{\underline{p}}$, $\mathbf{A}_{\underline{q}}$ and $\mathbf{A}_{\underline{r}}$ respectively.
Let $H$ be a compact graph. Moving forward to this section, we shall revisit the findings from [20] pertaining to the minimal generators of the edge cone of $H$, and subsequently establish the composition of $\mathcal{N}_{H}$.
3.1. Type one
Given positive integers $m\geq 2$ and $p_{1},\ldots,p_{m}$, we use $A$ to denote the graph $\mathbf{A}_{(p_{1},\ldots,p_{m})}$ for short. Thus $A$ has vertex set
$V(A)=\{u\}\cup\{u_{i,j}\mid 1\leq i\leq m,1\leq j\leq 2p_{i}\}$
and edge set
[…]
$\cup\{\{u,u_{i,1}\},\{u,u_{i,2p_{i}}\}\mid 1\leq i\leq m\}.$
Recall that an integral vector in $\mathrm{relint}({\mathbb{R}}_{+}(G))$ is called minimal if it cannot written as the sum of a vector in $\mathrm{relint}({\mathbb{R}}_{+}(G))\cap\mathbb{Z}^{|V(G)|}$ and a nonzero vector of ${\mathbb{R}}_{+}(G)\cap\mathbb{Z}^{|V(G)|}$. We showed in [20] that the following vectors
$\alpha_{\ell}:=\sum\limits_{i=1}^{m}\sum\limits_{j=1}^{2p_{i}}{\mathbf{u}}_{i,j}+2\ell{\mathbf{u}},\ \ell=1,\ldots,m-1$
are all the minimal vectors of $\mathrm{relint}({\mathbb{R}}_{+}(A))\cap{\mathbb{Z}}^{|V(A)|}$.
It is clear that $D_{H}=\Theta_{H}^{2}u^{2m}$. Therefore, the subsequent result follows directly from Formula 2 in Subsection 2.5.
Lemma 3.2.
For each compact graph $H$ of type 1,
$\mathcal{N}_{H}=\{\Theta_{H}u^{2\ell}\mid\ell=1,\ldots,m-1\}.$
In the expression of $\mathcal{N}_{H}$, the symbol $u$ represents the sole big vertex of the graph $H$. Notably, when considering $H$ as an induced subgraph of a compact graph belonging to either type 2 or type 3, it is possible that the unique big vertex of $H$ could alternatively be designated as $v$ or $w$. In such cases, it is necessary to substitute $u$ with either $v$ or $w$ accordingly.
3.2. type two
Assume that $n,m$ and $p_{1},\ldots,p_{m},q_{1},\ldots,q_{n}$ are given positive integers. Let $s\geq 0$ be an even number. We use $B$ denote the graph $\mathbf{B}^{s}_{(p_{1},\ldots,p_{m}):(q_{1},\ldots,q_{n})}$ for short. Then, we may assume that $B$ has vertex set
[…]
$\cup\{u_{i,j}\mid 1\leq i\leq m,1\leq j\leq 2p_{i}\}$$\cup\{v_{i,j}\mid 1\leq i\leq n,1\leq j\leq 2q_{i}\}$
and edge set
[…]
$\cup\{\{u,u_{i,1}\},\{u,u_{i,2p_{i}}\}\mid 1\leq i\leq m\}$
$\cup\{\{u,v\},\{u,w_{1}\},\{v,w_{s-1}\}\}\cup\{\{w_{i},w_{i+1}\}\mid 1\leq i\leq s-2\}$
$\cup\{\{v_{i,j},v_{i,j+1}\}\mid 1\leq i\leq n,1\leq j\leq 2q_{i}-1\}$
$\cup\{\{v,v_{i,1}\},\{v,v_{i,2q_{i}}\}\mid 1\leq i\leq n\}.$
Let $B^{0}$ and $B^{s}$ denote the compact graph $\mathbf{B}_{\underline{p}:\underline{q}}^{0}$ and $\mathbf{B}_{\underline{p}:\underline{q}}^{s}$ respectively, where $s\geq 2$ is an even number. We may write by Conventions 2.6
${\mathbb{R}}^{|V(B^{0})|}=\{\sum\limits_{i=1}^{m}\sum\limits_{j=1}^{2p_{i}}a_{i,j}{\mathbf{u}}_{i,j}+\sum\limits_{i=1}^{m}\sum\limits_{j=1}^{2p_{i}}b_{i,j}{\mathbf{v}}_{i,j}+a{\mathbf{u}}+b{\mathbf{v}}\mid\mbox{ all }a_{i,j},b_{i,j},a,b\in{\mathbb{R}}\}.$
We showed in [20, Section 5] that the following vectors
$\alpha_{\ell}:=\sum\limits_{i=1}^{m}\sum\limits_{j=1}^{2p_{i}}{\mathbf{u}}_{i,j}+\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{2q_{i}}{\mathbf{v}}_{i,j}+{\mathbf{v}}+(2\ell+1){\mathbf{u}},\quad\ell=0,\ldots,m-1$
and
$\beta_{\ell}:=\sum\limits_{i=1}^{m}\sum\limits_{j=1}^{2p_{i}}{\mathbf{u}}_{i,j}+\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{2q_{i}}{\mathbf{v}}_{i,j}+{\mathbf{u}}+(2\ell+1){\mathbf{v}},\quad\ell=1,\ldots,n-1$
are all the minimal vectors of $\mathrm{relint}({\mathbb{R}}_{+}(B^{0}))\cap{\mathbb{Z}}^{|V(B^{0})|}$.
Similarly we may write
${\mathbb{R}}^{|V(B^{s})|}=\{\sum\limits_{i=1}^{m}\sum\limits_{j=1}^{2p_{i}}a_{i,j}{\mathbf{u}}_{i,j}+\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{2q_{i}}b_{i,j}{\mathbf{v}}_{i,j}+\sum\limits_{i=1}^{s-1}c_{i}{\mathbf{w}}_{i}+a{\mathbf{u}}+b{\mathbf{v}}\mid\mbox{ all }a_{i,j},b_{i,j},c_{i},a,b\in{\mathbb{R}}\}.$
and showed in [20, Section 5] that the following vectors
$\alpha_{\ell}:=\sum\limits_{i=1}^{m}\sum\limits_{j=1}^{2p_{i}}{\mathbf{u}}_{i,j}+\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{2q_{i}}{\mathbf{v}}_{i,j}+\sum\limits_{i=1}^{s-1}{\mathbf{w}}_{i}+{\mathbf{v}}+2\ell{\mathbf{u}},\ell=1,\ldots,m$
and
$\beta_{\ell}:=\sum\limits_{i=1}^{m}\sum\limits_{j=1}^{2p_{i}}{\mathbf{u}}_{i,j}+\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{2q_{i}}{\mathbf{v}}_{i,j}+\sum\limits_{i=1}^{s-1}{\mathbf{w}}_{i}+{\mathbf{u}}+2\ell{\mathbf{v}},\ell=1,\ldots,n$
are all the minimal vectors of $\mathrm{relint}({\mathbb{R}}_{+}(B^{s}))\cap{\mathbb{Z}}^{|V(B^{s})|}$.
Lemma 3.3.
For each compact graph $H$ of type 2,
  1. (1)
    If $H=\mathbf{B}_{\underline{p}:\underline{q}}^{0}$, then
    $\mathcal{N}_{H}=\{\Theta_{H}u^{2m}v^{2\ell}\mid\ell=1,\ldots,n-1\}\sqcup\{\Theta_{H}u^{2\ell}v^{2n}\mid\ell=1,\ldots,m\}.$
  2. (2)
    If $H=\mathbf{B}_{\underline{p}:\underline{q}}^{s}$, then
    $\mathcal{N}_{H}=\{\Theta_{H}u^{2m+1}v^{2\ell}\mid\ell=1,\ldots,n\}\sqcup\{\Theta_{H}u^{2\ell}v^{2n+1}\mid\ell=1,\ldots,m\}.$
Proof.
If $H=\mathbf{B}_{\underline{p}:\underline{q}}^{0}$, then $D_{H}=\Theta_{H}^{2}u^{2m+1}v^{2n+1}$, and if $H=\mathbf{B}_{\underline{p}:\underline{q}}^{s}$, the expression for $D_{H}$ becomes $\Theta_{H}^{2}u^{2m+2}v^{2n+2}$. Given the minimal generators within the relative interior of the edge cones of $H$ as enumerated previously, the result is straightforwardly deduced.
Analogously in Lemma 3.3, the notation $u,v$ utilized in the expression of $\mathcal{N}_{H}$ represents the two big vertices of the graph $H$. Nevertheless, upon considering $H$ as an induced subgraph of a compact graph of type 3, there arises a possibility that the two big vertices of $H$ might alternately be denoted as $w,u$ or $w,v$. In such cases, it becomes necessary to substitute $w,u$ with either $v$ or $w,v$ accordingly.
3.3. type three
Given positive integers $m,n,k$, as well as the tuples $\underline{p}=(p_{1},\ldots,p_{m})$, $\underline{q}=(q_{1},\ldots,q_{n})$ and $\underline{r}=(r_{1},\ldots,r_{k})$, we denote the graph $\mathbf{C}_{\underline{p}:\underline{q}:\underline{r}}$ as $C$ for brevity. Here, $p_{i},q_{i},r_{i}$ are all positive integers. By definition, we may assume $C$ has vertex set
$V(C)=$$\{\{u,v,w\}\cup\{u_{i,j}\mid 1\leq i\leq m,1\leq j\leq 2p_{i}\}$
$\cup\{v_{i,j}\mid 1\leq i\leq n,1\leq j\leq 2q_{i}\}\cup\{w_{i,j}\mid 1\leq i\leq k,1\leq j\leq 2r_{i}\},$
and edge set
$E(C)=$$\{\{u_{i,j},u_{i,j+1}\}\mid 1\leq i\leq m,1\leq j\leq 2p_{i}-1\}$
$\cup\{\{u,u_{i,1}\},\{u,u_{i,2p_{i}}\}\mid 1\leq i\leq m\}$
$\cup\{\{v_{i,j},v_{i,j+1}\}\mid 1\leq i\leq n,1\leq j\leq 2q_{i}-1\}$
$\cup\{\{v,v_{i,1}\},\{v,v_{i,2q_{i}}\}\mid 1\leq i\leq n\}$
$\cup\{\{w_{i,j},w_{i,j+1}\}\mid 1\leq i\leq k,1\leq j\leq 2r_{i}-1\}$
$\cup\{\{w,w_{i,1}\},\{w,w_{i,2r_{i}}\}\mid 1\leq i\leq k\}$
$\cup\{\{u,v\},\{v,w\},\{w,u\}\}.$
It follows from Conventions 2.6 that
${\mathbb{R}}^{|V(C)|}=\{\sum\limits_{i=1}^{m}\sum\limits_{j=1}^{2p_{i}}{a_{i,j}{\mathbf{u}}_{i,j}}+a{\mathbf{u}}+\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{2q_{i}}{b_{i,j}{\mathbf{v}}_{i,j}}+b{\mathbf{v}}+\sum\limits_{i=1}^{k}\sum\limits_{j=1}^{2r_{i}}{c_{i,j}{\mathbf{w}}_{i,j}}+c{\mathbf{w}}\mid$
$a_{i,j},a,b_{i,j},b,c_{i,j},c\in{\mathbb{R}}\mbox{ for all }i,j\}.$
According to [20, Subsection 1], we have the following result.
Lemma 3.4.
The minimal vectors of $\mathrm{relint}({\mathbb{R}}_{+}(C))\cap\mathbb{Z}^{|V(C)|}$ are following $m+n+k$ vectors: $\alpha_{\ell}$ for $\ell=1,\ldots,m,$ $\beta_{\ell}$ for ${\ell}=1,\ldots,n$ and $\gamma_{\ell}$ for ${\ell}=1,\ldots,k$. Here, For $\ell=1,\ldots,m$,
$\alpha_{\ell}:=\sum\limits_{i=1}^{m}\sum\limits_{j=1}^{2p_{i}}{\mathbf{u}}_{i,j}+\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{2q_{i}}{\mathbf{v}}_{i,j}+\sum\limits_{i=1}^{k}\sum\limits_{j=1}^{2r_{i}}{\mathbf{w}}_{i,j}+{\mathbf{v}}+{\mathbf{w}}+2{\ell}{\mathbf{u}}.$
For ${\ell}=1,\ldots,n$,
$\beta_{\ell}=\sum\limits_{i=1}^{m}\sum\limits_{j=1}^{2p_{i}}{\mathbf{u}}_{i,j}+\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{2q_{i}}{\mathbf{v}}_{i,j}+\sum\limits_{i=1}^{k}\sum\limits_{j=1}^{2r_{i}}{\mathbf{w}}_{i,j}+{\mathbf{w}}+{\mathbf{u}}+2{\ell}{\mathbf{v}}.$
For ${\ell}=1,\ldots,k$,
$\gamma_{\ell}=\sum\limits_{i=1}^{m}\sum\limits_{j=1}^{2p_{i}}{\mathbf{u}}_{i,j}+\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{2q_{i}}{\mathbf{v}}_{i,j}+\sum\limits_{i=1}^{k}\sum\limits_{j=1}^{2r_{i}}{\mathbf{w}}_{i,j}+{\mathbf{u}}+{\mathbf{v}}+2{\ell}{\mathbf{w}}.$
In other words, the canonical module $\omega_{\mathbb{K}[C]}$ of $\mathbb{K}[C]$ is a monomial ideal of $\mathbb{K}[C]$ generated by
$\Theta_{C}u^{2}vw,\Theta_{C}u^{4}vw,\ldots,\Theta_{C}u^{2m}vw,\quad\Theta_{C}uv^{2}w,\Theta_{C}uv^{4}w,\ldots,\Theta_{C}uv^{2n}w$
and
$\quad\Theta_{C}uvw^{2},\Theta_{C}uvw^{4},\ldots,\Theta_{C}uvw^{2k}.$
The computation of $D_{C}$ is straightforward, resulting in $D_{C}=\Theta_{C}u^{2m+2}v^{2n+2}w^{2k+2}$. From this equality and applying Formula (1) in Subsection 2.5, we obtain the following result immediately.
Lemma 3.5.
If $H=\mathbf{C}_{\underline{p}:\underline{q}:\underline{r}}$, then
$\mathcal{N}_{H}=\{\Theta_{H}u^{2\ell}v^{2n+1}w^{2k+1}\mid\ell=1,\ldots,m\}\sqcup\{\Theta_{H}u^{2m+1}v^{2\ell}w^{2k+1}\mid\ell=1,\ldots,n\}$
$\sqcup\{\Theta_{H}u^{2m+1}v^{2n+1}w^{2\ell}\mid\ell=1,\ldots,k\}.$
Recall that $\mathfrak{t}(G)$ represents the count of minimal cycles in graph $G$. The lemma below highlights properties of $\mathcal{N}_{H}$ that are useful to our discussion.
Lemma 3.6.
Let $H$ be a compact graph with the projection dimension of its edge rings denoted as $p$. Then the following statements hold:
  1. (1)
    The projective dimension $p$ is equal to $\mathfrak{t}(H)-1$, which is also the cardinality of $\mathcal{N}_{H}$, i.e., $p=\mathfrak{t}(H)-1=|\mathcal{N}_{H}|$;
  2. (2)
    For each element $\alpha$ belonging to $\mathcal{N}_{H}$, its support set $\mathrm{supp}(\alpha)$ is identical to the vertex set $V(H)$ of the graph $H$.
Proof.
The equality $p=\mathfrak{t}(H)-1$ is a direct consequence of [20, Corollary 4.2]. The remaining assertions are derived by examining the members of $\mathcal{N}_{H}$.
4. Betti numbers of compact graphs
In this section, we present an explicit formula for the multi-graded Betti-numbers of $\mathbb{K}[G]$, whenever $G$ is a compact graph.
Lemma 4.1.
Let $G$ be a compact graph. Then there is a monomial order $<$ such that $\beta_{i}(\mathrm{in}_{<}(I_{G}))=(i+1)\binom{\mathfrak{t}(G)}{i+2}$ for all $i\geq 0$. Thus, for all $i\geq 1$, one has
$\beta_{i}(\mathbb{K}[G])\leq\beta_{i}(\mathbb{K}[E(G)]/\mathrm{in}_{<}(I_{G}))=i\binom{\mathfrak{t}(G)}{i+1}.$
Proof.
The result is immediate from [20, Theorem 4.1] and Theorem 2.1.
Proposition 4.2.
Let $G$ be a compact graph of either type 1 or type 2. Then, for any $10\leq i\leq\mathfrak{t}(G)-$ 1, we have
$\beta_{i}(\mathbb{K}[G])=\beta_{i}(\mathbb{K}[E(G)]/\mathrm{in}_{<}(I_{G})).$
Proof.
If $i=0$, then $\beta_{i}(\mathbb{K}[G])=\beta_{i}(\mathbb{K}[E(G)]/\mathrm{in}_{<}(I_{G}))=1.$ Fix $11\leq i\leq\mathfrak{t}(G)-$ 1. Since there are $\mathfrak{t}(G)$ minimal cycles, there are $\binom{\mathfrak{t}(G)}{i+1}$ compact induced subgraphs of $G$ that contain exactly $i+1$ minimal cycles of $G$. These induced subgraphs are of either type one or type two, and are denoted by $H_{1},\ldots,H_{r}$, respectively. Here, $r=\binom{\mathfrak{t}(G)}{i+1}$. Note that $\beta_{i,\alpha}(\mathbb{K}[H_{\ell}])=1$ for any $\alpha\in\mathcal{N}_{H_{\ell}}$ by Lemmas 3.2 and 3.3, it follows that $\beta_{i,\alpha}(\mathbb{K}[G])=1$ for any $\alpha\in\bigcup_{\ell=1}^{r}\mathcal{N}_{H_{\ell}}$ in view of Lemma 2.7. Put $\mathcal{N}_{i}(G)=\bigcup_{\ell=1}^{r}\mathcal{N}_{H_{\ell}}$.
Note that $\mathrm{supp}(\alpha)=V(H_{k})$ for any $\alpha\in\mathcal{N}_{H_{k}}$, see Lemma 3.6, we have $\mathcal{N}_{H_{1}},\ldots,\mathcal{N}_{H_{r}}$ are pairwise disjoint. Hence,
$|\mathcal{N}_{i}(G)|=\sum_{\ell=1}^{r}|\mathcal{N}_{H_{\ell}}|=ir=i\binom{\mathfrak{t}(G)}{i+1}.$
This implies
$i\binom{\mathfrak{t}(G)}{i+1}=\sum_{\alpha\in\mathcal{N}_{i}(G)}\beta_{i,\alpha}(\mathbb{K}[G])\leq\beta_{i}(\mathbb{K}[G])\leq\beta_{i}(\mathbb{K}[E(G)]/\mathrm{in}_{<}(I_{G}))=i\binom{\mathfrak{t}(G)}{i+1},$
completing the proof.
From the proof of Proposition 4.2, it is evident that
$\beta_{i}(\mathbb{K}[G])=\sum_{\alpha\in\mathcal{N}_{i}(G)}\beta_{i,\alpha}(\mathbb{K}[G]).$
In particular, if $\beta_{i,\alpha}(\mathbb{K}[G])\neq 0$, then $\alpha\in\mathcal{N}_{i}(G)$. Therefore, we can derive the multi-graded and graded Betti numbers of $\mathbb{K}[G]$ as follows.
Corollary 4.3.
Let $G$ be a compact graph of either type 1 or type 2. For any monomial $\alpha$ in $\mathbb{K}[G]$, we have
$\beta_{i,\alpha}(\mathbb{K}[G])=\left\{\begin{array}[]{ll}1,&\hbox{$\alpha\in\mathcal{N}_{i}(G)$;}\\ 0,&\hbox{otherwise.}\end{array}\right.$
In particular, $\beta_{i,j}(\mathbb{K}[G])$ is equal to the number of the set $\{\alpha\in\mathcal{N}_{i}(G)\mid|\mathrm{deg}(\alpha)|=2j\}.$
Corollary 4.3 has an interesting implication.
Corollary 4.4.
Let $G$ be a compact graph of either type 1 or type 2. Let $\alpha$ be a monomial of $\mathbb{K}[G]$. Then for every $11\leq i\leq m-$ 1 with $\beta_{i,\alpha}(\mathbb{K}[G])\neq 0$, the induced subgraph on $\mathrm{supp}(\alpha)$ of $G$ contains exactly $i+1$ minimal cycles of $G$.
We now turn to compact graphs of type 3. In the case when $G$ is of type one or type two, there are exactly $\binom{\mathfrak{t}(G)}{i+1}$ distinct induced subgraphs of $G$ which contains exactly $i+1$ minimal cycles for all $11\leq i\leq\mathfrak{t}(G)-$ 1. This is the reason why the proof of Theorem 4.2 holds. But this is not the case when considering compact graphs of type 3. Let us see the following example.
Example 4.5.
Assume that $G=\mathbf{C}_{(1):(1):(1,1)}$, as shown in Figure 1. Then $G$ contains 5 induced cycles. But there are only 4 distinct induced subgraphs of $G$ which contains 4 induced cycles. This is because the 4 cycles around the triangle formed three big vertices $\{u,v,w\}$ does not form an induced subgraph of $G$.
$x_{1}$$x_{2}$$w_{1}$$w_{2}$$u_{2}$$u_{1}$$v_{1}$$v_{2}$$u$$v$$w$
Figure 1. The graph $\mathbf{C}_{(1):(1):(1,1)}$
Based on the observation in Example 4.5, it is clear that the approach used to determine the multi-graded Betti numbers for edge rings of compact graphs of type 1 or type 2 falls short when applied to compact graphs of type 3. To address this issue for type 3 compact graphs, we need to compute the second top graded Betti numbers, leveraging Lemma 2.2. To apply Lemma 2.2, we make an observation related to the computation of simplicial complexes $\Gamma_{h}(I)$. Notably, we adhere to the notation established in Lemma 2.2 throughout our analysis. Let $\Omega$ be a set comprised of a number of monomials from $T$, and assume that the ideal $I$ is generated by all the monomials in $\Omega$. For each $\alpha\in\Omega$, we introduce a simplicial complex $\Gamma_{h}^{\alpha}(I)$ as follows. If $\alpha$ divides $Y^{h}$ and $Y^{h}/\alpha=y_{1}^{a_{1}}y_{2}^{a_{2}}\cdots y_{n}^{a_{n}}$ for some $a_{i}\geq 0$, then $\Gamma_{h}^{\alpha}(I)$ is specified as the complex: $F$ is a facet of $\Gamma_{h}^{\alpha}(I)$Longleftrightarrow there are $a_{i}\geq 0$ for $i\in[n]$ such that $Y^{h}=\alpha y_{1}^{a_{1}}y_{2}^{a_{2}}\cdots y_{n}^{a_{n}}$ and $F=\{i\in[n]\mid a_{i}\neq 0\}.$ Otherwise, $\Gamma_{h}^{\alpha}(I)$ is simply an empty set. We can easily check that
$\Gamma_{h}(I)=\bigcup_{\alpha\in\Omega}\Gamma_{h}^{\alpha}(I).$
Example 4.6.
Let $\mathbb{K}[G]$ be the edge ring of $G=\mathbf{C}_{(1):(1):(1,1)}$. Note that all the vertices of $G$ are labeled, as shown in Figure 1. Thus, $\Theta_{G}=x_{1}x_{2}w_{1}w_{2}u_{1}u_{2}v_{2}v_{2}$. In the following, we denote $\Theta_{G}$ by $\Theta$ for short. Moreover, in view of Lemma 3.4, the canonical module $\omega_{\mathbb{K}[G]}$, as an ideal of $\mathbb{K}[G]$, is generated by monomials
$\Theta uvw^{2},\Theta uvw^{4},\Theta u^{2}vw,\Theta uv^{2}w.$
The four monomials are denoted by $\beta_{i}$ for $i=1,\cdots,4$. Put $\alpha_{1}=\Theta u^{2}v^{2}w^{2}$ and $\alpha_{2}=\Theta u^{2}v^{2}w^{4}.$ Denote $uv,vw,wu$ by $y_{1},y_{2},y_{3}$ respectively.
Since $\Theta u^{2}v^{2}w^{2}=\beta_{1}y_{1}$, it follows that $\Gamma_{\alpha_{1}}^{\beta_{1}}(\omega_{\mathbb{K}[G]})=\langle\{1\}\rangle$;
Since $\beta_{2}$ does not divide $\Theta u^{2}v^{2}w^{2}$, it follows that $\Gamma_{\alpha_{1}}^{\beta_{2}}(\omega_{\mathbb{K}[G]})=\emptyset$;
Since $\Theta u^{2}v^{2}w^{2}=\beta_{3}y_{2}$, it follows that $\Gamma_{\alpha_{1}}^{\beta_{3}}(\omega_{\mathbb{K}[G]})=\langle\{2\}\rangle$;
Since $\Theta u^{2}v^{2}w^{2}=\beta_{4}y_{3}$, it follows that $\Gamma_{\alpha_{1}}^{\beta_{4}}(\omega_{\mathbb{K}[G]})=\langle\{3\}\rangle$.
Thus, $\Gamma_{\alpha_{1}}(\omega_{\mathbb{K}[G]})=\langle\{1\},\{2\},\{3\}\rangle$, and so $\dim_{\mathbb{K}}\widetilde{H}_{0}(\Gamma_{\alpha_{1}}(\omega_{\mathbb{K}[G]}),\mathbb{K})=2$.
Similarly, we have $\Gamma_{\alpha_{2}}^{\beta_{1}}(\omega_{\mathbb{K}[G]})=\langle\{2,3\}\rangle$, $\Gamma_{\alpha_{2}}^{\beta_{2}}(\omega_{\mathbb{K}[G]})=\langle\{1\}\rangle$ and $\Gamma_{\alpha_{2}}^{\beta_{3}}(\omega_{\mathbb{K}[G]})=\Gamma_{\alpha_{2}}^{\beta_{4}}=\emptyset$. Hence, $\Gamma_{\alpha_{2}}(\omega_{\mathbb{K}[G]})=\langle\{2,3\},\{1\}\rangle$ and it follows $\dim_{\mathbb{K}}\widetilde{H}_{0}(\Gamma_{\alpha_{2}}(\omega_{\mathbb{K}[G]}),\mathbb{K})=1$.
Note that $D_{G}=\Theta^{2}u^{4}v^{4}w^{6}$. Since $D_{G}/\alpha_{1}=\Theta u^{2}v^{2}w^{4}$ and $D_{G}/\alpha_{2}=\Theta u^{2}v^{2}w^{2}$, it follows from Lemma 2.2 as well as Lemma 2.3 that
$\beta_{3,\Theta u^{2}v^{2}w^{4}}(\mathbb{K}[G])=\beta_{1,\alpha_{1}}(\omega_{\mathbb{K}[G]})=2$
and
$\beta_{3,\Theta u^{2}v^{2}w^{2}}(\mathbb{K}[G])=\beta_{1,\alpha_{2}}(\omega_{\mathbb{K}[G]})=1.$
There are precisely four induced subgraphs of $G$, that consist of four minimal cycles of $G$. These four induced subgraphs are induced on $V(G)\setminus\{x_{1},x_{2}\}$ and $V(G)\setminus\{w_{1},w_{2}\}$, $V(G)\setminus\{u_{1},u_{2}\}$ and $V(G)\setminus\{v_{1},v_{2}\}$, denoted by $G_{1}$ and $G_{2}$, $G_{3}$ and $G_{4}$, respectively. According to the Lemmas 3.4 and 3.6, we have
$\mathcal{N}_{G_{1}}=\Theta_{G_{1}}\{u^{2}v^{3}w^{3},u^{3}v^{2}w^{3},u^{3}v^{3}w^{2}\},\quad\mathcal{N}_{G_{2}}=\Theta_{G_{2}}\{u^{2}v^{3}w^{3},u^{3}v^{2}w^{3},u^{3}v^{3}w^{2}\}$
and
$\mathcal{N}_{G_{3}}=\Theta_{G_{3}}\{v^{3}w^{2},v^{3}w^{4},v^{2}w^{5}\},\quad\mathcal{N}_{G_{4}}=\Theta_{G_{4}}\{u^{3}w^{2},u^{3}w^{4},u^{2}w^{5}\}.$
Here, for a set $A$ of monomials and for a monomial $\alpha$, the notation $\alpha A$ denotes the set $\{\alpha a\mid a\in A\}$.
Combining the above facts, the sum of Betti numbers $\beta_{3,\alpha}(\mathbb{K}[G])$ with $\mbox{supp}(\alpha)=V(G)$ is greater than or equal to 3, and the sum of Betti numbers $\beta_{3,\alpha}(\mathbb{K}[G])$ with $\mbox{supp}(\alpha)=V(G_{i})$ is also greater than or equal to $\geq$ 3 for $i=1,2,3,4$. Therefore, it follows that
$1515\leq\beta_{3}(\mathbb{K}[G])\leq\beta_{3}(\mathbb{K}[E(G)]/\mathrm{in}_{<}(I_{G}))=(2+1)\binom{5}{4}=$ 15 .
Hence, $\beta_{3}(\mathbb{K}[G])=\beta_{3}(\mathbb{K}[E(G)]/\mathrm{in}_{<}(I_{G}))$, and the multi-graded Betti numbers $\beta_{3,\alpha}(\mathbb{K}[G])$ could also be written down.
The approach outlined in Example 4.6 can be applied to the general case.
Definition 4.7.
Let $H$ be the graph $\mathbf{C}_{\underline{p}:\underline{q}:\underline{r}}$ with its vertices of $H$ labeled as in Subsection 3.3. We define $\mathcal{M}_{H}^{1}$, $\mathcal{M}_{H}^{2}$ and $\mathcal{M}_{H}$ to be the following sets:
$\mathcal{M}^{1}_{H}=$$\{\Theta_{H}u^{2\ell}v^{2n}w^{2k}\mid\ell=1,\ldots,m-1\}\sqcup\{\Theta_{H}u^{2m}v^{2\ell}w^{2k}\mid\ell=1,\ldots,n-1\}$
$\sqcup\{\Theta_{H}u^{2m}v^{2n}w^{2\ell}\mid\ell=1,\ldots,k-1\},$
$\mathcal{M}_{H}^{2}=\{\Theta_{H}u^{2m}v^{2n}w^{2k}\},$
and
$\mathcal{M}_{H}=\mathcal{M}_{H}^{1}\sqcup\mathcal{M}_{H}^{2}.$
Lemma 4.8.
Let $H=\mathbf{C}_{\underline{p}:\underline{q}:\underline{r}}$ denote a compact graph of type 3 and let $p=m+n+k$ represent the projective dimension of $\mathbb{K}[H]$. Then,
$\beta_{p-1,\alpha}(\mathbb{K}[H])=\left\{\begin{array}[]{ll}1,&\hbox{$\alpha\in\mathcal{M}^{1}_{H}$;}\\ 2,&\hbox{$\alpha\in\mathcal{M}^{2}_{H}$.}\end{array}\right.$
In particular, we have $\sum\limits_{\alpha}\beta_{p-1,\alpha}(\mathbb{K}[H])\geq p-1$, where the summation is over all $\alpha\in\mathrm{Mon}(\mathbb{K}[H])$ with $\mathrm{supp}(\alpha)=V(H)$.
Proof.
Denote $uv,vw$ and $wu$ by $y_{1},y_{2}$ and $y_{3}$ respectively. Let $\omega$ denote the canonical module $\omega_{\mathbb{K}[H]}$. Note that $\omega$ is a monomial ideal of $\mathbb{K}[H]$, whose generators are given in Lemma 3.4. Put $\alpha=\Theta u^{2}v^{2}w^{2}$, where $\Theta$ is used as a shorthand for $\Theta_{H}$. Then
$\Gamma_{\alpha}^{\Theta u^{2}vw}(\omega)=\langle\{2\}\rangle,\Gamma_{\alpha}^{\Theta uv^{2}w}(\omega)=\langle\{3\}\rangle,\Gamma_{\alpha}^{\Theta uvw^{2}}(\omega)=\langle\{1\}\rangle.$
For other generators $\beta$ of $\omega$, one easily sees that $\Gamma_{\alpha}^{\beta}(\omega)=\emptyset$. This implies $\Gamma_{\alpha}(\omega)=\langle\{2\},\{3\},\{1\}\rangle$ and so $\beta_{1,\alpha}(\omega)=2$ by Lemma 2.2. In the same vein, we can prove
$\beta_{1,\Theta u^{2\ell}v^{2}w^{2}}(\omega)=1\mbox{ for }\ell=2,3,\ldots,m,$
$\beta_{1,\Theta u^{2}v^{2\ell}w^{2}}(\omega)=1\mbox{ for }\ell=2,3,\ldots,n,$
and
$\beta_{1,\Theta u^{2}v^{2}w^{2\ell}}(\omega)=1\mbox{ for }\ell=2,3,\ldots,k.$
Note that $D_{H}=\Theta^{2}u^{2m+2}v^{2n+2}w^{2k+2},$ the first assertion follows from Lemma 2.3. For the second assertion, it follows from the facts that $|\mathcal{M}^{1}_{H}|=m+n+k-3$, $|\mathcal{M}^{2}_{H}|=1$, and $\mathrm{supp}(\alpha)=V(H)$ for each $\alpha\in\mathcal{M}^{1}_{H}\sqcup\mathcal{M}^{2}_{H}$.
It is important to note that the inequality stated in the above lemma does not apply to compact graphs of type one or type two. To elaborate further, we present the following remark:
Remark 4.9.
Assume $H$ is a compact graph of type 1 or type 2 and denote by $p$ the projective dimension of $\mathbb{K}[H]$. Then $\beta_{p-1,\alpha}(\mathbb{K}[H])=0$ holds true for each $\alpha\in\mathrm{Mon}(\mathbb{K}[H])$ with $\mathrm{supp}(\alpha)=V(H).$
In fact, if $\beta_{p-1,\alpha}(\mathbb{K}[H])\neq 0$, then the induced subgraph of $H$ on $\mathrm{supp}(\alpha)$ contains $p$ minimal cycles by Corollary 4.4. This is a contradiction since $H$ has $p+1$ minimal cycles.
Proposition 4.10.
Let $G=\mathbf{C}_{\underline{p}:\underline{q}:\underline{r}}$ denote a compact graph of type 3. Then
$\beta_{i}(\mathbb{K}[G])=\beta_{i}(\mathbb{K}[E(G)]/\mathrm{in}_{<}(I_{G}))$
for all $i\geq 0$.
Proof.
We categorize the minimal cycles of $G$ into four distinct classes. The cycle encompassing the three big vertices is designated as the big cycle. A minimal cycle that is distinct from the big cycle but includes the vertex $u$ (or $v$, or $w$ again) is referred to as a cycle belonging to $u$ (or $v$, or $w$ respectively).
Fix $i\geq 1$. Then the number of compact subgraphs of $G$ that contain exactly $i+1$ minimal cycles is given by $\binom{m+n+k+1}{i+1}$. But not all of these subgraphs are induced subgraphs. Consider a subgraph $H$ of $G$ and let $C_{1},\ldots,C_{i+1}$ represent all its minimal cycles. Then one easily sees that $H$ is not an induced subgraph if and only if
  • None of $C_{j}$ for $j=1,\ldots,i+1$ is the big cycle;
  • For each $x\in\{u,v,w\}$, there is at least one index $j\in[i+1]$ such that $C_{j}$ is a minimal cycle belonging to $x$.
Hence, there are $\sum\limits_{a+b+c=i+1}^{a,b,c\geq 1}\binom{m}{a}\binom{n}{b}\binom{k}{c}$ such subgraphs and it follows that
$\sum_{\alpha}\beta_{i,\alpha}(\mathbb{K}[G])\geq(\binom{m+n+k+1}{i+1}-\sum\limits_{a+b+c=i+1}^{a,b,c\geq 1}\binom{m}{a}\binom{n}{b}\binom{k}{c})i,$
where the summation is over all $\alpha\in\mathrm{Mon}(\mathbb{K}[G])$ such that the induced subgraph on $\mathrm{supp}(\alpha)$ consisting of $i+1$ minimal cycles. On the other hand, we notice the following fact: if $H$ is an induced subgraph of $G$ containing exactly $i+2$ minimal cycles, say $C_{1},\ldots,C_{i+2}$, then $H$ belongs to type 3 if and only if
  • One of $C_{j}$ for $j=1,\ldots,i+2$ is the big cycle. For simplicity, let us assume that $C_{i+2}$ is the big cycle;
  • For each $x\in\{u,v,w\}$, there is at least one index $j\in[i+1]$ such that $C_{j}$ is a minimal cycle belonging to $x$.
Thus, there are $\sum\limits_{a+b+c=i+1}^{a,b,c\geq 1}\binom{m}{a}\binom{n}{b}\binom{k}{c}$ compact induced subgraphs of $G$ that belong to type 3 and contain exactly $i+2$ minimal cycles. It follows from Lemma 4.8 that
$\sum_{\alpha}\beta_{i,\alpha}(\mathbb{K}[G])\geq\sum\limits_{a+b+c=i+1}^{a,b,c\geq 1}\binom{m}{a}\binom{n}{b}\binom{k}{c}i,$
where the summation is over all $\alpha\in\mathrm{Mon}(\mathbb{K}[G])$ such that the induced subgraph on $\mathrm{supp}(\alpha)$ is a compact graph of type 3 consisting of $i+2$ minimal cycles. Hence, we have
$\beta_{i}(\mathbb{K}[G])\geq\binom{m+n+k+1}{i+1}i.$
Now, the result follows from Lemma 4.1.
From the proof of Proposition 4.10, we see that the inequality in Lemma 4.8 is actually an equality. By combining all the findings we have made so far, we can immediately obtain the main result of this section.
Theorem 4.11.
Let $G$ be a compact graph, and denote by $p=\mathfrak{t}(G)-1$ the projective dimension of $\mathbb{K}[G]$. Consider an integer $i$ satisfying $1\leq i\leq p$. Let $H_{j}$ for $j=1,\ldots,s$ represent all the compact induced subgraphs of $G$ having exactly $i+1$ minimal cycles, and let $L_{j}$ for $j=1,\ldots,t$ be all the compact induced subgraphs of $G$ belonging to type 3 with $i+2$ minimal cycles. Define the following sets:
$\mathcal{N}_{i}(G)=\bigcup_{j=1}^{s}\mathcal{N}_{H_{j}},\quad\mathcal{M}_{i}^{1}(G)=\bigcup_{j=1}^{t}\mathcal{M}_{L_{j}}^{1},\quad\text{and}\quad\mathcal{M}_{i}^{2}(G)=\bigcup_{j=1}^{t}M_{L_{j}}^{2}.$
Then, the $i$-th multi-graded Betti numbers of $\mathbb{K}[G]$ can be expressed as follows:
$\beta_{i,\alpha}(\mathbb{K}[G])=\begin{cases}2,&\text{if }\alpha\in\mathcal{M}_{i}^{2}(G),\\ 1,&\text{if }\alpha\in\mathcal{M}_{i}^{1}(G)\cup\mathcal{N}_{i}(G),\\ 0,&\text{otherwise}.\end{cases}$
5. Betti numbers of $\mathbf{G}_{\underline{\ell}}$
In this section, we introduce the concept of multi-path graphs and then provide an explicit formula for computing the multi-graded Betti numbers associated with their edge rings.
Definition 5.1.
Let $m\geq 2$ be an integer, and ¯$\underline{\ell}=(\ell_{1},\ldots,\ell_{m})$ a vector in $\mathbb{Z}_{>0}^{m}$. We define $\mathbf{G}_{\underline{\ell}}$ as a simple graph consisting of $m$ paths of lengths $\ell_{1},\ldots,\ell_{m}$, respectively, originating from a common vertex and converging to another common vertex distinct from the starting point, and intersecting only at these endpoints. We call $\mathrm{\mathbf{G}}_{\underline{\ell}}$ to be a multi-path graph. It is classified as even type if all $\ell_{i}$ are even, and as odd type if all $\ell_{i}$ are odd. Conversely, a multi-path graph is designated as mixed type if it falls outside both even and odd categories. Furthermore, a multi-path graph is designated as true-mixed type when it possesses at least two even $\ell_{i}$ and at least two odd $\ell_{i}$. The last notion is justified by Lemma 5.12.
The study of the edge ring of a multi-path graph can be reduced to the investigation of the edge rings of multi-path graphs of both even and odd types, as demonstrated by the following lemma.
Lemma 5.2.
Let $G=\mathbf{G}_{\underline{\ell}}$ be a multi-path graph of mixed type, with $\ell_{1},\ldots,\ell_{s}$ being even and $\ell_{s+1},\ldots,\ell_{m}$ being odd. Define $G_{1}$ as the multi-path graph $\mathbf{G}_{(\ell_{1},\ldots,\ell_{s})}$, and $G_{2}$ as $\mathbf{G}_{(\ell_{s+1},\ldots,\ell_{m})}$. Then the edge ring of $G$ decomposes as the tensor product of the edge rings of $G_{1}$ and $G_{2}$ over the base field $\mathbb{K}$, i.e.,
$\mathbb{K}[G]=\mathbb{K}[G_{1}]\bigotimes_{\mathbb{K}}\mathbb{K}[G_{2}].$
Proof.
By [16, Lemma 3.2], every primitive even closed walks of $G$ is contained in either $G_{1}$ or $G_{2}$. It follows that the defining ideal $\mathbb{K}[G]$ is the sum of the defining ideals of $\mathbb{K}[G_{1}]$ and $\mathbb{K}[G_{2}]$, yielding the desired decomposition.
As a direct consequence, we have for any $i\geq 0$ and any monomial $\alpha\in\mathbb{K}[G]$, the following formula holds:
(3) $\beta_{i,\alpha}(\mathbb{K}[G])=\sum_{j+k=i}^{\beta\gamma=\alpha}\beta_{j,\beta}(\mathbb{K}[G_{1}])\cdot\beta_{k,\gamma}(\mathbb{K}[G_{2}])$
5.1. Odd type
Let $G=\mathbf{G}_{2\underline{\ell}+1}$ be a multi-path graph of odd type, where $\ell=(\ell_{1},\ldots,\ell_{m})\in\mathbb{Z}_{\geq 0}^{m}$. Note that there is at most one $j$ such that $\ell_{j}=0$. If $\ell_{j}=0$ for some $j$, then the Betti number of the edge ring of $\mathbf{G}_{2\underline{\ell}+1}$ can be calculated using the Betti splitting, see [5, Corollary 3.12] for details. So it is reasonable to assume all $\ell_{i}\geq 1$ from the begining. We may assume that the vertex set of $G$ is given by
$V(G)=\{v_{1},v_{2}\}\cup\{u_{i,j}\mid 1\leq i\leq m,1\leq j\leq 2\ell_{i}\}$
while the edge set of $G$ is given by
[…]
$\quad\cup\{\{u_{i,j},u_{i,j+1}\mid 1\leq i\leq m,1\leq j\leq 2\ell_{i}-1\}\}.$
Let us compare the edge ring $\mathbb{K}[G]$ with the edge ring $\mathbb{K}[A]$, where $A=\mathbf{A}_{\ell}$ is the compact graph of type one. There exists a natural bijection between the edge sets $E(G)$ and $E(A)$. Moreover, through this bijection, the primitive even closed walks of $G$ correspond to the primitive even closed walks of $A$. From this correspondence, it follows that the edge rings $\mathbb{K}[G]$ and $\mathbb{K}[A]$ possess identical defining ideals. Therefore, as rings, $\mathbb{K}[G]$ is isomorphic to $\mathbb{K}[A]$. However, as multi-graded rings, they are not isomorphic, as $\mathbb{K}[G]$ is $\mathbb{Z}^{|V(G)|}$-graded, while $\mathbb{K}[A]$ is $\mathbb{Z}^{|V(A)|}$-graded, and $|V(G)|=|V(A)|+1$. In view of all these facts, the results pertaining to compact graphs of type one can be seamlessly transferred to the realm of multi-path graphs of odd type, with appropriate adjustments made accordingly.
Lemma 5.3.
If $H=\mathbf{G}_{2\underline{\ell}+1}$, then,
$\mathcal{N}_{H}=\{\Theta_{H}v_{1}^{\ell}v_{2}^{\ell}\mid\ell=1,\ldots,m-1\}.$
Proposition 5.4.
Let $G=\mathbf{G}_{2\underline{\ell}+1}$ be a multi-path graph of odd type. Then, the following assertions hold:
  1. $(1)$
    $\mathrm{pdim}(\mathbb{K}[G])=m-1$,
  2. $(2)$
    For each $i$ with $11\leq i\leq m-$ 1, consider all multi-path induced subgraphs of $G$ consisting of $i+1$ paths, denoted by $H_{1},\ldots,H_{r}$. Put $\mathcal{N}_{i}(G)=\bigcup_{j=1}^{r}\mathcal{N}_{H_{j}}$. The multi-graded Betti numbers of $\mathbb{K}[G]$ are then given by
    $\beta_{i,\alpha}(\mathbb{K}[G])=\begin{cases}1,&\text{if }\alpha\in\mathcal{N}_{i}(G),\\ 0,&\text{otherwise}.\end{cases}$
5.2. Even type
Let $G=\mathbf{G}_{2\underline{\ell}}$ be a multi-path graph of even type. Here, ¯$\underline{\ell}=(\ell_{1},\ldots,\ell_{m})\in\mathbb{Z}_{>0}^{m}$. We may assume the vertex set of $G$ is given by
$V(G)=\{v_{1},v_{2}\}\cup\{u_{i,j}\mid 1\leq i\leq m,1\leq j\leq 2\ell_{i}-1\}$
while the edge set of $G$ is given by
[…]
$\quad\cup\{\{u_{i,j},u_{i,j+1}~{}|~{}1\leq i\leq m,1\leq j\leq 2\ell_{i}-2\}\}.$
We label the edges of $G$ as follows: $e_{i,1}=\{v_{1},u_{i,1}\},e_{i,j}=\{u_{i,j-1},u_{i,j}\},e_{i,2\ell_{i}}=\{v_{2},u_{i,2\ell_{i}-1}\}$, for $i\in[m]$ and $j\in\{2,\ldots,2\ell_{i}-1\}.$
It is evident that $G$ is a bipartite graph with bipartition $(V_{1},V_{2})$, where
$V_{1}=\bigcup\limits_{j=1}^{m}\bigcup\limits_{i=1}^{\ell_{j}}\{u_{j,2i-1}\}$
and
$V_{2}=\{v_{1},v_{2}\}\cup\bigcup\limits_{j=1}^{m}\bigcup\limits_{i=1}^{\ell_{j}}\{u_{j,2i}\}.$
Firstly, we establish an upper bound for the total Betti number of $\mathbb{K}[G]$, which is actually the precise value as shown later.
Lemma 5.5.
Let $G$ be the graph as described above. Then there is a monomial order $<$ such that $\beta_{i}(\mathrm{in}_{<}(I_{G}))\leq(i+1)\binom{m}{i+2}$ for all $i\geq 0$. Thus,
$\beta_{i}(\mathbb{K}[G])\leq\beta_{i}(\mathbb{K}[E(G)]/\mathrm{in}_{<}(I_{G}))\leq i\binom{m}{i+1}.$
Proof.
According to [11, Proposition 4.3], there is a monomial order $<$ such that $\mathrm{in}_{<}(I_{G})=(u_{1},\ldots,u_{\frac{m(m-1)}{2}})$, and the colon ideal $(u_{1},\ldots,u_{j-1}):u_{j}$ is generated by a regular sequence of length $k_{j}$ for $j=2,\ldots,\frac{m(m-1)}{2}$, where
$(k_{2},\ldots,k_{\frac{m(m-1)}{2}})=(\underbrace{{1,1}}_{2},\underbrace{2,2,2}_{3},\ldots,\underbrace{m-2,m-2,\ldots,m-2}_{m-1}).$
In view of Formula (1), we have,
$\beta_{i}(\mathrm{in}_{<}(I_{G}))$ $\leq\sum_{k=1}^{m-2}(k+1)\binom{k}{i}+\binom{0}{i}=(i+1)\sum_{k=1}^{m-2}\binom{k+1}{i+1}+\binom{0}{i}$
$=(i+1)\sum_{k=0}^{m-2}\binom{k+1}{i+1}=(i+1)\binom{m}{i+2}.$
Here, the third-to-last equality follows from the identity $k\binom{n}{k}=n\binom{n-1}{k-1}$, and the last equality follows from the identity $\sum\limits_{j=0}^{n}\binom{j}{k}=\binom{n+1}{k+1}$ for $k\geq 0$. From this bound and Lemma 2.1, the second assertion is immediate.
Next, we compute the minimal generators of the canonical module $\omega_{\mathbb{K}[G]}$.
It is easy to see that $|V(G)|=2\sum\limits_{i=1}^{m}\ell_{i}-m+2$. It follows from Conventions 2.6 that:
${\mathbb{R}}^{|V(G)|}=\{\sum\limits_{i=1}^{m}\sum\limits_{j=1}^{2\ell_{i}-1}{a_{i,j}{\mathbf{u}}_{i,j}}+b_{1}{\mathbf{v}}_{1}+b_{2}{\mathbf{v}}_{2}\mid a_{i,j},b_{i}\in{\mathbb{R}}\mbox{ for all }i,j\}.$
In what follows, we will construct $m-1$ integral vectors in ${\mathbb{R}}^{|V(G)|}$ and then show that they are minimal vectors of $\mathrm{relint}({\mathbb{R}}_{+}(G))\cap\mathbb{Z}^{|V(G)|}$. The construction is as follows:
For $\ell=1,\ldots,m-1$, let
$\alpha_{\ell}:=\sum\limits_{i=1}^{m}\sum\limits_{j=1}^{2\ell_{i}-1}{\mathbf{u}}_{i,j}+\ell{\mathbf{v}}_{1}+(m-\ell){\mathbf{v}}_{2}.$
We now verify that $\alpha_{\ell}\in\mathrm{relint}({\mathbb{R}}_{+}(G))$ for all possible $\ell$.
We see the following:
  • A subset $T$ of $V_{1}$ such that $H_{T}\cap{\mathbb{R}}_{+}(G)$ is the facet of ${\mathbb{R}}_{+}(G)$ if and only if $T$ is one of the following sets:
    1. $(i)$
      $\bigcup\limits_{j=1}^{s}\bigcup\limits_{i=1}^{k_{p_{j}}}\{u_{p_{j},2i-1}\}$, where $\{p_{1},\ldots,p_{s}\}\subseteq[m]$ and $1\leq k_{p_{j}}\leq\ell_{p_{j}}-1;$
    2. $(ii)$
      $\bigcup\limits_{j=1}^{s}\bigcup\limits_{i=k_{p_{j}}}^{\ell_{p_{j}}}\{u_{p_{j},2i-1}\}$, where $\{p_{1},\ldots,p_{s}\}\subseteq[m]$ and $2\leq k_{p_{j}}\leq\ell_{p_{j}};$
    3. $(iii)$
      $\bigcup\limits_{j\neq p}\bigcup\limits_{i=1}^{\ell_{j}}\{u_{j,2i-1}\}$, where $p\in[m];$
    4. $(iv)$
      $\bigcup\limits_{j\neq p}\bigcup\limits_{i=1}^{\ell_{j}}\{u_{j,2i-1}\}\cup\bigcup\limits_{i=1}^{f_{p}}\{u_{p,2i-1}\}$, where $p\in[m]$ and $11\leq f_{p}\leq\ell_{p}-$ 1;
    5. $(v)$
      $\bigcup\limits_{j\neq p}\bigcup\limits_{i=1}^{\ell_{j}}\{u_{j,2i-1}\}\cup\bigcup\limits_{i=k_{p}}^{\ell_{p}}\{u_{p,2i-1}\}$, where $p\in[m]$ and $2\leq k_{p}\leq\ell_{p}$;
    6. $(vi)$
      $\bigcup\limits_{j\neq p}\bigcup\limits_{i=1}^{\ell_{j}}\{u_{j,2i-1}\}\cup\bigcup\limits_{i=1}^{f_{p}}\{u_{p,2i-1}\}\cup\bigcup\limits_{i=k_{p}}^{\ell_{p}}\{u_{p,2i-1}\}$, where $p\in[m]$ and $11\leq f_{p}<k_{p}-1\leq\ell_{p}-$ 1;
    7. $(vii)$
      $\bigcup\limits_{i=f_{j}}^{k_{j}}\{u_{j,2i-1}\}$, where $j\in[m]$ and $12\leq f_{j}\leq k_{j}\leq\ell_{j}-$ 1.
Hence, it follows from ($\Delta_{1}$) that a vector of ${\mathbb{R}}^{|V(G)|}$ of the form:
$\sum\limits_{i=1}^{m}\sum\limits_{j=1}^{2\ell_{i}-1}{a_{i,j}{\mathbf{u}}_{i,j}}+b_{1}{\mathbf{v}}_{1}+b_{2}{\mathbf{v}}_{2}$
belongs to ${\mathbb{R}}_{+}(G)$ if and only if the following formulas are satisfied:
  1. (1)
    $\sum\limits_{j=1}^{m}\sum\limits_{i=1}^{\ell_{j}-1}a_{j,2i}+b_{1}+b_{2}=\sum\limits_{j=1}^{m}\sum\limits_{i=1}^{\ell_{j}}a_{j,2i-1}$;
  2. (2)
    $a_{i,j}\geq 0$ for any $1\leq i\leq m$ and $1\leq j\leq 2\ell_{i}-1;$
  3. (3)
    $b_{1},b_{2}\geq 0;$
  4. (4)
    $\sum\limits_{j=1}^{s}\sum\limits_{i=1}^{k_{p_{j}}}a_{p_{j},2i}+b_{1}\geq\sum\limits_{j=1}^{s}\sum\limits_{i=1}^{k_{p_{j}}}a_{p_{j},2i-1}$ for any $\{p_{1},\ldots,p_{s}\}\subseteq[m]$ and $11\leq k_{p_{j}}\leq\ell_{p_{j}}-$ 1;
  5. (5)
    $\sum\limits_{j=1}^{s}\sum\limits_{i=k_{p_{j}}-1}^{\ell_{p_{j}}-1}a_{p_{j},2i}+b_{2}\geq\sum\limits_{j=1}^{s}\sum\limits_{i=k_{p_{j}}}^{\ell_{p_{j}}}a_{p_{j},2i-1}$ for any $\{p_{1},\ldots,p_{s}\}\subseteq[m]$ and $2\leq k_{p_{j}}\leq\ell_{p_{j}}$;
  6. (6)
    $\sum\limits_{j\neq p}\sum\limits_{i=1}^{\ell_{j}-1}a_{j,2i}+b_{1}+b_{2}\geq\sum\limits_{j\neq p}\sum\limits_{i=1}^{\ell_{j}}a_{j,2i-1}$ for any $p\in[m]$;
  7. (7)
    $\sum\limits_{j\neq p}\sum\limits_{i=1}^{\ell_{j}-1}a_{j,2i}+\sum\limits_{i=1}^{f_{p}}a_{p,2i}+b_{1}+b_{2}\geq\sum\limits_{j\neq p}\sum\limits_{i=1}^{\ell_{j}}a_{j,2i-1}+\sum\limits_{i=1}^{f_{p}}a_{p,2i-1}$ for any $p\in[m]$ and $11\leq f_{p}\leq\ell_{p}-$ 1;
  8. (8)
    $\sum\limits_{j\neq p}\sum\limits_{i=1}^{\ell_{j}-1}a_{j,2i}+\sum\limits_{i=k_{p}}^{\ell_{p}}a_{p,2i-2}+b_{1}+b_{2}\geq\sum\limits_{j\neq p}\sum\limits_{i=1}^{\ell_{j}}a_{j,2i-1}+\sum\limits_{i=k_{p}}^{\ell_{p}}a_{p,2i-1}$ for any $p\in[m]$ and $2\leq k_{p}\leq\ell_{p}$;
  9. (9)
    $\sum\limits_{j\neq p}\sum\limits_{i=1}^{\ell_{j}-1}a_{j,2i}+\sum\limits_{i=1}^{f_{p}}a_{p,2i}+\sum\limits_{i=k_{p}}^{\ell_{p}}a_{p,2i-2}+b_{1}+b_{2}\geq\sum\limits_{j\neq p}\sum\limits_{i=1}^{\ell_{j}}a_{j,2i-1}+\sum\limits_{i=1}^{f_{p}}a_{p,2i-1}+\sum\limits_{i=k_{p}}^{\ell_{p}}a_{p,2i-1}$ for any $p\in[m]$ and $11\leq f_{p}<k_{p}-1\leq\ell_{p}-$ 1;
  10. (10)
    $\sum\limits_{i=f_{j}-1}^{k_{j}}a_{j,2i}\geq\sum\limits_{i=f_{j}}^{k_{j}}a_{j,2i-1}$ for any $j\in[m]$ and $12\leq f_{j}\leq k_{j}\leq\ell_{j}-$ 1.
It is straightforward to check that $\alpha_{\ell}$ satisfies these inequalities, with strict inequalities holding for each $\alpha_{\ell}$. This implies that $\alpha_{\ell}\in\mathrm{relint}({\mathbb{R}}_{+}(G))\cap\mathbb{Z}^{|V(G)|}$.
Next, we show that $\alpha_{\ell}$ is a minimal vector in $\mathrm{relint}({\mathbb{R}}_{+}(G))\cap\mathbb{Z}^{|V(G)|}$, i.e., it cannot be written as a sum of an element in $\mathrm{relint}({\mathbb{R}}_{+}(G))\cap\mathbb{Z}^{|V(G)|}$ and an element in ${\mathbb{R}}_{+}(G)\cap\mathbb{Z}^{|V(G)|}\setminus\{{\bf 0}\}$ for all $\ell=1,\ldots,m$.
Suppose on the contrary that $\alpha_{\ell}=\alpha^{\prime}+\alpha^{\prime\prime}$ for some $\alpha^{\prime}\in\mathrm{relint}({\mathbb{R}}_{+}(G))\cap\mathbb{Z}^{|V(G)|}$ and $\alpha^{\prime\prime}\in{\mathbb{R}}_{+}(G)\cap\mathbb{Z}^{|V(G)|}\setminus\{{\bf 0}\}$. Write
$\alpha^{\prime}=\sum\limits_{i=1}^{m}\sum\limits_{j=1}^{2\ell_{i}-1}{a^{\prime}_{i,j}{\mathbf{u}}_{i,j}}+b^{\prime}_{1}{\mathbf{v}}_{1}+b^{\prime}_{2}{\mathbf{v}}_{2},$
$\alpha^{\prime\prime}=\sum\limits_{i=1}^{m}\sum\limits_{j=1}^{2\ell_{i}-1}{a^{\prime\prime}_{i,j}{\mathbf{u}}_{i,j}}+b^{\prime\prime}_{1}{\mathbf{v}}_{1}+b^{\prime\prime}_{2}{\mathbf{v}}_{2}.$
In view of the inequalities (2) and (3), we see that $a^{\prime}_{i,j},b^{\prime}_{i}\geq 1$ for all $i,j$. Hence, $a^{\prime}_{i,j}=1$ and $a^{\prime\prime}_{i,j}=0$ for all possible $i,j$. Because of (6), we have $b^{\prime}_{1}+b^{\prime}_{2}\geq m$. Hence $b^{\prime\prime}_{1}+b^{\prime\prime}_{2}\leq 0$. From this together with (3) it follows that $b^{\prime\prime}_{1}=b^{\prime\prime}_{2}=0$. Hence, $\alpha^{\prime\prime}=0.$ This is a contradiction, which shows that $\alpha_{\ell}$ is a minimal vector in $\mathrm{relint}({\mathbb{R}}_{+}(G))\cap\mathbb{Z}^{|V(G)|}$ for $\ell=1,\ldots,m-1$. We summarize the above computation in the following lemma.
Lemma 5.6.
Let $G=\mathbf{G}_{2\underline{\ell}}$ be a mutlti-path graph of even type, where ¯$\underline{\ell}=(\ell_{1},\ldots,\ell_{m})\in\mathbb{Z}_{>0}^{m}$. Then $\Theta_{G}v_{1}^{\ell}v_{2}^{m-\ell}$ for $\ell=1,\ldots,m-1$ are minimal generators of $\omega_{\mathbb{K}[G]}$.
Proof.
This is attributed to the existence of a bijection between the minimal vectors in $\mathrm{relint}({\mathbb{R}}_{+}(G))\cap\mathbb{Z}^{|V(G)|}$ and the minimal generators of $\omega_{\mathbb{K}[G]}$.
Actually, $\Theta_{G}v_{1}^{\ell}v_{2}^{m-\ell}$ for $\ell=1,\ldots,m-1$ are all minimal generators of $\omega_{\mathbb{K}[G]}$, as shown lately. For this, we also need the following observation.
Lemma 5.7.
Let $G$ be a connected simple graph with vertex set $V(G)$ and edge set $E(G)$. If $G$ fulfills the odd-cycle condition then
$\mathrm{pdim}(\mathbb{K}[G])=\left\{\begin{array}[]{ll}|E(G)|-|V(G)|+1,&\hbox{$G$ is a bipartite graph;}\\ |E(G)|-|V(G)|,&\hbox{$G$ is not a bipartite graph.}\end{array}\right.$
Proof.
Since $\mathbb{K}[G]$ is Cohen-Macaulay, it follows from [18, Corollary 10.1.21] that
$\mathrm{depth}(\mathbb{K}[G])=\mathrm{dim}(\mathbb{K}[G])=\left\{\begin{array}[]{ll}|V(G)|-1,&\hbox{$G$ is a bipartite graph;}\\ |V(G)|,&\hbox{$G$ is not a bipartite graph.}\end{array}\right.$
Now, the conclusion follows from Auslander-Buchsbaum formula, as stated, for instance, in [9, Corollary A.4.3].
Lemma 5.8.
Denote $H$ as the graph $G$ described above. Then
$\mathrm{pdim}(\mathbb{K}[H])=\mathrm{pdim}(\mathbb{K}[E(G)]/\mathrm{in}_{<}(I_{H}))=m-1.$
Moreover, we have
$\mathcal{N}_{H}=\{\Theta_{H}v_{1}^{\ell}v_{2}^{m-\ell}\mid\ell=1,\ldots,m-1\}.$
Proof.
By virtue of Lemma 5.7, we know that the projective dimension of $\mathbb{K}[H]$ is $m-1$. Furthermore, Lemma 5.5 asserts that the projective dimension of $\mathbb{K}[E(G)]/\mathrm{in}_{<}(I_{H})$ is no greater than $m-1$. Consequently, the first assertion is established in view of Lemma 2.1.
Additionally, since $\mathrm{pdim}(\mathbb{K}[H])=m-1$, it follows from Lemma 5.5 that the top Betti number of $\mathbb{K}[H]$ is bounded above by $m-1$. This implies that the set $\mathcal{N}_{H}$ contains no more than $m-1$ elements. On the other hand, Lemma 5.6 guarantees that $\mathcal{N}_{H}$ contains at least $m-1$ elements. Therefore, the second assertion is achieved by noting $D_{H}=\Theta_{H}^{2}(v_{1}v_{2})^{m}$ in conjunction with Formula (2) in Subsection 2.5.
Proposition 5.9.
Let $G$ be the graph as described above. Then, for any $10\leq i\leq m-$ 1, we have
$\beta_{i}(\mathbb{K}[G])=\beta_{i}(\mathbb{K}[E(G)]/\mathrm{in}_{<}(I_{G})).$
Proof.
Fix $11\leq i\leq m-$ 1. Since $G$ has $m$ even paths, there are $\binom{m}{i+1}$ induced subgraphs of $G$ that contain exactly $i+1$ even paths of $G$. These induced subgraphs are all graphs of the same type as $G$ with a small number of even paths, and they are denoted as $H_{1},\ldots,H_{r}$, respectively. Here, $r=\binom{m}{i+1}$. Note that $\beta_{i,\alpha}(\mathbb{K}[H_{\ell}])=1$ for any $\alpha\in\mathcal{N}_{H_{\ell}}$ by Lemma 5.8, it follows that $\beta_{i,\alpha}(\mathbb{K}[G])=1$ for any $\alpha\in\bigcup_{\ell=1}^{r}\mathcal{N}_{H_{\ell}}$ in view of Lemma 2.7. Set $\mathcal{N}_{i}(G)=\bigcup_{\ell=1}^{r}\mathcal{N}_{H_{\ell}}$.
Note that $\mathrm{supp}(\alpha)=V(H_{k})$ for any $\alpha\in\mathcal{N}_{H_{k}}$, we have $\mathcal{N}_{H_{1}},\ldots,\mathcal{N}_{H_{r}}$ are pairwise distinct. Hence,
$|\mathcal{N}_{i}(G)|=\sum_{\ell=1}^{r}|\mathcal{N}_{H_{\ell}}|=ir=i\binom{m}{i+1}.$
This implies
$i\binom{m}{i+1}=\sum_{\alpha\in\mathcal{N}_{i}(G)}\beta_{i,\alpha}(\mathbb{K}[G])\leq\beta_{i}(\mathbb{K}[G])\leq\beta_{i}(\mathbb{K}[E(G)]/\mathrm{in}_{<}(I_{G}))\leq i\binom{m}{i+1}.$
The last inequality of the above equation comes from Lemma 5.5.
Hence,
$\beta_{i}(\mathbb{K}[G])=\sum_{\alpha\in\mathcal{N}_{i}(G)}\beta_{i,\alpha}(\mathbb{K}[G]).$
In particular, we have if $\beta_{i,\alpha}(\mathbb{K}[G])\neq 0$, then $\alpha\in\mathcal{N}_{i}(G)$ and so $\alpha\in\mathcal{N}_{H_{k}}$ for some $k\in[r]$. Now, the result follows.
Corollary 5.10.
Let $\mathcal{N}_{i}(G)$ denote the set $\sqcup_{H}\mathcal{N}_{H}$, where the union is over all the multi-path induced subgraphs $H$ of $G$ that contain exactly $i+1$ even paths. Then for any monomial $\alpha$ in $\mathbb{K}[G]$, we have
$\beta_{i,\alpha}(\mathbb{K}[G])=\left\{\begin{array}[]{ll}1,&\hbox{$\alpha\in\mathcal{N}_{i}(G)$;}\\ 0,&\hbox{otherwise.}\end{array}\right.$
In particular, $\beta_{i,j}(\mathbb{K}[G])$ is equal to the number of the set
$\{\alpha\in\mathcal{N}_{i}(G)\mid|\mathrm{deg}(\alpha)|=2j\}.$
5.3. Mixed type
Let $H=\mathbf{G}_{\underline{\ell}}$ be a multi-path graph of mixed type, with
¯$\underline{\ell}=(\ell_{1},\ldots,\ell_{m},\ell_{m+1},\ldots,\ell_{m+n}).$
Assume $\ell_{1},\ldots,\ell_{m}$ are all odd numbers $\geq 3$ and $\ell_{m+1},\ldots,\ell_{m+n}$ are all even numbers $\geq 2$. If $m=n=1$, then $\mathbb{K}[H]$ is actually a polynomial ring. Suppose now that either $m\geq 2$ or $n\geq 2$. Let $H_{1},H_{2}$ be induced subgraphs of $H$ such that $H_{i}=\mathbf{G}_{\underline{\ell}_{i}}$ for $i=1,2$, where ¯$\underline{\ell}_{1}=(\ell_{1},\ldots,\ell_{m})$ and ¯$\underline{\ell}_{2}=(\ell_{m+1},\ldots,\ell_{m+n})$. If $m=1$, then $\mathbb{K}[H_{1}]$ is a polynomial ring. It follows that $\mathbb{K}[H]$ and $\mathbb{K}[H_{2}]$ have the same Betti numbers and $\mathcal{N}_{H}=\mathcal{N}_{H_{2}}$. Thus, we may look up $H$ as a multi-path graph of odd type. If $n=1$, then $H$ may be regarded as a multi-path graph of even type. These facts justifies the introduction of multi-path graphs of true-mixed type. Recall that $H$ is of true-mixed type if both $m\geq 2$ and $n\geq 2$. Let us see the following example.
Example 5.11.
Let ¯$\underline{\ell}=(2,2,2,3,3,3)$ and consider the graph $G=\mathbf{G}_{\underline{\ell}}$. Then $G_{1}=\mathbf{G}_{2,2,2}$ and $G_{2}=\mathbf{G}_{3,3,3}$ are both induced subgraphs of $G$. According to Lemmas 5.3 and 5.8, it follows that $\mathcal{N}_{G_{1}}=\Theta_{G_{1}}\{v_{1}v_{2}^{2},v_{1}^{2}v_{2}\}$ and $\mathcal{N}_{G_{2}}=\Theta_{G_{2}}\{v_{1}v_{2},(v_{1}v_{2})^{2}\}$ respectively.
Using Formula (3), we can compute the top multi-graded Betti numbers of $\mathbb{K}[G]$ as follows: $\beta_{4,\alpha_{i}}(\mathbb{K}[G])=1$ for $i=1,\ldots,4$, where $\alpha_{1}=\Theta_{G}v_{1}^{2}v_{2}^{3},\alpha_{2}=\Theta_{G}v_{1}^{3}v_{2}^{2},\alpha_{3}=\Theta_{G}v_{1}^{3}v_{2}^{4}$, and $\alpha_{4}=\Theta_{G}v_{1}^{4}v_{2}^{3}$.
The above example suggests that $\mathcal{N}_{G}=\mathcal{N}_{G_{1}}\bullet\mathcal{N}_{G_{2}}.$ From this, the following result is immediate.
Lemma 5.12.
Let $H$ be a multi-path graph of true-mixed type which consists of $m$ odd-length paths and $n$ even-length paths with $m,n\geq 2$. Let $v_{1}$ and $v_{2}$ be vertices of $H$ with a degree $\geq 3$. Then
$\mathcal{N}_{H}=\{\Theta_{H}v_{1}^{s+t}v_{2}^{n+s-t}\mid s=1,\ldots,m-1,t=1,\ldots,n-1\}.$
In particular, $|\mathcal{N}_{H}|=(m-1)(n-1)$ and $\mathrm{supp}(\alpha)=V(H)$ for all $\alpha\in\mathcal{N}_{H}.$
In the case when $G$ is a multi-path graph of either even type or odd type, the notion $\mathcal{N}_{i}(G)$ has been defined in Proposition 5.4 and Corollary 5.10, respectively. Suppose now that $G$ is a multi-path graph of mixed type. We define $\mathcal{N}_{i}(G)$ as follows. Let $H_{1},\ldots,H_{r}$ denote all the multi-path induced subgraphs of $G$ belonging to true-mixed type that are composed of $i+2$ paths, and let $H_{r+1},\ldots,H_{s}$ denote all the non-mixed multi-path induced subgraphs of $G$ that contain $i+1$ paths. Then, $\mathcal{N}_{i}(G)$ is defined as:
$\mathcal{N}_{i}(G)=\bigcup_{j=1}^{s}\mathcal{N}_{H_{j}}.$
Now, we are ready to present the main result of this section.
Theorem 5.13.
Let $G$ be a multi-path graph that consists of $m$ even-length paths and $n$ odd-length paths. Then
  1. $(1)$
    The projective dimension of $\mathbb{K}[G]$ equals $m+n-2$ if $m$ and $n$ are both $\geq 1$, otherwise it is $m+n-1$;
  2. $(2)$
    For any integer $1\leq i\leq\mathrm{pdim}(\mathbb{K}[G])$, the $i$-th multi-graded Betti numbers of $\mathbb{K}[G]$ are given by
    $\beta_{i,\alpha}(\mathbb{K}[G])=\left\{\begin{array}[]{ll}1,&\hbox{$\alpha\in\mathcal{N}_{i}(G)$;}\\ 0,&\hbox{otherwise.}\end{array}\right.$
Proof.
(1) It is clear from Lemma 5.7.
(2) The cases in which either $m=0$ or $n=0$ are direct consequences of Proposition 5.4 and Corollary 5.10, respectively. Therefore, we may assume without loss of generality that $m\geq 1$ and $n\geq 1$. Under these assumptions, the $i$-th Betti number of $\mathbb{K}[G]$ is given by
$\beta_{i}(\mathbb{K}[G])=\sum_{\begin{subarray}{c}j+k=i\\ j\geq 1,k\geq 1\end{subarray}}\binom{n}{j+1}\binom{m}{k+1}jk+i(\binom{m}{i+1}+\binom{n}{i+1}).$
On the other hand, according to the definition of $\mathcal{N}_{i}(G)$, the Betti number $\beta_{i,\alpha}(\mathbb{K}[G])$ is equal to 1 for every $\alpha$ belonging to $\mathcal{N}_{i}(G)$, hence, to complete the proof, it suffices to demonstrate that the cardinality of $\mathcal{N}_{i}(G)$ is equal to the $i$-th Betti number of $\mathbb{K}[G]$. Let $H_{1},\ldots,H_{r},H_{r+1},\ldots,H_{s}$ be as in the definition of $\mathcal{N}_{i}(G)$. Note that all these subsets $\mathcal{N}_{H_{i}}$ for $i=1,\ldots,s$ are pairwise disjoint. Fix integers $j$ and $k$ such that $j\geq 2$, $k\geq 2$, and $j+k=i+2$, there exist exactly $\binom{m}{j}\binom{n}{k}$ elements $t$ in the set $[r]$ such that $H_{t}$ contains $j$ even-length paths and $k$ odd-length paths, furthermore, for such $t\in[r]$, the cardinality of $\mathcal{N}_{H_{t}}$ is $(j-1)(k-1)$. This implies
$|\bigcup_{t=1}^{r}\mathcal{N}_{H_{j}}|=\sum_{\begin{subarray}{c}j+k=i+2\\ j\geq 2,k\geq 2\end{subarray}}\binom{m}{j}\binom{n}{k}(j-1)(k-1).$
Next, it is easy to see that $|\mathcal{N}_{H_{t}}|=i$ for $t=r+1,\ldots,s$ by Lemmas 5.3 as well as 5.8. Moreover, we have $s-r=\binom{m}{i+1}+\binom{n}{i+1}$. It follows that
$|\bigcup_{t=r+1}^{s}\mathcal{N}_{H_{j}}|=i(\binom{m}{i+1}+\binom{n}{i+1}).$
Therefore, we have
$|\mathcal{N}_{i}(G)|=\sum_{\begin{subarray}{c}j+k=i+2\\ j\geq 2,k\geq 2\end{subarray}}\binom{m}{j}\binom{n}{k}(j-1)(k-1)+i(\binom{m}{i+1}+\binom{n}{i+1})=\beta_{i}(\mathbb{K}[G]).$
This completes the proof.
6. Betti numbers of $\mathbf{G}_{m}$
In [10], a specific class of simple graphs designated as $\mathbf{G}_{m}$ was introduced. We call such graphs two-ear graphs. In this section, we offer a precise formula for computing the multi-graded Betti numbers of $\mathbb{K}[\mathbf{G}_{m}]$, further elucidating the algebraic properties of this graph class.
Definition 6.1.
Let $m\geq 1$ be an integer. Following [10], we define $\mathbf{G}_{m}$ as a simple graph with vertex set
$V(\mathbf{G}_{m})=\{v_{1},v_{2},x_{1},x_{2},y_{1},y_{2}\}\cup\{u_{i}~{}|~{}1\leq i\leq m-1\}$
and with edge set
[…]
$\quad\cup\{\{v_{1},u_{i}~{}|~{}1\leq i\leq m-1\}\}\cup\{\{v_{2},u_{i}~{}|~{}1\leq i\leq m-1\}\}.$
Let $\mathbf{P}_{m}$ denote the induced subgraph $\mathbf{G}_{m}\setminus x_{2}$ of $\mathbf{G}_{m}$. We call a graph $G$ a two-ear graph of type $m$ or an one-ear graph of type $m$ if it is isomorphic to the graph $\mathbf{G}_{m}$ or $\mathbf{P}_{m}$, respectively.
By Lemma 5.7, the following result is immediate.
Lemma 6.2.
For all $m\geq 1$, we have $\mathrm{pdim}(\mathbb{K}[\mathbf{G}_{m}])=m$ and $\mathrm{pdim}(\mathbb{K}[\mathbf{P}_{m}])=m-1$.
$v_{1}$$v_{2}$$x_{1}$$u_{m-1}$$y_{1}$$u_{2}$$u_{1}$$x_{2}$$y_{2}$$e_{m-1,1}$$e_{m-1,2}$$e_{2,1}$$e_{2,2}$$e_{1,1}$$e_{1,2}$$f_{1}$$g_{1}$$f_{2}$$g_{2}$$f_{3}$$g_{3}$$w$
Figure 2. The two-ear graph $\mathbf{G}_{m}$
Let us denote $\mathbf{G}_{m}$ as $G$. To write out the minimal generators of $\mathrm{in}_{<}(I_{G})$, we label the edges of $G$ as follows: for $i\in[m-1]$, $e_{i,1}=\{v_{1},u_{i}\},e_{i,2}=\{v_{2},u_{i}\}$, and $f_{1}=\{v_{1},x_{1}\},f_{2}=\{x_{1},x_{2}\},f_{3}=\{v_{1},x_{2}\},g_{1}=\{v_{2},y_{1}\},g_{2}=\{y_{1},y_{2}\},g_{3}=\{v_{2},y_{2}\},w=\{x_{1},y_{1}\}$. In [10, Corollary 4.3 ], the minimal set of monomial generators of $\mathrm{in}_{<}(I_{G})$ is specified for a particular monomial order $<$. Note that the monomial order $<$ in this section refers to this order. We record this result in the following lemma for later use.
Lemma 6.3.
There is a monomial order $<$ such that the minimal set of monomial generators of $\mathrm{in}_{<}(I_{G})$ is given by $\mathcal{M}=\mathcal{M}_{1}\cup\mathcal{M}_{2}\cup\mathcal{M}_{3}\cup\mathcal{M}_{4}$, where
  1. $(i)$
[…]
  2. $(ii)$
    $\mathcal{M}_{2}=\left\{f_{1}f_{2}g_{1}g_{2}\right\}$,
  3. $(iii)$
[…]
  4. $(iv)$
[…]
We sort the set of minimal generators of $\mathrm{in}_{<}(I_{G})$ as follows:
$\begin{split}&e_{1,1}e_{2,2}\prec e_{1,1}e_{2,3}\prec\cdots\prec e_{1,1}e_{m-1,2}\prec e_{2,1}e_{3,2}\prec e_{2,1}e_{4,2}\prec\cdots\prec e_{2,1}e_{m-1,2}\prec\cdots\\ &\prec e_{m-2,1}e_{m-1,2}\prec g_{1}f_{2}e_{1,1}\prec g_{1}f_{2}e_{2,1}\prec\cdots\prec g_{1}f_{2}e_{m-1,1}\prec f_{1}f_{2}g_{1}g_{2}\prec g_{2}f_{1}e_{m-1,2}\\ &\prec g_{2}f_{1}e_{m-2,2}\prec\cdots\prec g_{2}f_{1}e_{1,2}.\end{split}$
For brevity, we label each term in this sequence consecutively: $e_{1,1}e_{2,2}$ as $u_{1}$, $e_{1,1}e_{2,3}$ as $u_{2}$, and so forth, culminating with $g_{2}f_{1}e_{1,2}$ designated as $u_{r}$. Here, $r=\frac{m(m+1)}{2}$.
Lemma 6.4.
The monomial ideal $I=\mathrm{in}_{<}(I_{G})$ has regular quotients with respect to the above order. More precisely, the colon ideal $(u_{1},\ldots,u_{j-1}):u_{j}$ is generated by a regular sequence of length $k_{j}$ for $j=2,\ldots,\frac{m(m+1)}{2}$, where
$(k_{2},\ldots,k_{\frac{m(m+1)}{2}})=(\underbrace{{1,1}}_{2},\underbrace{2,2,2}_{3},\ldots,\underbrace{m-1,m-1,\ldots,m-1}_{m}).$
Proof.
Let $I_{j-1}=(u_{1},\ldots,u_{j-1})$ be the ideal generated by the first $j-1$ elements of the sequence. Now, we consider several cases based on the form of $u_{j}$:
1. If $u_{j}=e_{a,1}e_{b,2}$ for some $11\leq a<b\leq m-$ 1, then the colon ideal $I_{j-1}:u_{j}$ is given by
$I_{j-1}:u_{j}=(e_{1,1},e_{2,1},\ldots,e_{a-2,1},e_{a,2},e_{a+1,2},\ldots,e_{b-2,2}).$
2. If $u_{j}=g_{1}f_{2}e_{a,1}$ for some $11\leq a\leq m-$ 1, the colon ideal becomes
$I_{j-1}:u_{j}=(e_{1,1},e_{2,1},\ldots,e_{a-2,1},e_{a,2},e_{a+1,2},\ldots,e_{m-1,2}).$
3. If $u_{j}=f_{1}f_{2}g_{1}g_{2}$, the colon ideal simplifies to
$I_{j-1}:u_{j}=(e_{1,1},e_{2,1},\ldots,e_{m-1,1}).$
4. Finally, if $u_{j}=g_{2}f_{1}e_{a,2}$ for some $11\leq a\leq m-$ 1, the colon ideal is
$I_{j-1}:u_{j}=(g_{1}f_{2},e_{1,1},e_{2,1},\ldots,e_{a-2,1},e_{a,2},e_{a+1,2},\ldots,e_{m-1,2}).$
By enumerating the minimal generators of these colon ideals, the desired result follows.
Lemma 6.5.
Let $G$ be the graph as described above. Then there is a monomial order $<$ such that $\beta_{i}(\mathrm{in}_{<}(I_{G}))\leq(i+1)\binom{m+1}{i+2}$ for all $i\geq 0$. Thus,
$\beta_{i}(\mathbb{K}[G])\leq\beta_{i}(\mathbb{K}[E(G)]/\mathrm{in}_{<}(I_{G}))\leq i\binom{m+1}{i+1}.$
Proof.
In view of Lemma 6.4 the result follows by the same argument as in Lemma 5.5.
Next, we compute the minimal generators of the canonical module $\omega_{\mathbb{K}[G]}$. It is easy to see that $|V(G)|=m+5$. It follows from Conventions 2.6 that:
${\mathbb{R}}^{|V(G)|}=\{\sum\limits_{i=1}^{m-1}{a_{i}{\mathbf{u}}_{i}}+\sum\limits_{i=1}^{2}({b_{i}{\mathbf{v}}_{i}+c_{i}{\mathbf{x}}_{i}+d_{i}{\mathbf{y}}_{i}})\mid a_{i},b_{i},c_{i},d_{i}\in{\mathbb{R}}\mbox{ for all }i\},$
where ${\mathbf{u}}_{i},{\mathbf{v}}_{i},{\mathbf{x}}_{i},{\mathbf{y}}_{i}$ are the unit vectors of ${\mathbb{R}}^{|V(G)|}$, each ${\mathbf{u}}_{i}$ (resp. ${\mathbf{v}}_{i},{\mathbf{x}}_{i},{\mathbf{y}}_{i}$) corresponds to $u_{i}$ (resp. $v_{i},x_{i},y_{i}$) for all $i$.
In what follows, we will construct $m$ integral vectors in ${\mathbb{R}}^{|V(G)|}$ and then show that they are minimal vectors of $\mathrm{relint}({\mathbb{R}}_{+}(G))\cap\mathbb{Z}^{|V(G)|}$. The construction is as follows:
For $\ell=1,\ldots,m$, let
$\alpha_{\ell}:=\sum\limits_{i=1}^{m-1}{\mathbf{u}}_{i}+\sum\limits_{i=1}^{2}({\mathbf{x}}_{i}+{\mathbf{y}}_{i})+\ell{\mathbf{v}}_{1}+(m+1-\ell){\mathbf{v}}_{2}.$
We now verify that $\alpha_{\ell}\in\mathrm{relint}({\mathbb{R}}_{+}(G))$ for all $\ell=1,\ldots,m$.
We see the following:
  • The vertices of $G$ are all regular vertices.
  • An independent subset $T$ of $V(G)$ is fundamental if and only if $T$ is one of the following sets:
    1. $(i)$
      $\{u_{1},u_{2},\cdots,u_{m-1},x_{1},y_{2}\};$
    2. $(ii)$
      $\{u_{1},u_{2},\cdots,u_{m-1},x_{2},y_{1}\};$
    3. $(iii)$
      $\{v_{1},v_{2}\};$
    4. $(iv)$
      $\{v_{1},y_{1}\};$
    5. $(v)$
      $\{v_{2},x_{1}\}.$
Hence, it follows from ($\Delta_{2}$) that a vector of ${\mathbb{R}}^{|V(G)|}$ of the form:
$\sum\limits_{i=1}^{m-1}{a_{i}{\mathbf{u}}_{i}}+\sum\limits_{i=1}^{2}({b_{i}{\mathbf{v}}_{i}+c_{i}{\mathbf{x}}_{i}+d_{i}{\mathbf{y}}_{i}})$
belongs to ${\mathbb{R}}_{+}(G)$ if and only if the following inequalities are satisfied:
  1. (1)
    $a_{i}\geq 0$ for any $11\leq i\leq m-$ 1 ;
  2. (2)
    $b_{i},c_{i},d_{i}\geq 0$ for any $1\leq i\leq 2$ ;
  3. (3)
    $b_{1}+b_{2}+c_{2}+d_{1}\geq\sum\limits_{i=1}^{m-1}a_{i}+c_{1}+d_{2}$ ;
  4. (4)
    $b_{1}+b_{2}+c_{1}+d_{2}\geq\sum\limits_{i=1}^{m-1}a_{i}+c_{2}+d_{1}$ ;
  5. (5)
    $\sum\limits_{i=1}^{m-1}a_{i}+c_{1}+c_{2}+d_{1}+d_{2}\geq b_{1}+b_{2}$ ;
  6. (6)
    $\sum\limits_{i=1}^{m-1}a_{i}+b_{2}+c_{1}+c_{2}+d_{2}\geq b_{1}+d_{1}$ ;
  7. (7)
    $\sum\limits_{i=1}^{m-1}a_{i}+b_{1}+c_{2}+d_{1}+d_{2}\geq b_{2}+c_{1}$
It is straightforward to check that $\alpha_{\ell}$ satisfies these inequalities, with strict inequalities holding for each $\alpha_{\ell}$. This implies that $\alpha_{\ell}\in\mathrm{relint}({\mathbb{R}}_{+}(G))\cap\mathbb{Z}^{|V(G)|}$.
Next, we show that $\alpha_{\ell}$ is a minimal vector in $\mathrm{relint}({\mathbb{R}}_{+}(G))\cap\mathbb{Z}^{|V(G)|}$.
Suppose on the contrary that $\alpha_{\ell}=\alpha^{\prime}+\alpha^{\prime\prime}$ for some $\alpha^{\prime}\in\mathrm{relint}({\mathbb{R}}_{+}(G))\cap\mathbb{Z}^{|V(G)|}$ and $\alpha^{\prime\prime}\in{\mathbb{R}}_{+}(G)\cap\mathbb{Z}^{|V(G)|}\setminus\{{\bf 0}\}$. Write
$\alpha^{\prime}=\sum\limits_{i=1}^{m-1}{a^{\prime}_{i}{\mathbf{u}}_{i}}+\sum\limits_{i=1}^{2}({b^{\prime}_{i}{\mathbf{v}}_{i}+c^{\prime}_{i}{\mathbf{x}}_{i}+d^{\prime}_{i}{\mathbf{y}}_{i}}),$
$\alpha^{\prime\prime}=\sum\limits_{i=1}^{m-1}{a^{\prime\prime}_{i}{\mathbf{u}}_{i}}+\sum\limits_{i=1}^{2}({b^{\prime\prime}_{i}{\mathbf{v}}_{i}+c^{\prime\prime}_{i}{\mathbf{x}}_{i}+d^{\prime\prime}_{i}{\mathbf{y}}_{i}}).$
In view of the inequalities (1) and (2), we see that $a^{\prime}_{i},c^{\prime}_{i},d^{\prime}_{i}\geq 1$ and $a^{\prime\prime}_{i},c^{\prime\prime}_{i},d^{\prime\prime}_{i}\geq 0$ for all $i,j$. Hence, $a^{\prime}_{i}=c^{\prime}_{i}=d^{\prime}_{i}=1$ and $a^{\prime\prime}_{i}=c^{\prime\prime}_{i}=d^{\prime\prime}_{i}=0$ for all possible $i,j$. Because of (3) ,we have $b^{\prime}_{1}+b^{\prime}_{2}\geq m$. Hence $b^{\prime\prime}_{1}+b^{\prime\prime}_{2}\leq 1$. Because of (6) and (7), we have $b^{\prime\prime}_{1}=b^{\prime\prime}_{2}$. But from (2) we know $b^{\prime\prime}_{1},b^{\prime\prime}_{2}\geq 0$, so we have $b^{\prime\prime}_{1}=b^{\prime\prime}_{2}=0$.
Hence, $\alpha^{\prime\prime}=0.$ This is a contradiction, which shows that $\alpha_{\ell}$ is a minimal vector in $\mathrm{relint}({\mathbb{R}}_{+}(G))\cap\mathbb{Z}^{|V(G)|}$ for $\ell=1,\ldots,m$.
Actually, $\alpha_{\ell}$ for $\ell=1,\ldots,m$ are all minimal vector in $\mathrm{relint}({\mathbb{R}}_{+}(G))\cap\mathbb{Z}^{|V(G)|}$, as shown below.
Lemma 6.6.
Denote by $H$ the graph $G$ described above. Then
$\mathrm{pdim}(\mathbb{K}[H])=\mathrm{pdim}(\mathbb{K}[E(H)]/\mathrm{in}_{<}(I_{H}))=m,$
and moreover,
$\mathcal{N}_{H}=\{\Theta_{H}x_{1}y_{1}v_{1}^{\ell}v_{2}^{m+1-\ell}\mid\ell=1,\ldots,m\}.$
Proof.
By Lemma 6.2, we have $\mathrm{pdim}(\mathbb{K}[H])=m$. Furthermore, according to Lemma 6.5, it can be inferred that $\mathrm{pdim}(\mathbb{K}[E(H)]/\mathrm{in}_{<}(I_{H}))\leq m$. Combined with Lemma 2.1, the first assertion can be proved.
Moreover, given that $\mathrm{pdim}(\mathbb{K}[H])=m$, Lemma 6.5 implies that the top Betti number of $\mathbb{K}[H]$ is at most $m$. It follows that $\mathcal{N}_{H}$ contains no more than $m$ elements. On the other hand, as demonstrated earlier, $\mathcal{N}_{H}$ contains at least $m$ elements. The desired result then follows by observing that $D_{H}=\Theta_{H}^{2}(x_{1}y_{1})^{3}(v_{1}v_{2})^{m+1}$ together with Formula (2) in Subsection 2.5.
Lemma 6.6 enables us to express $\mathcal{N}_{H}$ for all graphs $H$ which are isomorphic to $\mathbf{G}_{m}$.
Let $G^{\prime}$ (resp. $G^{\prime\prime}$) denote the induced subgraph $G\setminus x_{2}$ (resp. $G\setminus y_{2}$) of $G$. According to Definition 6.1, both $G^{\prime}$ and $G^{\prime\prime}$ are isomorphic to $\mathbf{P}_{m},$ which are one-ear graphs of type $m$. We need to identify the elements in the set $\mathcal{N}_{H}$ for an one-ear graph $H$.
It follows from Conventions 2.6 that:
${\mathbb{R}}^{|V(G^{\prime})|}=\{\sum\limits_{i=1}^{m-1}{a_{i}{\mathbf{u}}_{i}}+\sum\limits_{i=1}^{2}({b_{i}{\mathbf{v}}_{i}+d_{i}{\mathbf{y}}_{i}})+c_{1}{\mathbf{x}}_{1}\mid a_{i},b_{i},c_{1},d_{i}\in{\mathbb{R}}\mbox{for all }i\},$
here, ${\mathbf{u}}_{i},{\mathbf{v}}_{i},{\mathbf{x}}_{1},{\mathbf{y}}_{i}$ correspond the vertices of $G^{\prime}$ in a natural way. We could show the following vectors employing the same process as in the case of a two-ear graph:
$\alpha_{\ell}:=\sum\limits_{i=1}^{m-1}{\mathbf{u}}_{i}+{\mathbf{x}}_{1}+2{\mathbf{y}}_{1}+{\mathbf{y}}_{2}+\ell{\mathbf{v}}_{1}+(m-\ell){\mathbf{v}}_{2},\ \ell=1,\ldots,m-1$
are the minimal vectors of $\mathrm{relint}({\mathbb{R}}_{+}(G^{\prime}))\cap{\mathbb{Z}}^{|V(G^{\prime})|}$.
It is clear that $D_{H}=\Theta_{H}^{2}y_{1}^{3}v_{1}^{m}v_{2}^{m+1}$. Therefore, the subsequent result follows directly from Formula 2 in Subsection 2.5.
Lemma 6.7.
Denote $H=\mathbf{P}_{m}$ as graph $G^{\prime}$ described above, then
$\{\Theta_{H}y_{1}^{2}v_{1}^{\ell}v_{2}^{m-\ell}\mid\ell=1,\ldots,m-1\}\subseteq\mathcal{N}_{H}.$
Complete bipartite graphs of the form $\mathbf{K}_{2,d}$ for some $d\geq 2$ may potentially be induced subgraphs of $G$. Observe that $\mathbf{K}_{2,d}$ is isomorphic to $\mathbf{G}_{\underline{\ell}}$, where $\underline{\ell}$ consists of $d$ occurrences of $2$. Consequently, if $H$ is isomorphic to $\mathbf{K}_{2,d}$, then $\mathrm{pdim}(\mathbb{K}[H])=d-1$ and the set $\mathcal{N}_{H}$ comprises $d-1$ monomials. Furthermore, each monomial in $\mathcal{N}_{H}$ has a support that coincides with the vertex set $V(H)$.
Proposition 6.8.
Let $G$ be the graph as described above. Then, for any $0\leq i\leq m$, we have
$\beta_{i}(\mathbb{K}[G])=\beta_{i}(\mathbb{K}[E(G)]/\mathrm{in}_{<}(I_{G})).$
Proof.
From Lemma 6.6, it can be seen that $\mathrm{pdim}(\mathbb{K}[G])=\mathrm{pdim}(\mathbb{K}[E(G)]/\mathrm{in}_{<}(I_{G}))=m$.
If $i=0$, then $\beta_{i}(\mathbb{K}[G])=\beta_{i}(\mathbb{K}[E(G)]/\mathrm{in}_{<}(I_{G}))=1.$ Suppose now that $1\leq i\leq m$. We now identify all induced subgraphs of $G$ whose edge rings have a projective dimension of $i$. Note that $\mathrm{pdim}(\mathbb{K}[\mathbf{G}_{i}])=\mathrm{pdim}(\mathbb{K}[\mathbf{P}_{i+1}])=\mathrm{pdim}(\mathbb{K}[\mathbf{K}_{2,i+1}])$. Given that $G$ has $\binom{m-1}{i-1}$ induced subgraphs that are isomorphic to $\mathbf{G}_{i}$, 2$\binom{m-1}{i}$ induced subgraphs that are isomorphic to $\mathbf{P}_{i+1}$, and $\binom{m-1}{i+1}$ induced subgraphs which are isomorphic to $\mathbf{K}_{2,i+1}$, there are at least $r$ induced subgraphs $H$ of $G$ such that $\mathrm{pdim}(\mathbb{K}[H])=i$. Here, $r$ is given by
$r=\binom{m-1}{i-1}+2\binom{m-1}{i}+\binom{m-1}{i+1}=\binom{m+1}{i+1}.$
Let $\mathcal{N}_{i}(G)$ denote the union of all $\mathcal{N}_{H}$ such that $H$ is an induced subgraph of $G$ satisfying $\mathrm{pdim}(\mathbb{K}[H])=i$. Observe that if $H$ is isomorphic to either $\mathbf{G}_{i}$ or $\mathbf{K}_{2,i+1}$, then, according to Lemma 6.6 and the preceding statement of this proposition, $\mathcal{N}_{H}$ comprises $i$ monomials, with each of their supports aligning with $V(H)$. While if $H$ is isomorphic to $\mathbf{P}_{i+1}$, then Lemma 6.7 ensures that $\mathcal{N}_{H}$ contains at least $i$ monomials, with their supports coinciding with $V(H)$ as well. Hence,
$|\mathcal{N}_{i}(G)|\geq ir=i\binom{m+1}{i+1}.$
This implies
$i\binom{m+1}{i+1}\leq\sum_{\alpha\in\mathcal{N}_{i}(G)}\beta_{i,\alpha}(\mathbb{K}[G])\leq\beta_{i}(\mathbb{K}[G])\leq\beta_{i}(\mathbb{K}[E(G)]/\mathrm{in}_{<}(I_{G}))\leq i\binom{m+1}{i+1},$
as desired.
From the proof of Proposition 6.8, we obtain the following corollaries immediately. The first is an improvement of Lemma 6.7, which allows us to write all elements of $\mathcal{N}_{H}$ for all one-ear graphs $H$.
Corollary 6.9.
Denote by $H$ the graph $G^{\prime}$ described above, then
$\mathcal{N}_{H}=\{\Theta_{H}y_{1}^{2}v_{1}^{\ell}v_{2}^{m-\ell}\mid\ell=1,\ldots,m-1\}.$
The second is the main result of this section. As in the proof of Proposition 6.8, $\mathcal{N}_{i}(G)$ denotes the union of all $\mathcal{N}_{H}$ such that $H$ is an induced subgraph of $G$ satisfying $\mathrm{pdim}(\mathbb{K}[H])=i$.
Theorem 6.10.
Let $G$ be a two-ear graph of type $m$. Then for any monomial $\alpha$ in $\mathbb{K}[G]$ and for any integer $i\geq 1$, we have
$\beta_{i,\alpha}(\mathbb{K}[G])=\left\{\begin{array}[]{ll}1,&\hbox{$\alpha\in\mathcal{N}_{i}(G)$;}\\ 0,&\hbox{otherwise.}\end{array}\right.$
In particular, $\beta_{i,j}(\mathbb{K}[G])$ is equal to the number of the set $\{\alpha\in\mathcal{N}_{i}(G)\mid|\mathrm{deg}(\alpha)|=2j\}.$
7. Closing Remarks
The three classes of graphs we study are representative, each exhibiting distinct characteristics. The compact graphs are devoid of even cycles. Multi-path graphs, on the other hand, lack odd cycles in the sense that no primitive even closed walks traverse an odd cycle. Lastly, two-ear graphs embrace both even and odd cycles. reflecting a different flavor.
Remark 7.1.
In the three theorems computing multi-graded Betti numbers of edge rings, the concept $\mathcal{N}_{i}(G)$ is consistently employed. In essence, the definitions of $\mathcal{N}_{i}(G)$ are identical across these theorems. Namely, $\mathcal{N}_{i}(G)$ is always the union of all sets $\mathcal{N}_{H}$, where $H$ is an induced subgraph of $G$ such that $\mathrm{pdim}(\mathbb{K}[H])=i$. However, when calculating $\mathcal{N}_{i}(G)$, in order to avoid duplication, we do not consider all induced subgraphs of $G$ satisfying the above condition. For example, if $G$ is a multi-path graph of mixed type, we do not need to consider the induced subgraphs of $G$ that contain exactly one even-length path and $i+1$ odd-length paths in the process of obtaining $\mathcal{N}_{i}(G)$.
Remark 7.2.
Let $G$ be a simple graph satisfying the odd-condition. The prerequisite for our method to hold is to find an appropriate monomial order $<$ such that $\mathrm{in}_{<}(I_{G})$ and $I_{G}$ have the same Betti numbers. In particular, we have the following result.
Proposition 7.3.
Let $G$ be either a compact graph, a multi-path graph, or a two-ear graph. There is a monomial order $<$ (i.e., the monomial orders considered in this paper) such that
$\beta_{i,j}(\mathrm{in}_{<}(I_{G}))=\beta_{i.j}(I_{G})\mbox{ for all }i,j.$
In particular, for such monomial orders, $\mathrm{in}_{<}(I_{G})$ is a Cohen-Macaulay ideal.
Remark 7.4.
Let $G$ be a graph considered in this paper. In most cases, every non-zero multi-graded Betti number of $\mathbb{K}[G]$ is equal to a top multi-graded Betti number of the edge ring of some induced subgraph of $G$.
Proposition 7.5.
Let $G$ be either a multi-path graph, a two-ear graph, or a compact graph of either type one or type two. Consider a monomial $\alpha\in\mathbb{K}[G]$ and let $H$ be the induced graph of $G$ on $\mathrm{supp}(\alpha)$. Then the following statements are equivalent:
  1. $(1)$
    $\beta_{i,\alpha}(\mathbb{K}[G])\neq 0$;
  2. $(2)$
    $\beta_{i,\alpha}(\mathbb{K}[G])=1$;
  3. $(3)$
    $\mathrm{pdim}(\mathbb{K}[H])=i$ and $\beta_{i,\alpha}(\mathbb{K}[H])=1$.
It is noteworthy to mention that if $G$ is a compact graph of type three, the previously stated result ceases to be valid. Firstly, the Betti number $\beta_{i,\alpha}(\mathbb{K}[G])$ may take a value of 2 in this case. Secondly, a non-zero multi-graded Betti number of $\mathbb{K}[G]$ could be a second top Betti number of an induced subgraph of $G$.
Remark 7.6.
In most cases, the projective dimension of $\mathbb{K}[G]$ coincides with the type of $\mathbb{K}[G]$, which, by definition, corresponds to the number of $\mathcal{N}_{G}$.
Proposition 7.7.
Let $G$ be either a compact graph, multi-path graphs of even or odd type, or a two-ear graph. Then $\mathrm{pdim}(\mathbb{K}[G])$ is equal to the type of $\mathbb{K}[G].$
However, for multi-path graphs of mixed type, this equivalence does not generally hold: the type may be much larger than the projective dimension.
Remark 7.8.
Notably, the three classes of graphs we investigate are planar, presenting a natural starting point for our exploration. As we move forward, the next challenge lies in tackling with non-planar graphs. One of such graphs are complete bipartite graphs $K_{m,n}$ with $m,n\geq 3$. For such graphs, the main obstacle is to obtain a suitable upper bound for the total Betti numbers of their edge rings.
Remark 7.9.
In the quest to derive upper bounds for $\mathbb{K}[G]$ when $G$ is a multi-path graph or a two-ear graph, the finding that there is a monomial order $<$ such that $\mathrm{in}_{<}(I_{G})$ has regular quotients with respect to some order of its generators plays a key role. Indeed, the same discovery remains valid even when G is a compact graph. These facts lead us ask the following question
Question: For any normal edge ring $\mathbb{K}[G]$, does there always exist a monomial order $<$ such that $\mathrm{in}_{<}(I_{G})$ has regular quotients with respect to some order of its generators?
Acknowledgment: This project is supported by NSFC (No. 11971338) The authors are grateful to the software systems [4] for providing us with a large number of examples to test our ideas.