### v2

# 111* Corresponding author

Abstract.
Motivated by the study of square-free Gröbner degenerations conducted by Conca and Varbaro in their 2020 paper [8], we embark on an investigation of Betti numbers of toric ideals associated with normal edge rings by introducing a novel method. We utilize this method to prove that, for two classes of graphs satisfying the odd-cycle condition, all the Betti numbers (not merely the extremal ones) of the toric ideals of their edge rings, as well as those of certain square-free initial ideals derived from these toric ideals, are identical. As a result, we can completely determine their multi-graded Betti numbers. These classes consist of compact graphs, i.e., the graphs devoid of even cycles and satisfying the odd-cycle condition, and the graphs—as introduced in [14]—that possess both even and odd cycles. We speculate this method can be applicable to every normal edge ring.
Key words and phrases:
Induced-subgraph method, Multi-graded Betti number, Top-Betti graph, Compact graph, Canonical module, Square-free Gröbner degeneration
2010 Mathematics Subject Classification:
Primary 13A02, 05E40 Secondary 06D50.
1. Introduction
Let $R:=\mathbb{K}[X_{1},\ldots,X_{n}]$ be a standard graded polynomial ring over a field $\mathbb{K}$. In the realms of algebraic geometry and commutative algebra, one of the classical and well-studied problems is the investigation of the graded Betti numbers of graded ideals in the ring $R$. When considering a square-free monomial ideal of $R$, Hochster, in the reference [12], elegantly presented a formula. This formula establishes a connection between the multi-graded Betti numbers of the ideal and the simplicial homologies of relevant simplicial complexes. In more recent times, the scope of this formula has been broadened. Specifically, it has been extended to accommodate an arbitrary monomial ideal, as detailed in [17, Theorem 1.34].
[…]
Though much progress has been made on the natural problem, we are still far from obtaining a complete answer. To shed more light on this problem, we pose the following task.
Task: Find more examples of graded ideals $I$ for which there exists a term order $<$ such that $\mathrm{in}_{<}(I)$ is square-free and, simultaneously, the equality $\beta_{i,j}(R/I)=\beta_{i,j}(R/\mathrm{in}_{<}(I))$ holds for all $i,j$.
[…]
Let $G=(V(G),E(G))$ be a simple graph with vertex set $V(G)=\{x_{1},\ldots,x_{n}\}$ and edge set $E(G)=\{e_{1},\ldots,e_{m}\}$ and let $\mathbb{K}$ be a field. The edge ring $\mathbb{K}[G]$ is the toric ring $\mathbb{K}[x_{e}\mid e\in E(G)]\subseteq\mathbb{K}[x_{1},\ldots,x_{n}]$. Here, $x_{e}=\prod_{x_{i}\in e}x_{i}$ for all $e\in E(G)$. Let $\mathbb{K}[e_{1},\ldots,e_{m}]$ or $\mathbb{K}[E(G)]$ denote the polynomial ring over $\mathbb{K}$ in variables $e_{1},\ldots,e_{m}$. We can define a ring homomorphism $\phi:\mathbb{K}[E(G)]\rightarrow\mathbb{K}[G]$ by letting $\phi(e_{i})=x_{e_{i}}$ for all $i$. The toric ideal of $\mathbb{K}[G]$ or $G$ is the ideal $I_{G}=ker(\phi)$. It follows that $\mathbb{K}[G]\cong\mathbb{K}[E(G)]/I_{G}$. The study of Betti numbers and other algebraic properties of toric ideals of graphs is an active area of research. These studies emphasize the use of combinatorial properties of graphs to describe the algebraic properties of toric ideals of graphs. Relevant research can be found in [6], [9], [10], [11], and [18] and the references therein.
Influenced jointly by the research $\ddot{\text{o}}$bner degeneration in [8] as well as the remarkable result in [22, Proposition 13.15] that if a toric ideal has a square-free initial ideal, then its quotient ring is a normal ring, we consider how to compute the Betti numbers of normal edge rings. In 1999, Ohsugi and Hibi established in [19] that the edge ring $\mathbb{K}[G]$ is a normal algebra precisely when the graph $G$ fulfills the odd-cycle condition. By definition, a simple graph satisfies the odd-cycle condition whenever any two minimal cycles $C_{1}$ and $C_{2}$ either share at least one vertex or are connected by an edge bridging a vertex from $C_{1}$ to a vertex from $C_{2}$. In this paper, we introduce a novel approach for computing the multi-graded Betti numbers of normal edge rings. This method not only expresses the Betti numbers explicitly but also reveals the connection between the combinatorial properties of the graph and the Betti numbers of its edge ring.
We first provide some definitions and notions to facilitate the introduction of our approach and the main results of this paper. Assuming the normality of $\mathbb{K}[G]$, [3, Theorem 6.31] reveals that the multi-graded canonical module of the multi-graded $\mathbb{K}[E(G)]$-algebra $\mathbb{K}[G]$, denoted as $\omega_{\mathbb{K}[G]}$, is isomorphic to an ideal of $\mathbb{K}[G]$ with a generating set that possesses a geometric interpretation, thereby simplifying its computation. Leveraging this understanding, we can determine the top multi-graded Betti numbers of $\mathbb{K}[G]$ by invoking Formula 6.6 from [3]. Note that the top multi-graded Betti numbers of a finitely generated $\mathbb{Z}^{n}$-graded module $M$ refer to the numbers $\beta_{p,h}(M)$ for $h\in\mathbb{Z}^{n}$ if $M$ has a projective dimension $p$. We will also utilize the terminology the second top multi-graded Betti numbers of $M$ to refer to the numbers $\beta_{p-1,h}(M)$ for $h\in\mathbb{Z}^{n}$.
Besides [3, Formula 6.6], our method for computing the Betti numbers of normal edge rings also relies on the following results. The first result is as follows.
Lemma 1.1.
(Lemma 2.8) Let $H$ be an induced subgraph of a simple graph $G$. Then for $h\in\mathbb{Z}_{\geq 0}^{V(G)}$ with $\mathrm{supp}(h)\subseteq V(H)$, we have
$\beta_{i,h}(\mathbb{K}[H])=\beta_{i,h}(\mathbb{K}[G])\qquad\forall i\geq 0.$
The second result is a generalization of Hochster’s formula. Let $T$ represent the subalgebra of the polynomial ring $\mathbb{K}[Y_{1},\ldots,Y_{m}]$ generated by monomials $y_{1},\ldots,y_{n}$. Within $T$, all the monomials form a monoid under multiplication, denoted by $\mathcal{U}$. Let $R$ be the polynomial ring $\mathbb{K}[X_{1},\ldots,X_{n}]$. We endow it with a multi-grading in $\mathbb{Z}^{m}$ by assigning the multi-degree of $y_{i}$ to $X_{i}$, where $i=1,\ldots,n$. There is a multi-graded homomorphism from $R$ to $T$ sending $X_{i}$ to $y_{i}$ for $i=1,\ldots,n$. Consider an ideal $I$ of $T$ generated by certain monomials from $\mathcal{U}$. The formula for expressing $\beta_{i,h}^{R}(T/I)$ has been provided in [4]. However, to compute the Betti numbers of the canonical module of a normal edge ring, we require the following lemma.
Lemma 1.2.
(Lemma 2.2) Let $I$ be a monomial ideal of $T$. To each multi-degree $h\in\mathbb{Z}_{\geq 0}^{m}$, we associate a simplicial complex
$\Gamma_{h}(I):=\{F\subseteq[n]\mid Y^{h}=y^{F}\alpha\mbox{ for some monomial }\alpha\in I\}.$
Then, for all $i\geq 0$, the multi-graded Betti numbers of $I$ are expressed as
$\beta_{i,h}^{R}(I)=\dim_{\mathbb{K}}\widetilde{H}_{i-1}(\Gamma_{h}(I),\mathbb{K}).$
We are ready to introduce our method. Briefly speaking, our method expresses every multi-graded Betti number of a normal edge ring $\mathbb{K}[G]$ as the top or second-top multi-graded Betti number of the edge ring associated with certain induced subgraph of $G$, and the latter can be derived from the data of its canonical module.
The detailed steps of our approach is as follows. Let $G$ be a simple graph that satisfies the odd-cycle condition. Assuming that the projective dimension of $\mathbb{K}[G]$ is $p$, and that there exists a term order $<$ such that $\mathrm{in}_{<}(I_{G}))$ is square-free, and for each $1\leq i\leq p$, a suitable upper bound for the total Betti number $\beta_{i}(\mathbb{K}[E(G)]/\mathrm{in}_{<}(I_{G}))$ has been derived.
  1. (1)
    Initially, we compute the top multi-graded Betti numbers $\beta_{p,h}(\mathbb{K}[G])$ of $\mathbb{K}[G]$ by determining the minimal generators of $\omega_{\mathbb{K}[G]}$.
  2. (2)
    For indices $i$ with $0<i<p$, we identify all of the induced subgraphs $H$ of the graph $G$ such that (i) the projective dimension of $\mathbb{K}[H]$ is equal to $i$; (ii) if $\beta_{i,h}(\mathbb{K}[H])\neq 0$ for some $h\in\mathbb{Z}_{\geq 0}^{V(H)}$, then $\mbox{supp}(h)=V(H)$. Subsequently, by leveraging [3, Formula 6.6], we compute the top multi-graded Betti numbers of the edge rings associated with these induced subgraphs, and then sum them up.
  3. (3)
    If the sum obtained in the second step matches the given upper bound for $\beta_{i}(\mathbb{K}[E(G)]/\mathrm{in}_{<}(I_{G}))$, then we have achieved our goal of computing the multi-graded Betti numbers $\beta_{i,h}(\mathbb{K}[G])$ for $h\in\mathbb{Z}_{\geq 0}^{V(G)}$. In this case, $\beta_{i,h}(\mathbb{K}[G])$ is equal to the top Betti number $\beta_{i,h}(\mathbb{K}[H])$, where $H$ is the induced graph of $G$ on $\mathrm{supp}(h)$. Moreover, if $\beta_{i,h}(\mathbb{K}[G])\neq 0$, then the projective dimension of $\mathbb{K}[H]$ is equal to $i$.
  4. (4)
    Otherwise, we have to consider induced subgraphs with edge rings of projective dimension $i+1$ and to compute the second top Betti numbers of the edge rings of these graphs. This is conducted by [3, Formula 6.6] together with Lemma 1.2.
We refer to this method as the induced-subgraph method. In view of the first step of this method, we introduce the following definition.
Definition 1.3.
Let $G$ be a connected graph that fulfills the odd-cycle condition. Denote by $p$ the projective dimension of $\mathbb{K}[G]$. We call $G$ to be a top-Betti graph provided that if $\beta_{p,h}(\mathbb{K}[G])\neq 0$ for some $h\in\mathbb{Z}_{\geq 0}^{V(G)}$, then $\mathrm{supp}(h)=V(G)$. Here and after, for $h\in\mathbb{Z}_{\geq 0}^{V(G)}$, $\mathrm{supp}(h)$ represents the set of $x_{i}\in V(G)$ with $h(x_{i})\neq 0$.
A combinatorial characterization for top-Betti graphs was presented in Proposition 3.1. Based on this result, it is not difficult to determine all the top-Betti induced subgraphs of any given graph that meets the odd-condition. This fact eases our study.
In the fourth step, we need to discern strict second-Betti induced subgraphs of a given graph, which is defined below.
Definition 1.4.
Let $G$ be a connected graph that fulfills the odd-cycle condition. Denote by $p$ the projective dimension of $\mathbb{K}[G]$. For a subset $S\subseteq V(G)$, let $G[S]$ stand for the induced graph of $G$ on $S$. We call $G$ to be a second-Betti graph provided that if $\beta_{p-1,h}(G)\neq 0$ for some multi-degree $h$, then either $\mathrm{supp}(h)=V(G)$ or the projective dimension of $\mathbb{K}[G[\mathrm{supp}(h)]]$ is equal to $p-1.$ A second-Betti graph is termed strict if there is a multi-degree $h$ such that $\mathrm{supp}(h)=V(G)$ and $\beta_{p-1,h}(G)\neq 0$.
A sufficient condition for $G$ to be a second-Betti subgraph are given in Proposition 3.5, which enables us to discern all strict second-Betti induced subgraphs of the graphs we are studying.
In this paper, we successfully utilize this approach to compute the multi-graded Betti numbers of edge rings associated with two distinct classes of simple graphs that adhere to the odd-cycle condition. To introduce our main results, we need an additional notation.
Definition 1.5.
Let $H$ be a simple graph satisfying the odd-cycle condition. We define $\mathcal{N}_{H}$, which we call the top support of $H$, as the set of multi-degrees $h$ for which $\beta_{p,h}(\mathbb{K}[H])$ is non-zero, where $p$ is the projective dimension of $\mathbb{K}[H]$. It is evident that the top multi-graded Betti numbers of $\mathbb{K}[H]$ are determined by the set $\mathcal{N}_{H}$.
The first class of graphs under consideration is $\mathbf{G}_{m}$, which was introduced in [14] for all $m\geq 1$. Its precise definition could be found in Definition 4.1. Owing to its shape, as depicted in Figure 1, we refer to the graph $\mathbf{G}_{m}$ as a two-ear graph. In this paper, we show that the Betti numbers of the edge ring of $\mathbf{G}_{m}$ could be computed using the initial three steps of the induced-subgraph method.
Theorem 1.6.
(Theorems 4.8 and 4.11) Let $G$ denote a two-ear graph $\mathbf{G}_{m}$ for some $m\geq 3$. For $i\geq 1$, let $\mathcal{N}_{i}(G)$ denote the disjoint union of all $\mathcal{N}_{H}$ such that $H$ is a top-Betti induced subgraph of $G$ whose edge ring has a projective dimension of $i$. Then for any multi-degree $\alpha\in\mathbb{Z}_{\geq 0}^{V(G)}$ and for any integer $i\geq 1$, we have
$\beta_{i,\alpha}(\mathbb{K}[G])=\left\{\begin{array}[]{ll}1,&\hbox{$\alpha\in\mathcal{N}_{i}(G)$;}\\ 0,&\hbox{otherwise.}\end{array}\right.$
In particular, there exists a square-free initial ideal $\mathrm{in}_{<}(I_{G})$ of $I_{G}$ such that
$\beta_{i,j}(\mathrm{in}_{<}(I_{G}))=\beta_{i,j}(I_{G})$ for all $i,j$.
We note that there are exactly three types of top-Betti induced subgraphs in a two-ear graph, namely, two-ear graphs, one-ear graphs and complete bipartite graphs of the form $\mathbf{K}_{2,d}$. The top supports for these subgraphs are provided in Lemma 4.5, and Corollaries 4.9 and 4.10, respectively.
The second class of graphs under consideration is compact graphs. According to [25], a simple graph is compact if it lacks even cycles and satisfies the odd-cycle condition. A vertex of a compact graph is said to be a big vertex if it has a degree greater than 2. To a large extent, the structure of a compact graph is determined by the quantity of its big vertices. In view of [25, Proposition 3.3], there are at most three big vertices in a compact graph. In the scenario where $G$ possesses three big vertices, i.e., $G$ is a compact graph of type 3, the fourth step of the induced-graph method becomes necessary. The multi-graded Betti numbers of a compact graph are presented in the following theorem.
Theorem 1.7.
(Theorem 6.9 and Corollary 6.10) Let $G$ be a compact graph, and denote by $p$ the projective dimension of $\mathbb{K}[G]$. Consider an integer $i$ satisfying $1\leq i\leq p$. Let $H_{j}$ for $j=1,\ldots,s$ represent all the compact induced subgraphs of $G$ having exactly $i+1$ minimal cycles, and let $L_{j}$ for $j=1,\ldots,t$ be all the compact induced subgraphs of $G$ belonging to type 3 with $i+2$ minimal cycles. Define the following sets:
$\mathcal{N}_{i}(G)=\bigcup_{j=1}^{s}\mathcal{N}_{H_{j}},\quad\mathcal{M}_{i}^{1}(G)=\bigcup_{j=1}^{t}\mathcal{M}_{L_{j}}^{1},\quad\text{and}\quad\mathcal{M}_{i}^{2}(G)=\bigcup_{j=1}^{t}M_{L_{j}}^{2}.$
Then, the $i$-th multi-graded Betti numbers of $\mathbb{K}[G]$ can be expressed as follows:
$\beta_{i,\alpha}(\mathbb{K}[G])=\begin{cases}2,&\text{if }\alpha\in\mathcal{M}_{i}^{2}(G),\\ 1,&\text{if }\alpha\in\mathcal{M}_{i}^{1}(G)\cup\mathcal{N}_{i}(G),\\ 0,&\text{otherwise}.\end{cases}$
Here, $\mathcal{N}_{H},\mathcal{M}_{H}^{1}$ and $\mathcal{M}_{H}^{2}$ are sets of monomials that could be read from the compact graph $H$ directly if $H$ is labeled through Lemma 5.3, Lemma 5.5 and Lemma 6.7.
In particular, there exists a square-free initial ideal $\mathrm{in}_{<}(I_{G})$ of $I_{G}$ such that
$\beta_{i,j}(\mathrm{in}_{<}(I_{G}))=\beta_{i,j}(I_{G})$ for all $i,j$.
We note that compact graphs of type 3 are strict second-Betti graphs, while compact graphs of type 1 or 2 are not. Theorems 1.3 and 1.4 demonstrate that the toric ideals of the edge rings associated with compact graphs and two-ear graphs precisely satisfy the requirements of Task. These results suggest that the induced-subgraph method is applicable to all normal edge rings. In the subsequent work [26] of this paper, we shift our focus to another class of normal edge rings, namely the edge rings of multi- path graphs. We demonstrate by the induced-subgraph method that the toric ideals of these rings meet the requirements specified in Task as well. This further verifies the validity of our speculation.
The paper is structured as follows. In Section 2, we lay the groundwork by introducing fundamental concepts and preliminary results. In Section 3, we present basic properties of top-Betti graphs and second-Betti graphs. Subsequently, in section 4 we compute the multi-graded Betti numbers of edge rings associated with two-ear graphs. As induced subgraphs of two-ear graphs, the multi-graded Betti numbers of edge rings associated with one-ear graphs and complete bipartite graphs with one part consisting of two elements are also given. In Sections 5 and 6, we explore the Betti-numbers of the edge rings of compact graphs, culminating in the presentation of Theorem 6.9.
2. Preliminaries
In this section, we provide a brief review of the notation and fundamental facts that will be utilized later on.
2.1. Betti numbers and Monomial ideals with regular quotients
Let $\mathbb{K}$ be a field, and let $R:=\mathbb{K}[X_{1},\ldots,X_{n}]$ be the polynomial ring in variables $X_{1},\ldots,X_{n}$ that are $\mathbb{Z}^{m}$-graded. Here, we do not require that $m=n$. For a finitely generated $\mathbb{Z}^{m}$-graded $R-$module $M$, there exists the minimal multi-graded free resolution of $M$ that has the form:
$0\rightarrow\underset{h\in\mathbb{Z}^{m}}{\bigoplus}R[-h]^{\beta_{p,h}(M)}\rightarrow\cdots\rightarrow\underset{h\in\mathbb{Z}^{m}}{\bigoplus}R[-h]^{\beta_{1,h}(M)}\rightarrow\underset{h\in\mathbb{Z}^{m}}{\bigoplus}R[-h]^{\beta_{0,h}(M)}\rightarrow M\rightarrow 0$
Note that $R[-h]$ is the cyclic free $R$-module generated in degree $h$. The projective dimension of $M$ are defined to be $\mathrm{pdim}\,(M):={\mbox{max}}\,\{i\mid\beta_{i,\,h}(M)\neq 0\mbox{ for some}h\}.$ The number $\beta_{i,h}(M):={\rm{dim}}_{\mathbb{K}}\mathrm{Tor}_{i}^{R}(M,\mathbb{K})_{h}$ is called the $(i,h)$-th multi-graded Betti number of $M$ and $\beta_{i}(M):=\sum_{h\in\mathbb{Z}^{m}}\beta_{i,h}(M)$ is called the $i$-th total Betti number of $M$.
Denote by $p$ the projective dimension of the module $M$. Subsequently, $\beta_{p}(M)$ is referred to as the top total Betti number of $M$, while $\beta_{p,h}(M)$ for $h\in\mathbb{Z}^{m}$ are designated as the top multi-graded Betti numbers of $M$. On the other hand, $\beta_{p-1}(M)$ is labeled as the second top total Betti number of $M$, and $\beta_{p-1,h}(M)$ for $h\in\mathbb{Z}^{m}$ are designated as the second top multi-graded Betti numbers of $M$.
To derive upper bounds for Betti numbers of toric ideals under our investigation, we need a notion that was originally introduced in [16] and subsequently explored in [15]. Recall that a monomial ideal $I$ in $R$ is said to possess regular quotients if there exists a linear ordering $\alpha_{1}\prec\alpha_{2}\prec\ldots\prec\alpha_{r}$ among the minimal generators $\{\alpha_{1},\alpha_{2},\ldots,\alpha_{r}\}$ of $I$ such that, for each $s=2,\ldots,r$, the quotient ideal $(\alpha_{1},\ldots,\alpha_{s-1}):\alpha_{s}$ is generated by a regular sequence. Notably, a sequence of monomials $\beta_{1},\ldots,\beta_{r}$ is regular precisely when the supports of distinct monomials are disjoint, i.e., $\mathrm{supp}(\beta_{i})\cap\mathrm{supp}(\beta_{j})=\emptyset$ for all $1\leq s<t\leq r$. Here, the support of a monomial $\alpha$, denoted by $\mathrm{supp}(\alpha)$, refers to the set of variables that divide $\alpha$.
Assuming that the quotient ideal $(\alpha_{1},\ldots,\alpha_{s-1}):\alpha_{s}$, denoted as $L_{s}$, is generated by a regular sequence of length $k_{s}$ for each $s$ ranging from $2$ to $r$, we can apply Koszul theory to deduce that the total Betti numbers $\beta_{i}(R/L_{s})$ are given by $\beta_{i}(R/L_{s})=\binom{k_{s}}{i}$ for all non-negative integers $i$. Let $I_{s}$ represent the monomial ideal generated by $\alpha_{1},\ldots,\alpha_{s}$, and let $d_{s}$ represent the degree of $\alpha_{s}$ for each $s$ from $1$ to $r$.
Now, consider the following short exact sequences:
$0\rightarrow R/L_{s}[-d_{s}]\rightarrow R/I_{s-1}\rightarrow R/I_{s}\rightarrow 0,\quad\text{for all}\ s=2,\ldots,r.$
Utilizing the mapping cone construction (see e.g. [21, section 27] for the details), we can deduce that
$\beta_{i}(R/I_{s})\leq\beta_{i-1}(R/L_{s})+\beta_{i}(R/I_{s-1}).$
Repeatedly applying the above inequality, we arrive at
$\beta_{i}(R/I)\leq\beta_{i-1}(R/L_{r})+\cdots+\beta_{i-1}(R/L_{2})+\beta_{i}(R/(\alpha_{1}))$
and
$\beta_{i}(I)\leq\beta_{i}(R/L_{r})+\cdots+\beta_{i}(R/L_{2})+\beta_{i}(R).$
This implies that, for all non-negative integers $i$, we have the following inequality, which will be crucial in our subsequent discussions:
(1) $\beta_{i}(I)\leq\sum_{j=2}^{r}\binom{k_{j}}{i}+\binom{0}{i}.$
Apart from this formula, the following result, as mentioned, for instance, in [13, Corollary 3.3.3], is necessary for us to derive the upper bound for Betti numbers of edge rings.
Lemma 2.1.
Let $I$ be a homogeneous ideal of the polynomial ring $R=\mathbb{K}[X_{1},\ldots,X_{n}]$, and let $<$ be any monomial order on $R$. Then for all $i,j\geq 0$, we have
$\beta_{i,j}(R/I)\leq\beta_{i,j}(R/\mathrm{in}_{<}(I)).$
In particular,
$\beta_{i}(R/I)\leq\beta_{i}(R/\mathrm{in}_{<}(I)).$
2.2. Betti numbers of monomial ideals in a monomial algebras
Let $T$ represent the subalgebra of the polynomial ring $\mathbb{K}[Y_{1},\ldots,Y_{m}]$ generated by monomials $y_{1},\ldots,y_{n}$. In [23], $T$ is called a monomial algebra. Within $T$, all the monomials form a monoid under multiplication, denoted by $\mathcal{U}$, and $T$ is isomorphic to the positive affine monoid ring $\mathbb{K}[\mathcal{U}]$. We define $R$ as the polynomial ring $\mathbb{K}[X_{1},\ldots,X_{n}]$ and endow it with a multi-grading in $\mathbb{Z}^{m}$ by assigning the multi-degree of $y_{i}$ to $X_{i}$. Consider an ideal $I$ of $T$ generated by certain monomials from $\mathcal{U}$.
By the existence of a surjective homomorphism from $R$ to $T$ mapping $X_{i}$ to $y_{i}$, it follows that $T$, $I$ and $T/I$ are all finitely generated $\mathbb{Z}^{m}$-graded $R$-modules.
A formula for computing of the multi-graded Betti numbers of $T/I$ was previously established in [4] using relative simplicial homology. However, in this paper, we require to calculate the multi-graded Betti numbers of $I$. For this, we associate every degree $h\in\mathbb{Z}_{\geq 0}^{m}$ with the following simplicial complex:
$\Gamma_{h}(I):=\{F\subseteq[n]\mid Y^{h}=y^{F}\alpha\mbox{ for some monomial }\alpha\in I\}.$
Here, $Y^{h}$ is defined as $Y_{1}^{h_{1}}\cdots Y_{m}^{h_{m}}$ for $h=(h_{1},\ldots,h_{m})\in\mathbb{Z}_{\geq 0}^{m}$, and $y^{F}$ represents the product of all monomials $y_{i}$ with $i\in F$.
Lemma 2.2.
Let $I$ be a monomial ideal of $T$, and let $h\in\mathbb{Z}_{\geq 0}^{m}$ be a multi-degree. Then, for all $i\geq 0$, one has
$\beta_{i,h}^{R}(I)=\dim_{\mathbb{K}}\widetilde{H}_{i-1}(\Gamma_{h}(I),\mathbb{K}).$
Proof.
Let $K_{\bullet}$ denote the Koszul complex of $X_{1},\ldots,X_{n}$. Namely, $K_{\bullet}$ is the following complex
$0\rightarrow K_{n}\rightarrow\cdots\rightarrow K_{i}\rightarrow\cdots\rightarrow K_{0}\rightarrow 0$
Here, $K_{i}$ is a multi-graded free $R$-module that possesses a base composed of elements $e_{F}$ such that $|F|=i$ and $F\subseteq[n]$. Moreover,
$\deg(e_{F})=\sum_{i\in F}\deg(X_{i})=\sum_{i\in F}\deg(y_{i})\in\mathbb{Z}_{\geq 0}^{m}.$
In particular, $K_{0}=Re_{\emptyset}=R$. Since $K_{\bullet}$ is a multi-graded minimal free resolution of $\mathbb{K}$, we have
$\beta_{i,h}^{R}(I)=\dim_{\mathbb{K}}\mathrm{Tor}_{i}^{R}(\mathbb{K},I)_{h}=\dim_{\mathbb{K}}H_{i}(K_{\bullet}\otimes_{R}I)_{h}.$
On the other hand,
$(K_{i}\otimes I)_{h}=\bigoplus_{|F|=i}(Re_{F}\otimes_{R}I)_{h}=\bigoplus_{|F|=i}(I(\deg(e_{F}))_{h}.$
Note that
$(I(\deg(e_{F}))_{h}=I_{h-\deg(e_{F})}=\left\{\begin{array}[]{ll}\mathbb{K},&\hbox{$F\in\Gamma_{h}(I)$;}\\ 0,&\hbox{otherwise,}\end{array}\right.$
it follows that there is an isomorphism between the complex $K_{\bullet}\otimes_{R}I$ and the complex $\widetilde{C}_{\bullet}(\Gamma_{h}(I),\mathbb{K})[-1]$, yielding the desired formula.
We may look up this lemma as a generalization of [17, Theorem 1.34], which address the same question on monomial ideals in a polynomial ring.
2.3. Simple graphs and minimal cycles
By definition, a simple graph is a finite graph devoid of loops and multiple edges. Let $G$ be a simple graph and denote the vertex set and edge set of $G$ by $V(G)$ and $E(G)$ respectively. Recall that a walk of $G$ of length $q$ is a subgraph $W$ of $G$ such that $E(W)=\{\{i_{0},i_{1}\},\{i_{1},i_{2}\},\ldots,\{i_{q-1},i_{q}\}\}$, where $i_{0},i_{1},\ldots,i_{q}$ are vertices of $G$. The graph $G$ is connected if, for any two vertices $i_{j}$ and $i_{k}$ of $G$, there is a walk between $i_{j}$ and $i_{k}$.
A cycle of $G$ of length $q$ is a subgraph $C$ of $G$ such that
$E(C)=\{\{i_{1},i_{2}\},\{i_{2},i_{3}\},\ldots,\{i_{q-1},i_{q}\},\{i_{q},i_{1}\}\}$
where $i_{1},i_{2},\ldots,i_{q}$ are vertices of $G$ and where $i_{j}\neq i_{k}$ if $j\neq k$. A cycle is called even (resp. odd) if $q$ is even (resp. odd). For a subset $W$ of $V(G)$, the induced subgraph $G_{W}$ is the graph with vertex set $W$ and for every pair $j,k\in W$, they are adjacent in $G_{W}$ if and only if they are adjacent in $G$. For a vertex $j\in V(G)$, let $G\setminus j$ be the induced subgraph $G_{V(G)\setminus\{j\}}$. A cycle of $G$ is called an induced cycle if it is an induced subgraph of $G$. We also call an induced cycle of $G$ to be a minimal cycle of $G$, and use $\mathfrak{t}(G)$ to denote the number of the minimal cycles of $G$.
2.4. Conventions and notions
We set conventions and fix notions that play an important role in the rest of this paper.
Conventions 2.3.
Let $G$ be a simple graph, with its vertices labelled by letters or subscripted letters, such as $u,u_{i,j},v,v_{i,j},w,w_{i,j}$, and so forth. We establish a correspondence between these labels and unit vectors in ${\mathbb{R}}^{V(G)}$ as follows:
$u,u_{i,j},v,v_{i,j},w,w_{i,j},x,y$
correspond to
${\mathbf{u}},{\mathbf{u}}_{i,j},{\mathbf{v}},{\mathbf{v}}_{i,j},{\mathbf{w}},{\mathbf{w}}_{i,j},{\mathbf{x}},{\mathbf{y}}$
respectively. With this notation, we can express ${\mathbb{R}}^{V(G)}$ as the set
$\left\{\sum_{i,j}a_{i,j}{\mathbf{u}}_{i,j}+a{\mathbf{u}}+\sum_{i,j}b_{i,j}{\mathbf{v}}_{i,j}+b{\mathbf{v}}+\sum_{i,j}c_{i,j}{\mathbf{w}}_{i,j}+c{\mathbf{w}}\mid a_{i,j},a,b_{i,j},b,c_{i,j},c\in{\mathbb{R}}\right\}.$
By way of example, $\deg(uv)={\mathbf{u}}_{i}+{\mathbf{v}}_{j}$, and $\deg(u^{2}wu_{1,2})=2{\mathbf{u}}+{\mathbf{w}}+{\mathbf{u}}_{1,2}$, and so on.
We identify a monomial with its multi-degree. For example, we will write $\beta_{i,uv}(\mathbb{K}[G])$ for $\beta_{i,{\mathbf{u}}+{\mathbf{v}}}(\mathbb{K}[G])$. Under this identification, we look up $\mathbb{Z}^{V(H)}$ as a subset of $\mathbb{Z}^{V(G)}$ if $H$ is an induced subgraph of $G$.
Adopting these conventions, for a simple graph $G$ with vertex set $V(G)$, both the edge ring $\mathbb{K}[G]$ and its toric ideal $I_{G}$ possess an $\mathbb{Z}_{\geq 0}^{V(G)}$-grading. Consequently, it is permissible to consider the $\mathbb{Z}_{\geq 0}^{V(G)}$-graded Betti numbers of $\mathbb{K}[G]$ and $I_{G}$. Additionally, by assigning a degree of 1 to each edge $e$ in $E(G)$, both $\mathbb{K}[G]$ and $I_{G}$ acquire a $\mathbb{Z}$-grading. The relationship between the $\mathbb{Z}_{\geq 0}^{V(G)}$-graded Betti numbers and the $\mathbb{Z}_{\geq 0}$-graded Betti numbers is as follows:
$\beta_{i,j}(\mathbb{K}[G])=\sum_{\begin{subarray}{c}h\in\mathbb{Z}_{\geq 0}^{V(G)}\end{subarray}}^{|h|=2j}\beta_{i,h}(\mathbb{K}[G]).$
Notation 2.4.
Let $H$ be a simple graph satisfying the odd-cycle condition. We define $D_{H}$ to be the multi-degree $\sum_{e\in E(H)}\mathbf{v}_{e}$. Since we identify a monomial with its multi-degree, $D_{H}$ can alternatively be interpreted as the product of all edges of $H$. Here, an edge is considered as the product of its two vertices. Furthermore, we employ $\Theta_{H}$ to represent the product of all the vertices.
It is straightforward to observe that $\Theta_{H}^{2}$ divides $D_{H}$.
2.5. Edge cones and canonical modules
Let $G$ be a simple graph with vertex set $V(G)=\{1,\ldots,n\}$ and edge set $E(G)$. For any edge $f=\{i,j\}\in E(G)$, we define $\mathbf{v}_{f}=\mathbf{e}_{i}+\mathbf{e}_{j}$, where $\mathbf{e}_{i}$ represents the $i$th unit vector in $\mathbb{R}^{n}$.
The edge cone of $G$, denoted as $\mathbb{R}_{+}(G)$, is the cone in $\mathbb{R}^{n}$ spanned by $\{\mathbf{v}_{f}|f\in E(G)\}$. Specifically, it can be expressed as:
$\mathbb{R}_{+}(G):=\left\{\sum_{f\in E(G)}{a_{f}\mathbf{v}_{f}}\bigg{|}a_{f}\in\mathbb{R}_{\geq 0}\text{ for all }f\in E(G)\right\}.$
Let us express the edge cone ${\mathbb{R}}_{+}(G)$ in terms of linear inequalities, or equivalently, through its facets. To achieve this, additional notation is required. Let $S$ be a subset of $V(G)$. The induced subgraph of $G$ on $S$, denoted by $G[S]$, is the graph with vertex set $S$ and edge set $\{e\in E(G)|e\subseteq S\}$. A non-empty subset $T\subseteq V(G)$ is called an independent set of $G$ if $\{j,k\}\not\in E(G)$ for any $j,k\in T$. For an independent set $T\subseteq V(G)$, we use $N_{G}(T)$ (or simply $N(T)$ for brevity) to denote the neighborhood set of $T$, namely,
$N_{G}(T):=\{j\in V(G)\mid\{j,k\}\in E(G)\mbox{ for some }k\in T\}.$
Given a vertex $i$ of $G$, let $H_{i},H_{i}^{+}$ and $H_{i}^{>}$ denote the following equalities or inequalities respectively:
$H_{i}:x_{i}=0,\qquad H_{i}^{+}:x_{i}\geq 0,\qquad H_{i}^{>}:x_{i}>0.$
Given an independent set $T$ of $G$, let $H_{T},H_{T}^{-}$ and $H_{T}^{<}$ denote the following equalities or inequalities respectively:
$H_{T}:\sum_{i\in T}x_{i}=\sum_{i\in N(T)}x_{i},\qquad H_{T}^{-}:\sum_{i\in T}x_{i}\leq\sum_{i\in N(T)}x_{i},\qquad H_{T}^{<}:\sum_{i\in T}x_{i}<\sum_{i\in N(T)}x_{i}.$
Note that $H_{i}\neq H_{\{i\}}$, since the latter represents the equality $x_{i}=\sum_{j\in N(i)}x_{j}$. If $G$ is bipartite, the facets of ${\mathbb{R}}_{+}(G)$ are given in [24, Proposition 3.6], that we record in the following result for convenience.
Lemma 2.5.
If $G$ is a bipartite connected graph with bipartition $(V_{1},V_{2})$, then $F$ is a facet of ${\mathbb{R}}_{+}(G)$ if and only if either
  1. $(a)$
    $F=H_{i}\cap{\mathbb{R}}_{+}(G)$ for some $i\in V(G)$ with $G\setminus i$ connected; or
  2. $(b)$
    $F=H_{T}\cap{\mathbb{R}}_{+}(G)$ for some $T\varsubsetneq V_{1}$, such that $G[T\cup N(T)]$ and $G[V(G)\setminus(T\cup N(T))]$ connected.
Let $\mathcal{I}$ denote the set of all vertices satisfying condition (a) and $\mathcal{T}$ denote the set of all independent sets satisfying (b). Then, since $\mathrm{aff}(\mathbb{R}_{+}(G))=H_{V_{1}}=H_{V_{2}}$, we have
($\Delta_{1}$) $\mathbb{R}_{+}(G)=H_{V_{1}}\cap\bigcap_{i\in\mathcal{I}}H_{i}^{+}\cap\bigcap_{T\in\mathcal{T}}H_{T}^{-}$
and
$\mathrm{relint}(\mathbb{R}_{+}(G))=H_{V_{1}}\cap\bigcap_{i\in\mathcal{I}}H_{i}^{>}\cap\bigcap_{T\in\mathcal{T}}H_{T}^{<}.$
For the case when $G$ is non-bipartite, we need to introduce the following notions on graphs.
  • We call a vertex $j$ of $G$ regular if each connected component of $G\setminus j$ contains an odd cycle.
  • We say that an independent set $T$ of $V(G)$ is a fundamental set if
    • the bipartite graph on the vertex set $T\cup N(T)$ with the edge set $E(G)\cap\{\{j,k\}\mid j\in T,k\in N(T)\}$ is connected, and
    • either $T\cup N(T)=V(G)$ or each of the connected components of the graph $G[V(G)\setminus(T\cup N(T))]$ contains an odd cycle.
We herby record a result from [23, Theorem 10.7.6] as follows:
Lemma 2.6.
Let $G$ be a non-bipartite connected simple graph on $V(G)=\{1,\ldots,n\}$. Then ${\mathbb{R}}_{+}(G)$ consists of the elements $(x_{1},\ldots,x_{n})\in{\mathbb{R}}^{n}$ satisfying all the following inequalities:
($\Delta_{2}$) $\begin{split}x_{i}&\geq 0\;\;\text{ for any regular vertex }i;\\ \sum\limits_{i\in N(T)}x_{i}&\geq\sum\limits_{i\in T}x_{i}\;\;\text{ for any fundamental set }T.\end{split}$
It is established by [19, Corollary 2.3] that if $G$ satisfies the odd-cycle condition, then the edge ring $\mathbb{K}[G]$ is normal. Let $\mathcal{S}(G)$ represent the semigroup composed of all $\alpha\in\mathbb{Z}^{n}$ with $x^{\alpha}\in\mathbb{K}[G].$ According to [3, Theorem 6.31], the ideal of $\mathbb{K}[G]$ generated by all monomials $x^{h}$ with $h\in\text{relint}(\mathbb{R}{+}(G))\cap\mathcal{S}(G)$ constitutes the $\mathbb{Z}^{V(G)}$-graded canonical module of the $\mathbb{Z}^{V(G)}$-graded local algebra $\mathbb{K}[G]$, which we denote by $\omega_{\mathbb{K}[G]}$. Then, based on [3, Formula 6.6], we are able to derive the following result.
Lemma 2.7.
Let $G$ be a simple graph satisfying the odd-cycle condition and denote by $p$ the projective dimension of $\mathbb{K}[G]$. Then
$\beta_{p-i,h}(\mathbb{K}[G])=\beta_{i,D_{G}-h}(\omega_{\mathbb{K}[G]})\qquad\mbox{ for all }\quad 0\leq i\leq p\mbox{ and for all }h\in\mathbb{Z}_{\geq 0}^{V(G)}.$
The following two lemmas are important, especially the former, which serves as the cornerstone of the induced-subgraph method. Although they are known, we provide proofs for the completeness.
Lemma 2.8.
Let $H$ be an induced subgraph of a simple graph $G$. Then for $h\in\mathbb{Z}_{\geq 0}^{V(G)}$ with $\mathrm{supp}(h)\subseteq V(H)$, we have
$\beta_{i,h}(\mathbb{K}[H])=\beta_{i,h}(\mathbb{K}[G])\qquad\mbox{ for all }i\geq 0.$
Proof.
Since $\mathrm{supp}(h)\subseteq V(H)$ we have $\Gamma_{h}(\mathbb{K}[H])=\Gamma_{h}(\mathbb{K}[G])$ by definition. From this fact, the result follows directly in view of Lemma 2.2.
Let $c_{0}(G)$ denote the number of bipartite connected components of a simple $G$.
Lemma 2.9.
Let $G$ be a simple graph with vertex set $V(G)$ and edge set $E(G)$. If $G$ fulfills the odd-condition, then $\mathrm{pdim}(\mathbb{K}[G])=|E(G)|-|V(G)|+c_{0}(G).$
Proof.
First consider the case when $G$ is connected. By [22, Proposition 13.15] and [4, Theorem 6.3.5], $\mathbb{K}[G]$ is Cohen-Macaulay. It follows from [23, Corollary 10.1.21] that
$\mathrm{depth}(\mathbb{K}[G])=\mathrm{dim}(\mathbb{K}[G])=\left\{\begin{array}[]{ll}|V(G)|-1,&\text{if $G$ is bipartite;}\\ |V(G)|,&\text{otherwise.}\end{array}\right.$
According to the Auslander-Buchsbaum formula, as stated, for instance, in [13, Corollary A.4.3], we have
$\mathrm{pdim}(\mathbb{K}[G])=\left\{\begin{array}[]{ll}|E(G)|-|V(G)|+1,&\text{if $G$ bipartite;}\\ |E(G)|-|V(G)|,&\text{otherwise.}\end{array}\right.$
Next, assume $G$ has connected components $G_{1},\ldots,G_{k}$. Then,
$\mathrm{pdim}(\mathbb{K}[G])=\mathrm{pdim}(\mathbb{K}[G_{1}])+\cdots+\mathrm{pdim}(\mathbb{K}[G_{k}]).$
Now, the result follows by combining the above two equalities.
3. top-Betti graphs and second-Betti graphs
In the present section, we provide a combinatorial characterization of a top-Betti graph, and establish a sufficient condition for a graph to be a second-Betti graph. Regarding the definitions of a top-Betti graph and a (strict) second-Betti graph, we refer to Definitions 1.3 and 1.4.
Proposition 3.1.
Let $G$ be a connected graph that fulfills the odd-cycle condition.
$\mathrm{(1)}$ If $G$ is bipartite, then $G$ is a top-Betti graph if and only if for every $x\in V(G)$, one has $c_{0}(G\setminus x)<\deg x.$
$\mathrm{(2)}$ If $G$ is non-bipartite, then $G$ is a top-Betti graph if and only if for every $x\in V(G)$, one has $c_{0}(G\setminus x)<\deg x-1.$
Proof.
We only prove (1) since (2) can be proved similarly. Let $p$ be the projective dimension of $\mathbb{K}[G]$. Assume that there is $x\in V(G)$ such that $c_{0}(G\setminus x)\geq\deg x.$ Then $c_{0}(G\setminus x)=\deg x$. Note that $|E[G\setminus x]|=|E(G)|-\deg x$ and $|V(G\setminus x)|=|V(G)|-1$. Then, we have
$\mathrm{pdim}(\mathbb{K}[G\setminus x])=|E(G)|-\deg x-(|V(G)|-1)+c_{0}(G)=p.$
It follows that there is a multi-degree $h\in\mathbb{Z}_{\geq 0}^{V(G)}$ with $\mathrm{supp}(h)\subseteq V(G)\setminus\{x\}$ such that $\beta_{p,h}(\mathbb{K}[G\setminus x]))\neq 0$. Hence, we have $\beta_{p,h}(\mathbb{K}[G])\neq 0$ by Lemma 2.8, which implies $G$ is not a Betti-graph.
Conversely, if $G$ is not a top-Betti graph, then there is a multi-degree $h\in\mathbb{Z}_{\geq 0}^{V(G)}$ with $\mathrm{supp}(h)\subsetneq V(G)$ such that $\beta_{p,h}(\mathbb{K}[G])\neq 0$. Let $x\in V(G)\setminus\mathrm{supp}(h)$. Since $\mathrm{supp}(h)\subseteq V(G)\setminus\{x\}$, we have $\beta_{p,h}(\mathbb{K}[G\setminus x])\neq 0$ by Lemma 2.8. It follows that $\mathrm{pdim}(\mathbb{K}[G\setminus x])=p$, and thus $c_{0}(G\setminus x)=\deg x$, completing the proof.
Corollary 3.2.
Every vertex of a top-Betti graph $G$ has a degree at least two.
Proof.
Assume on the contrary that $\deg x=1$ for some $x\in V(G)$. Then, it follows that $c_{0}(G\setminus x)=1$ if $G$ is bipartite and $c_{0}(G\setminus\{x\})=0$ if $G$ is non-bipartite. This implies $G$ is not a top-Betti graph by Proposition 3.1, a contradiction.
Example 3.3.
Let $\mathbf{K}_{m,n}$ denote the complete bipartite graph with two parts consisting of $m$ and $n$ vertices respectively. Then $\mathbf{K}_{m,n}$ is top-Betti graph if and only if $m\geq 2$ and $n\geq 2$.
Example 3.4.
A 2-connected graph is a graph that remains connected even after the removal of any single vertex and its incident edges. Let $G$ be a 2-connected. Then $G$ is not a top-Betti graph if and only if $G$ is a non-bipartite and there is $x\in G$ such that $G\setminus x$ is a bipartite graph. One such graphs is the graph consisting of a triangle and a rectangle with a common edge, depicted as follows:
In general, we can use the following results to find all the strict second-Betti induced subgraphs of a given graph.
Proposition 3.5.
If $G$ is a top-Betti graph, then $G$ is a second-Betti graph
Proof.
From the proof of Proposition 3.1, we see that if $G$ fulfills the odd-cycle condition, then $G$ is a top-Betti graph if and only if $\mathrm{pdim}(\mathbb{K}[H])<\mathrm{pdim}(\mathbb{K}[G])$ for every induced subgraph $H$ of $G$ with $H\neq G$. Let $h$ be a multi-degree such that $\beta_{p-1,h}(\mathbb{K}[G])\neq 0$. We are done if $\mathrm{supp}(h)=V(G)$. Suppose that $\mathrm{supp}(h)\subsetneqq V(G)$. Let $H$ be the subgraph of $G$ induced on $\mathrm{supp}(h)$. Then we have $\mathrm{pdim}(\mathbb{K}[H])<\mathrm{pdim}(\mathbb{K}[G])$, and thus $\mathrm{pdim}(\mathbb{K}[H])\leq p-1$. However, we have $\beta_{p-1,h}((\mathbb{K}[H]))=\beta_{p-1,h}((\mathbb{K}[G]))\neq 0$. Hence, $\mathrm{pdim}(\mathbb{K}[H])=p-1$, as desired.
If $G$ has a vertex $x$ of degree 1, then the edge rings of $G$ and $G\setminus x$ have the same multi-graded Betti numbers. See [25, Lemma 2.1] for the details. This observation yields the following result.
Lemma 3.6.
If $G$ is a strict top-Betti graph, then every vertex of $G$ has a degree at least 2.
4. Betti numbers of $\mathbf{G}_{m}$
In [14], a specific class of simple graphs, designated as $\mathbf{G}_{m}$, was introduced for all $m\geq 1$. It was proved there that the toric ideal of $\mathbb{K}[\mathbf{G}_{m}]$ admits two different initial ideals. One of them is square-free, which implies that it is Cohen-Macaulay, but the other fails to be Cohen-Macaulay. We refer to such a graph as a two-ear graph due to its shape. In this section, we employ the induced-subgraph method to demonstrate that the toric ideal of $\mathbf{G}_{m}$ possesses exactly the same Betti numbers as the square-free initial ideal of the two initial ideals, as previously mentioned. Furthermore, we present all of its multi-graded Betti numbers.
Definition 4.1.
Let $m\geq 3$ be an integer. Following [14], we define $\mathbf{G}_{m}$ as a simple graph with vertex set
$V(\mathbf{G}_{m})=\{v_{1},v_{2},x_{1},x_{2},y_{1},y_{2}\}\cup\{u_{i}~{}|~{}1\leq i\leq m-1\}$
and with edge set
[…]
$\quad\cup\{\{v_{1},u_{i}~{}|~{}1\leq i\leq m-1\}\}\cup\{\{v_{2},u_{i}~{}|~{}1\leq i\leq m-1\}\}.$
Let $\mathbf{P}_{m}$ denote the induced subgraph $\mathbf{G}_{m}\setminus x_{2}$ of $\mathbf{G}_{m}$. We call a graph $G$ a two-ear graph of type $m$ or an one-ear graph of type $m$ if it is isomorphic to the graph $\mathbf{G}_{m}$ or $\mathbf{P}_{m}$, respectively.
$v_{1}$$v_{2}$$x_{1}$$u_{m-1}$$y_{1}$$u_{2}$$u_{1}$$x_{2}$$y_{2}$$e_{m-1,1}$$e_{m-1,2}$$e_{2,1}$$e_{2,2}$$e_{1,1}$$e_{1,2}$$f_{1}$$g_{1}$$f_{2}$$g_{2}$$f_{3}$$g_{3}$$w$
Figure 1. The two-ear graph $\mathbf{G}_{m}$
Let us denote $\mathbf{G}_{m}$ as $G$. To write out the minimal generators of $\mathrm{in}_{<}(I_{G})$, we label the edges of $G$ as follows: for $i\in[m-1]$, $e_{i,1}=\{v_{1},u_{i}\},e_{i,2}=\{v_{2},u_{i}\}$, and $f_{1}=\{v_{1},x_{1}\},f_{2}=\{x_{1},x_{2}\},f_{3}=\{v_{1},x_{2}\},g_{1}=\{v_{2},y_{1}\},g_{2}=\{y_{1},y_{2}\},g_{3}=\{v_{2},y_{2}\},w=\{x_{1},y_{1}\}$. In [14, Corollary 4.3 ], the minimal set of monomial generators of $\mathrm{in}_{<}(I_{G})$ is specified for a particular monomial order $<$. Note that the monomial order $<$ in this section refers to this order. We record this result in the following lemma for later use.
Lemma 4.2.
There is a monomial order $<$ such that the minimal set of monomial generators of $\mathrm{in}_{<}(I_{G})$ is given by $\mathcal{M}=\mathcal{M}_{1}\cup\mathcal{M}_{2}\cup\mathcal{M}_{3}\cup\mathcal{M}_{4}$, where
  1. $(i)$
[…]
  2. $(ii)$
    $\mathcal{M}_{2}=\left\{f_{1}f_{2}g_{1}g_{2}\right\}$,
  3. $(iii)$
[…]
  4. $(iv)$
[…]
We sort the set of minimal generators of $\mathrm{in}_{<}(I_{G})$ as follows:
$\begin{split}&e_{1,1}e_{2,2}\prec e_{1,1}e_{2,3}\prec\cdots\prec e_{1,1}e_{m-1,2}\prec e_{2,1}e_{3,2}\prec e_{2,1}e_{4,2}\prec\cdots\prec e_{2,1}e_{m-1,2}\prec\cdots\\ &\prec e_{m-2,1}e_{m-1,2}\prec g_{1}f_{2}e_{1,1}\prec g_{1}f_{2}e_{2,1}\prec\cdots\prec g_{1}f_{2}e_{m-1,1}\prec f_{1}f_{2}g_{1}g_{2}\prec g_{2}f_{1}e_{m-1,2}\\ &\prec g_{2}f_{1}e_{m-2,2}\prec\cdots\prec g_{2}f_{1}e_{1,2}.\end{split}$
For brevity, we label each term in this sequence consecutively: $e_{1,1}e_{2,2}$ as $u_{1}$, $e_{1,1}e_{2,3}$ as $u_{2}$, and so forth, culminating with $g_{2}f_{1}e_{1,2}$ designated as $u_{r}$. Here, $r=\frac{m(m+1)}{2}$.
Lemma 4.3.
The monomial ideal $I=\mathrm{in}_{<}(I_{G})$ has regular quotients with respect to the above order. More precisely, the colon ideal $(u_{1},\ldots,u_{j-1}):u_{j}$ is generated by a regular sequence of length $k_{j}$ for $j=2,\ldots,\frac{m(m+1)}{2}$, where
$(k_{2},\ldots,k_{\frac{m(m+1)}{2}})=(\underbrace{{1,1}}_{2},\underbrace{2,2,2}_{3},\ldots,\underbrace{m-1,m-1,\ldots,m-1}_{m}).$
Proof.
Let $I_{j-1}=(u_{1},\ldots,u_{j-1})$ be the ideal generated by the first $j-1$ elements of the sequence. Now, we consider several cases based on the form of $u_{j}$:
1. If $u_{j}=e_{a,1}e_{b,2}$ for some $11\leq a<b\leq m-$ 1, then the colon ideal $I_{j-1}:u_{j}$ is given by
$I_{j-1}:u_{j}=(e_{1,1},e_{2,1},\ldots,e_{a-2,1},e_{a,2},e_{a+1,2},\ldots,e_{b-2,2}).$
2. If $u_{j}=g_{1}f_{2}e_{a,1}$ for some $11\leq a\leq m-$ 1, the colon ideal becomes
$I_{j-1}:u_{j}=(e_{1,1},e_{2,1},\ldots,e_{a-2,1},e_{a,2},e_{a+1,2},\ldots,e_{m-1,2}).$
3. If $u_{j}=f_{1}f_{2}g_{1}g_{2}$, the colon ideal simplifies to
$I_{j-1}:u_{j}=(e_{1,1},e_{2,1},\ldots,e_{m-1,1}).$
4. Finally, if $u_{j}=g_{2}f_{1}e_{a,2}$ for some $11\leq a\leq m-$ 1, the colon ideal is
$I_{j-1}:u_{j}=(g_{1}f_{2},e_{1,1},e_{2,1},\ldots,e_{a-2,1},e_{a,2},e_{a+1,2},\ldots,e_{m-1,2}).$
By enumerating the minimal generators of these colon ideals, the desired result follows.
Lemma 4.4.
Let $G$ be the graph as described above. Then there is a monomial order $<$ such that $\beta_{i}(\mathrm{in}_{<}(I_{G}))\leq(i+1)\binom{m+1}{i+2}$ for all $i\geq 0$. Thus,
$\beta_{i}(\mathbb{K}[G])\leq\beta_{i}(\mathbb{K}[E(G)]/\mathrm{in}_{<}(I_{G}))\leq i\binom{m+1}{i+1}.$
Proof.
Based on Lemma 4.3 and considering Formula (1), we obtain,
$\beta_{i}(\mathrm{in}_{<}(I_{G}))$ $\leq\sum_{k=1}^{m-1}(k+1)\binom{k}{i}+\binom{0}{i}=(i+1)\sum_{k=1}^{m-1}\binom{k+1}{i+1}+\binom{0}{i}$
$=(i+1)\sum_{k=0}^{m-1}\binom{k+1}{i+1}=(i+1)\binom{m+1}{i+2}.$
Here, the third-to-last equality follows from the identity $k\binom{n}{k}=n\binom{n-1}{k-1}$, and the last equality follows from the identity $\sum\limits_{j=0}^{n}\binom{j}{k}=\binom{n+1}{k+1}$ for $k\geq 0$. From this bound and Lemma 2.1, the second assertion is immediate.
Notice that a two-ear graph can have top-Betti induced subgraphs of three types: two-ear graphs, one-ear graphs, and complete bipartite graphs of the form $\mathbf{K}_{2,d}$ for some $d\geq 2$. Next, we will determine the elements in the top support $\mathcal{N}_{H}$ for the three types of graphs by considering the canonical module of their respective edge rings. The most complex scenario is when $H$ is a two-ear graph, for which we will provide a detailed proof. When $H$ is a one-ear graph or a complete bipartite graph, the proofs are similar with only minor differences, so we will only outline the proofs for those cases.
4.1. The composition of $\mathcal{N}_{\mathbf{G}_{m}}$
In this subsection, we use $G$ to represent $\mathbf{G}_{m}$ for short. It is easy to see that $|V(G)|=m+5$. It follows from Conventions 2.3 that:
${\mathbb{R}}^{V(G)}=\{\sum\limits_{i=1}^{m-1}{a_{i}{\mathbf{u}}_{i}}+\sum\limits_{i=1}^{2}({b_{i}{\mathbf{v}}_{i}+c_{i}{\mathbf{x}}_{i}+d_{i}{\mathbf{y}}_{i}})\mid a_{i},b_{i},c_{i},d_{i}\in{\mathbb{R}}\mbox{ for all }i\},$
where ${\mathbf{u}}_{i},{\mathbf{v}}_{i},{\mathbf{x}}_{i},{\mathbf{y}}_{i}$ are the unit vectors of ${\mathbb{R}}^{V(G)}$, each ${\mathbf{u}}_{i}$ (resp. ${\mathbf{v}}_{i},{\mathbf{x}}_{i},{\mathbf{y}}_{i}$) corresponds to $u_{i}$ (resp. $v_{i},x_{i},y_{i}$) for all $i$.
In what follows, we will construct $m$ integral vectors in ${\mathbb{R}}^{V(G)}$ and then show that they are minimal vectors of $\mathrm{relint}({\mathbb{R}}_{+}(G))\cap\mathcal{S}(G)$. The construction is as follows:
For $\ell=1,\ldots,m$, let
$\alpha_{\ell}:=\sum\limits_{i=1}^{m-1}{\mathbf{u}}_{i}+\sum\limits_{i=1}^{2}({\mathbf{x}}_{i}+{\mathbf{y}}_{i})+\ell{\mathbf{v}}_{1}+(m+1-\ell){\mathbf{v}}_{2}.$
We now verify that $\alpha_{\ell}\in\mathrm{relint}({\mathbb{R}}_{+}(G))$ for all $\ell=1,\ldots,m$.
We see the following:
  • The vertices of $G$ are all regular vertices.
  • An independent subset $T$ of $V(G)$ is fundamental if and only if $T$ is one of the following sets:
    1. $(i)$
      $\{u_{1},u_{2},\cdots,u_{m-1},x_{1},y_{2}\};$
    2. $(ii)$
      $\{u_{1},u_{2},\cdots,u_{m-1},x_{2},y_{1}\};$
    3. $(iii)$
      $\{v_{1},v_{2}\};$
    4. $(iv)$
      $\{v_{1},y_{1}\};$
    5. $(v)$
      $\{v_{2},x_{1}\}.$
Hence, it follows from ($\Delta_{2}$) that a vector of ${\mathbb{R}}^{V(G)}$ of the form:
$\sum\limits_{i=1}^{m-1}{a_{i}{\mathbf{u}}_{i}}+\sum\limits_{i=1}^{2}({b_{i}{\mathbf{v}}_{i}+c_{i}{\mathbf{x}}_{i}+d_{i}{\mathbf{y}}_{i}})$
belongs to ${\mathbb{R}}_{+}(G)$ if and only if the following inequalities are satisfied:
  1. (1)
    $a_{i}\geq 0$ for any $11\leq i\leq m-$ 1 ;
  2. (2)
    $b_{i},c_{i},d_{i}\geq 0$ for any $1\leq i\leq 2$ ;
  3. (3)
    $b_{1}+b_{2}+c_{2}+d_{1}\geq\sum\limits_{i=1}^{m-1}a_{i}+c_{1}+d_{2}$ ;
  4. (4)
    $b_{1}+b_{2}+c_{1}+d_{2}\geq\sum\limits_{i=1}^{m-1}a_{i}+c_{2}+d_{1}$ ;
  5. (5)
    $\sum\limits_{i=1}^{m-1}a_{i}+c_{1}+c_{2}+d_{1}+d_{2}\geq b_{1}+b_{2}$ ;
  6. (6)
    $\sum\limits_{i=1}^{m-1}a_{i}+b_{2}+c_{1}+c_{2}+d_{2}\geq b_{1}+d_{1}$ ;
  7. (7)
    $\sum\limits_{i=1}^{m-1}a_{i}+b_{1}+c_{2}+d_{1}+d_{2}\geq b_{2}+c_{1}$
It is straightforward to check that $\alpha_{\ell}$ satisfies these inequalities, with strict inequalities holding for each $\alpha_{\ell}$. It is obvious that the sum of all components of $\alpha_{\ell}$ is even, and according to [23, Proposition 3.4], we have $\alpha_{\ell}\in\mathrm{relint}({\mathbb{R}}_{+}(G))\cap\mathcal{S}(G)$.
Next, we show that $\alpha_{\ell}$ is a minimal vector in $\mathrm{relint}({\mathbb{R}}_{+}(G))\cap\mathcal{S}(G)$.
Suppose on the contrary that $\alpha_{\ell}=\alpha^{\prime}+\alpha^{\prime\prime}$ for some $\alpha^{\prime}\in\mathrm{relint}({\mathbb{R}}_{+}(G))\cap\mathcal{S}(G)$ and $\alpha^{\prime\prime}\in{\mathbb{R}}_{+}(G)\cap\mathcal{S}(G)\setminus\{{\bf 0}\}$. Write
$\alpha^{\prime}=\sum\limits_{i=1}^{m-1}{a^{\prime}_{i}{\mathbf{u}}_{i}}+\sum\limits_{i=1}^{2}({b^{\prime}_{i}{\mathbf{v}}_{i}+c^{\prime}_{i}{\mathbf{x}}_{i}+d^{\prime}_{i}{\mathbf{y}}_{i}}),$
$\alpha^{\prime\prime}=\sum\limits_{i=1}^{m-1}{a^{\prime\prime}_{i}{\mathbf{u}}_{i}}+\sum\limits_{i=1}^{2}({b^{\prime\prime}_{i}{\mathbf{v}}_{i}+c^{\prime\prime}_{i}{\mathbf{x}}_{i}+d^{\prime\prime}_{i}{\mathbf{y}}_{i}}).$
Note that $\mathrm{relint}({\mathbb{R}}_{+}(G))\cap\mathcal{S}(G)\subseteq{\mathbb{R}}_{+}(G)\cap\mathcal{S}(G)\subseteq\mathbb{Z}^{V(G)}$. In view of the inequalities (1) and (2), we see that $a^{\prime}_{i},c^{\prime}_{i},d^{\prime}_{i}\geq 1$ and $a^{\prime\prime}_{i},c^{\prime\prime}_{i},d^{\prime\prime}_{i}\geq 0$ for all $i,j$. Hence, $a^{\prime}_{i}=c^{\prime}_{i}=d^{\prime}_{i}=1$ and $a^{\prime\prime}_{i}=c^{\prime\prime}_{i}=d^{\prime\prime}_{i}=0$ for all possible $i,j$. Because of (3) ,we have $b^{\prime}_{1}+b^{\prime}_{2}\geq m$. Hence $b^{\prime\prime}_{1}+b^{\prime\prime}_{2}\leq 1$. Because of (6) and (7), we have $b^{\prime\prime}_{1}=b^{\prime\prime}_{2}$. But from (2) we know $b^{\prime\prime}_{1},b^{\prime\prime}_{2}\geq 0$, so we have $b^{\prime\prime}_{1}=b^{\prime\prime}_{2}=0$.
Hence, $\alpha^{\prime\prime}=0.$ This is a contradiction, which shows that $\alpha_{\ell}$ is a minimal vector in $\mathrm{relint}({\mathbb{R}}_{+}(G))\cap\mathcal{S}(G)$ for $\ell=1,\ldots,m$.
Actually, $\alpha_{\ell}$ for $\ell=1,\ldots,m$ are all minimal vector in $\mathrm{relint}({\mathbb{R}}_{+}(G))\cap\mathcal{S}(G)$, as shown below.
Lemma 4.5.
Let $G$ represent the graph $\mathbf{G}_{m}$ with the labeling as shown in Figure 1. Then, we have
$\mathrm{pdim}(\mathbb{K}[G])=\mathrm{pdim}(\mathbb{K}[E(G)]/\mathrm{in}_{<}(I_{G}))=m.$
Furthermore,
$\mathcal{N}_{G}=\{\Theta_{G}x_{1}y_{1}v_{1}^{\ell}v_{2}^{m+1-\ell}\mid\ell=1,\ldots,m\}.$
Proof.
The first statement is a consequence of Lemma 2.9 combined with [8, Corollary 2.7].
In addition, since $\mathrm{pdim}(\mathbb{K}[G])=m$, according to Lemma 4.4, the top-Betti number of $\mathbb{K}[G]$ is at most $m$. This implies that the set $\mathcal{N}_{G}$ contains no more than $m$ elements. Conversely, as previously shown, $\mathcal{N}_{G}$ contains at least $m$ elements. The desired result is then obtained by the fact that $D_{G}=\Theta_{G}^{2}(x_{1}y_{1})^{3}(v_{1}v_{2})^{m+1}$ along with Lemma 2.7.
4.2. The composition of $\mathcal{N}_{\mathbf{P}_{m}}$.
Let $G^{\prime}$ (resp. $G^{\prime\prime}$) denote the induced subgraph $G\setminus x_{2}$ (resp. $G\setminus y_{2}$) of $G$. According to Definition 4.1, both $G^{\prime}$ and $G^{\prime\prime}$ are isomorphic to the one-ear graph $\mathbf{P}_{m}$.
It follows from Conventions 2.3 that:
${\mathbb{R}}^{V(G^{\prime})}=\{\sum\limits_{i=1}^{m-1}{a_{i}{\mathbf{u}}_{i}}+\sum\limits_{i=1}^{2}({b_{i}{\mathbf{v}}_{i}+d_{i}{\mathbf{y}}_{i}})+c_{1}{\mathbf{x}}_{1}\mid a_{i},b_{i},c_{1},d_{i}\in{\mathbb{R}}\mbox{for all }i\},$
here, ${\mathbf{u}}_{i},{\mathbf{v}}_{i},{\mathbf{x}}_{1},{\mathbf{y}}_{i}$ correspond the vertices of $G^{\prime}$ in a natural way. We could show the following vectors employing the same process as in the case of a two-ear graph:
$\alpha_{\ell}:=\sum\limits_{i=1}^{m-1}{\mathbf{u}}_{i}+{\mathbf{x}}_{1}+2{\mathbf{y}}_{1}+{\mathbf{y}}_{2}+\ell{\mathbf{v}}_{1}+(m-\ell){\mathbf{v}}_{2},\ \ell=1,\ldots,m-1$
are the minimal vectors of $\mathrm{relint}({\mathbb{R}}_{+}(G^{\prime}))\cap\mathcal{S}(G^{\prime})$.
It is clear that $D_{G^{\prime}}=\Theta_{G^{\prime}}^{2}y_{1}^{3}v_{1}^{m}v_{2}^{m+1}$. The following result is now immediate from Lemma 2.9 and Lemma 2.7.
Lemma 4.6.
Assume that $G^{\prime}$ is labelled as in Figure 1. Then
$\mathrm{pdim}(\mathbb{K}[G^{\prime}])=m-1\mbox{ and }\{\Theta_{G^{\prime}}y_{1}v_{1}^{\ell}v_{2}^{m-\ell}\mid\ell=1,\ldots,m-1\}\subseteq\mathcal{N}_{G^{\prime}}.$
4.3. The composition of $\mathcal{N}_{\mathbf{K}_{2,d}}$.
For certain values of $d\geq 2$, a complete bipartite graph of the form $\mathbf{K}_{2,d}$ may also be an induced subgraph of graph $G$.
It is easy to see that $|V(\mathbf{K}_{2,d})|=d+2$. Assume that $\mathbf{K}_{2,d}$ has two parts $\{v_{1},v_{2}\}$ and $\{u_{1},\ldots,u_{d}\}$. It follows from Conventions 2.3 that:
${\mathbb{R}}^{V(\mathbf{K}_{2,d})}=\{\sum\limits_{i=1}^{d}{a_{i}{\mathbf{u}}_{i}}+{b_{1}{\mathbf{v}}_{1}}+{b_{2}{\mathbf{v}}_{2}}\mid a_{i},b_{i}\in{\mathbb{R}}\mbox{ for all }i\},$
where ${\mathbf{u}}_{i},{\mathbf{v}}_{i}$ are the unit vectors of ${\mathbb{R}}^{V(\mathbf{K}_{2,d})}$, each ${\mathbf{u}}_{i}$ (resp. ${\mathbf{v}}_{i}$) corresponds to $u_{i}$ (resp. $v_{i}$) for all $i$.
In what follows, we will construct $(d-1)$ integral vectors in ${\mathbb{R}}^{V(\mathbf{K}_{2,d})}$ and then show that they are minimal vectors of $\mathrm{relint}({\mathbb{R}}_{+}(\mathbf{K}_{2,d}))\cap\mathcal{S}(\mathbf{K}_{2,d})$. The construction is as follows:
For $\ell=1,\ldots,d-1$, let
$\alpha_{\ell}:=\sum\limits_{i=1}^{d}{\mathbf{u}}_{i}+\ell{\mathbf{v}}_{1}+(d-\ell){\mathbf{v}}_{2}.$
According to Lemma 2.5, all the facets of ${\mathbb{R}}_{+}(\mathbf{K}_{2,d})$ are exactly $H_{\{v_{1}\}}\cap{\mathbb{R}}_{+}(\mathbf{K}_{2,d})$ and $H_{\{v_{2}\}}\cap{\mathbb{R}}_{+}(\mathbf{K}_{2,d})$.
Hence, it follows from ($\Delta_{1}$) that a vector of ${\mathbb{R}}^{V(\mathbf{K}_{2,d})}$ of the form:
$\sum\limits_{i=1}^{d}{a_{i}{\mathbf{u}}_{i}}+b_{1}{\mathbf{v}}_{1}+b_{2}{\mathbf{v}}_{2}$
belongs to ${\mathbb{R}}_{+}(\mathbf{K}_{2,d})$ if and only if the following formulas are satisfied:
  1. (1)
    $a_{i}\geq 0$ for any $1\leq i\leq d$ ;
  2. (2)
    $b_{i}\geq 0$ for any $i=1,2$ ;
  3. (3)
    $\sum\limits_{i=1}^{d}a_{i}\geq b_{1}$ ;
  4. (4)
    $\sum\limits_{i=1}^{d}a_{i}\geq b_{2}$
It is routine to prove that $\alpha_{\ell}$ is a minimal vector in $\mathrm{relint}({\mathbb{R}}_{+}(G))\cap\mathcal{S}(G)$ for $\ell=1,\ldots,d-1$. Denote $H=\mathbf{K}_{2,d}$. Then it is clear that $D_{H}=\Theta_{H}^{2}v_{1}^{d}v_{2}^{d}$. Therefore, the subsequent result follows directly from Lemma 2.9 and Lemma 2.7.
Lemma 4.7.
Denote $H=\mathbf{K}_{2,d}$, then
$\mathrm{pdim}(\mathbb{K}[H])=d-1\mbox{ and }\{\Theta_{H}v_{1}^{\ell}v_{2}^{d-\ell}\mid\ell=1,\ldots,d-1\}\subseteq\mathcal{N}_{H}.$
Thus far, we have computed the projective dimension of the edge rings for the three types of graphs under consideration. Moreover, we have determined the top support of a two-ear graph. As for the one-ear graph and the complete bipartite graph $\mathbf{K}_{2,d}$, we have respectively obtained two particular subsets of their top supports. We are now ready to present the first main result of this section.
Theorem 4.8.
Let $G$ be a two-ear graph of type $m$. Then, for any $0\leq i\leq m$, we have
$\beta_{i}(\mathbb{K}[G])=\beta_{i}(\mathbb{K}[E(G)]/\mathrm{in}_{<}(I_{G})).$
Proof.
From Lemma 4.5, it can be seen that $\mathrm{pdim}(\mathbb{K}[G])=\mathrm{pdim}(\mathbb{K}[E(G)]/\mathrm{in}_{<}(I_{G}))=m$. We only need to prove the two Betti numbers are identical.
If $i=0$, then $\beta_{i}(\mathbb{K}[G])=\beta_{i}(\mathbb{K}[E(G)]/\mathrm{in}_{<}(I_{G}))=1.$ Suppose now that $1<i\leq m$. We now identify all top-Betti inducedsubgraphs of $G$ whose edge rings have a projective dimension of $i$. Note that $\mathrm{pdim}(\mathbb{K}[\mathbf{G}_{i}])=\mathrm{pdim}(\mathbb{K}[\mathbf{P}_{i+1}])=\mathrm{pdim}(\mathbb{K}[\mathbf{K}_{2,i+1}])=i$. Observe that $G$ has $\binom{m-1}{i-1}$ induced subgraphs that are isomorphic to $\mathbf{G}_{i}$, 2$\binom{m-1}{i}$ induced subgraphs that are isomorphic to $\mathbf{P}_{i+1}$, and $\binom{m-1}{i+1}$ induced subgraphs which are isomorphic to $\mathbf{K}_{2,i+1}$. Hence, there are at least $r$ induced subgraphs $H$ of $G$ such that $\mathrm{pdim}(\mathbb{K}[H])=i$. Here, $r$ is given by
$r=\binom{m-1}{i-1}+2\binom{m-1}{i}+\binom{m-1}{i+1}=\binom{m+1}{i+1}.$
Let $\mathcal{N}_{i}(G)$ denote the union of all $\mathcal{N}_{H}$ such that $H$ is a top-Betti inducedsubgraph of $G$ satisfying $\mathrm{pdim}(\mathbb{K}[H])=i$.
Observe that if $H$ is isomorphic to $\mathbf{G}_{i}$, then according to Lemma 4.5, $\mathcal{N}_{H}$ comprises $i$ monomials. If $H$ is isomorphic to either $\mathbf{P}_{i+1}$ or $\mathbf{K}_{2,i+1}$, then the corresponding lemma (Lemma 4.6 for $\mathbf{P}_{i+1}$ and Lemma 4.7 for $\mathbf{K}_{2,i+1}$) ensures that $\mathcal{N}_{H}$ contains at least $i$ monomials.
Hence,
$|\mathcal{N}_{i}(G)|\geq ir=i\binom{m+1}{i+1}.$
This implies
$i\binom{m+1}{i+1}\leq\sum_{\alpha\in\mathcal{N}_{i}(G)}\beta_{i,\alpha}(\mathbb{K}[G])\leq\beta_{i}(\mathbb{K}[G])\leq\beta_{i}(\mathbb{K}[E(G)]/\mathrm{in}_{<}(I_{G}))\leq i\binom{m+1}{i+1},$
as desired.
From the proof of Proposition 4.8, we can immediately draw the following two corollaries. These represent an improvement over Lemmas 4.6 and 4.7, enabling us to enumerate all elements of $\mathcal{N}_{H}$ for all one-ear graphs and complete bipartite graphs $H$.
Corollary 4.9.
Denote by $H$ the graph $G^{\prime}$ described above, then
$\mathcal{N}_{H}=\{\Theta_{H}y_{1}v_{1}^{\ell}v_{2}^{m-\ell}\mid\ell=1,\ldots,m-1\}.$
Corollary 4.10.
Denote by $H$ the graph $\mathbf{K}_{2,d}$ described above, then
$\mathcal{N}_{H}=\{\Theta_{H}v_{1}^{\ell}v_{2}^{d-\ell}\mid\ell=1,\ldots,d-1\}.$
We now present the second main result of this section. As in the proof of Proposition 4.8, $\mathcal{N}_{i}(G)$ denotes the disjoint union of all $\mathcal{N}_{H}$ such that $H$ is a top-Betti induced subgraph of $G$ satisfying $\mathrm{pdim}(\mathbb{K}[H])=i$.
Theorem 4.11.
Let $G$ be a two-ear graph of type $m$. Then for any monomial $\alpha$ in $\mathbb{K}[G]$ and for any integer $i\geq 1$, we have
$\beta_{i,\alpha}(\mathbb{K}[G])=\left\{\begin{array}[]{ll}1,&\hbox{$\alpha\in\mathcal{N}_{i}(G)$;}\\ 0,&\hbox{otherwise.}\end{array}\right.$
In particular, there exists a square-free initial ideal $\mathrm{in}_{<}(I_{G})$ of $I_{G}$ such that
$\beta_{i,j}(\mathrm{in}_{<}(I_{G}))=\beta_{i,j}(I_{G})=|\{\alpha\in\mathcal{N}_{i+1}(G)\mid|\mathrm{deg}(\alpha)|=2j\}|$ for all $i,j$.
5. Basics for compact graphs
The primary objective of this section is to determine the elements of the top support $\mathcal{N}_{H}$ for every compact graph $H$, thereby facilitating the application of the induced-subgraph method for computing the Betti numbers of their edge rings in the subsequent section. To accomplish this, we revisit the precise definition and classification of compact graphs as presented in [25].
Definition 5.1.
Let $G$ be a connected simple graph where every vertex has a degree greater than one. We call $G$ to be a compact graph if it does not contain any even cycles and satisfies the odd-cycle condition.
We call a vertex of a compact graph to be a big vertex if it has a degree greater than 2. According to [25, Proposition 3.3], there are at most three big vertices in a compact graph. Therefore, compact graphs containing more than one cycle can be categorized into three classes, each determined by the number of big vertices. Specifically, we say a compact graph falls into type $i$ if it possesses $i$ big vertices for $i=1,2,3.$
We present a complete classification of compact graphs as follows, which is given in [25, Theorem 2.12]. A compact graph of type 1 is a finite collection of odd cycles that share a vertex.
A compact graph of type 2 consists of two disjoint compact graphs of type 1, where the two big vertices are connected either by an edge or by an edge as well as a path of even length.
A compact graph of type 3 consists of three disjoint compact graphs of type 1, where every pair of big vertices is connected by an edge.
Suppose $\underline{p}=(p_{1},\ldots,p_{m})$, $\underline{q}=(q_{1},\ldots,q_{n})$ and $\underline{r}=(r_{1},\ldots,r_{k})$ are positive integral vectors with dimensions $m,n$ and $k$ respectively. We denote a compact graph of type 1, where the odd cycles have lengths $12p_{1}+1,\ldots,2p_{m}+$ 1 respectively, as $\mathbf{A}_{\underline{p}}$ or $\mathbf{A}_{(p_{1},\ldots,p_{m})}$.
By $\mathbf{B}_{\underline{p}:\underline{q}}^{0}$ we mean a compact graph of type 2 where the two disjoint compact graphs of type 1 that compose it are $\mathbf{A}_{\underline{p}}$ and $\mathbf{A}_{\underline{q}}$ and where the two big vertices are connected by an edge. Furthermore, if $s>0$ is an even number, then $\mathbf{B}_{\underline{p};\underline{q}}^{s}$ represents the graph obtained by appending a path of length $s$ connecting two big vertices to $\mathbf{B}_{\underline{p}:\underline{q}}^{0}$.
A compact graph of type 3 is denoted by $\mathbf{C}_{\underline{p}:\underline{q}:\underline{r}}$ if the three disjoint compact graphs of type 1 that make up it are $\mathbf{A}_{\underline{p}}$, $\mathbf{A}_{\underline{q}}$ and $\mathbf{A}_{\underline{r}}$ respectively.
By Proposition 3.1, it is easy to see that all compact graphs are top-Betti graphs, which, according to Proposition 3.5, implies that they are also second-Betti graphs. We shall revisit the findings from [25] pertaining to the minimal generators of the edge cone of compact graphs, and subsequently establish the compositions of top supports of three types of compact graphs, respectively.
5.1. Type one
Given positive integers $m\geq 2$ and $p_{1},\ldots,p_{m}$, we use $A$ to denote the graph $\mathbf{A}_{(p_{1},\ldots,p_{m})}$ for short. We may assume that $A$ has vertex set
$V(A)=\{u\}\cup\{u_{i,j}\mid 1\leq i\leq m,1\leq j\leq 2p_{i}\}$
and edge set
[…]
$\cup\{\{u,u_{i,1}\},\{u,u_{i,2p_{i}}\}\mid 1\leq i\leq m\}.$
Recall that an integral vector in $\mathrm{relint}({\mathbb{R}}_{+}(G))$ is called minimal if it cannot written as the sum of a vector in $\mathrm{relint}({\mathbb{R}}_{+}(G))\cap\mathcal{S}(G)$ and a nonzero vector of ${\mathbb{R}}_{+}(G)\cap\mathcal{S}(G)$. We showed in [25, Section 6] that the following vectors
$\alpha_{\ell}:=\sum\limits_{i=1}^{m}\sum\limits_{j=1}^{2p_{i}}{\mathbf{u}}_{i,j}+2\ell{\mathbf{u}},\ \ell=1,\ldots,m-1$
are all the minimal vectors of $\mathrm{relint}({\mathbb{R}}_{+}(A))\cap\mathcal{S}(A)$.
It is clear that $D_{A}=\Theta_{A}^{2}u^{2m}$. Therefore, the subsequent result follows directly from Lemma 2.7.
Lemma 5.2.
Let $A$ be a compact graph of type 1 labelled as previously. Then
$\mathcal{N}_{A}=\{\Theta_{A}u^{2\ell}\mid\ell=1,\ldots,m-1\}.$
5.2. type two
Assume that $n,m$ and $p_{1},\ldots,p_{m},q_{1},\ldots,q_{n}$ are given positive integers. Let $s\geq 0$ be an even number. We use $B$ denote the graph $\mathbf{B}^{s}_{(p_{1},\ldots,p_{m}):(q_{1},\ldots,q_{n})}$ for short. We may assume that $B$ has vertex set
$V(B)=$$\{u,v\}\cup\{w_{1},\ldots,w_{s-1}\}$
$\cup\{u_{i,j}\mid 1\leq i\leq m,1\leq j\leq 2p_{i}\}\cup\{v_{i,j}\mid 1\leq i\leq n,1\leq j\leq 2q_{i}\}$
and edge set
[…]
$\cup\{\{u,u_{i,1}\},\{u,u_{i,2p_{i}}\}\mid 1\leq i\leq m\}$
$\cup\{\{u,v\},\{u,w_{1}\},\{v,w_{s-1}\}\}\cup\{\{w_{i},w_{i+1}\}\mid 1\leq i\leq s-2\}$
$\cup\{\{v_{i,j},v_{i,j+1}\}\mid 1\leq i\leq n,1\leq j\leq 2q_{i}-1\}$
$\cup\{\{v,v_{i,1}\},\{v,v_{i,2q_{i}}\}\mid 1\leq i\leq n\}.$
Let $B^{0}$ and $B^{s}$ denote the compact graph $\mathbf{B}_{\underline{p}:\underline{q}}^{0}$ and $\mathbf{B}_{\underline{p}:\underline{q}}^{s}$ respectively, where $s\geq 2$ is an even number. We may write by Conventions 2.3
${\mathbb{R}}^{V(B^{0})}=\{\sum\limits_{i=1}^{m}\sum\limits_{j=1}^{2p_{i}}a_{i,j}{\mathbf{u}}_{i,j}+\sum\limits_{i=1}^{m}\sum\limits_{j=1}^{2p_{i}}b_{i,j}{\mathbf{v}}_{i,j}+a{\mathbf{u}}+b{\mathbf{v}}\mid\mbox{ all }a_{i,j},b_{i,j},a,b\in{\mathbb{R}}\}.$
We showed in [25, Section 5] that the following vectors
$\alpha_{\ell}:=\sum\limits_{i=1}^{m}\sum\limits_{j=1}^{2p_{i}}{\mathbf{u}}_{i,j}+\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{2q_{i}}{\mathbf{v}}_{i,j}+{\mathbf{v}}+(2\ell+1){\mathbf{u}},\quad\ell=0,\ldots,m-1$
and
$\beta_{\ell}:=\sum\limits_{i=1}^{m}\sum\limits_{j=1}^{2p_{i}}{\mathbf{u}}_{i,j}+\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{2q_{i}}{\mathbf{v}}_{i,j}+{\mathbf{u}}+(2\ell+1){\mathbf{v}},\quad\ell=1,\ldots,n-1$
are all the minimal vectors of $\mathrm{relint}({\mathbb{R}}_{+}(B^{0}))\cap\mathcal{S}(B^{0})$.
Similarly we may write
${\mathbb{R}}^{V(B^{s})}=\{\sum\limits_{i=1}^{m}\sum\limits_{j=1}^{2p_{i}}a_{i,j}{\mathbf{u}}_{i,j}+\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{2q_{i}}b_{i,j}{\mathbf{v}}_{i,j}+\sum\limits_{i=1}^{s-1}c_{i}{\mathbf{w}}_{i}+a{\mathbf{u}}+b{\mathbf{v}}\mid\mbox{ all }a_{i,j},b_{i,j},c_{i},a,b\in{\mathbb{R}}\}.$
and showed in [25, Section 6] that the following vectors
$\alpha_{\ell}:=\sum\limits_{i=1}^{m}\sum\limits_{j=1}^{2p_{i}}{\mathbf{u}}_{i,j}+\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{2q_{i}}{\mathbf{v}}_{i,j}+\sum\limits_{i=1}^{s-1}{\mathbf{w}}_{i}+{\mathbf{v}}+2\ell{\mathbf{u}},\ell=1,\ldots,m$
and
$\beta_{\ell}:=\sum\limits_{i=1}^{m}\sum\limits_{j=1}^{2p_{i}}{\mathbf{u}}_{i,j}+\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{2q_{i}}{\mathbf{v}}_{i,j}+\sum\limits_{i=1}^{s-1}{\mathbf{w}}_{i}+{\mathbf{u}}+2\ell{\mathbf{v}},\ell=1,\ldots,n$
are all the minimal vectors of $\mathrm{relint}({\mathbb{R}}_{+}(B^{s}))\cap\mathcal{S}(B^{s})$.
Lemma 5.3.
For each compact graph $H$ of type 2,
  1. $(1)$
    If $H=\mathbf{B}_{\underline{p}:\underline{q}}^{0}$, then
    $\mathcal{N}_{H}=\{\Theta_{H}u^{2m}v^{2\ell}\mid\ell=1,\ldots,n-1\}\sqcup\{\Theta_{H}u^{2\ell}v^{2n}\mid\ell=1,\ldots,m\}.$
  2. $(2)$
    If $H=\mathbf{B}_{\underline{p}:\underline{q}}^{s}$, then
    $\mathcal{N}_{H}=\{\Theta_{H}u^{2m+1}v^{2\ell}\mid\ell=1,\ldots,n\}\sqcup\{\Theta_{H}u^{2\ell}v^{2n+1}\mid\ell=1,\ldots,m\}.$
Proof.
If $H=\mathbf{B}_{\underline{p}:\underline{q}}^{0}$, then $D_{H}=\Theta_{H}^{2}u^{2m+1}v^{2n+1}$, and if $H=\mathbf{B}_{\underline{p}:\underline{q}}^{s}$, the expression for $D_{H}$ becomes $\Theta_{H}^{2}u^{2m+2}v^{2n+2}$. Given the minimal generators within the relative interior of the edge cones of $H$ as enumerated previously, the result is straightforwardly deduced.
Analogously in Lemma 5.3, the notation $u,v$ utilized in the expression of $\mathcal{N}_{H}$ represents the two big vertices of the graph $H$. Nevertheless, upon considering $H$ as an induced subgraph of a compact graph of type 3, there arises a possibility that the two big vertices of $H$ might alternately be denoted as $w,u$ or $w,v$. In such cases, it becomes necessary to substitute $w,u$ with either $v$ or $w,v$ accordingly.
5.3. type three
Given positive integers $m,n,k$, as well as the positive integral tuples $\underline{p}=(p_{1},\ldots,p_{m})$, $\underline{q}=(q_{1},\ldots,q_{n})$ and $\underline{r}=(r_{1},\ldots,r_{k})$, we denote the graph $\mathbf{C}_{\underline{p}:\underline{q}:\underline{r}}$ as $C$ for brevity. Here, $p_{i},q_{i},r_{i}$ are all positive integers. By definition, we may assume that $C$ has vertex set
$V(C)=$$\{\{u,v,w\}\cup\{u_{i,j}\mid 1\leq i\leq m,1\leq j\leq 2p_{i}\}$
$\cup\{v_{i,j}\mid 1\leq i\leq n,1\leq j\leq 2q_{i}\}\cup\{w_{i,j}\mid 1\leq i\leq k,1\leq j\leq 2r_{i}\},$
and edge set
$E(C)=$$\{\{u_{i,j},u_{i,j+1}\}\mid 1\leq i\leq m,1\leq j\leq 2p_{i}-1\}$
$\cup\{\{u,u_{i,1}\},\{u,u_{i,2p_{i}}\}\mid 1\leq i\leq m\}$
$\cup\{\{v_{i,j},v_{i,j+1}\}\mid 1\leq i\leq n,1\leq j\leq 2q_{i}-1\}$
$\cup\{\{v,v_{i,1}\},\{v,v_{i,2q_{i}}\}\mid 1\leq i\leq n\}$
$\cup\{\{w_{i,j},w_{i,j+1}\}\mid 1\leq i\leq k,1\leq j\leq 2r_{i}-1\}$
$\cup\{\{w,w_{i,1}\},\{w,w_{i,2r_{i}}\}\mid 1\leq i\leq k\}$
$\cup\{\{u,v\},\{v,w\},\{w,u\}\}.$
It follows from Conventions 2.3 that
${\mathbb{R}}^{V(C)}=\{\sum\limits_{i=1}^{m}\sum\limits_{j=1}^{2p_{i}}{a_{i,j}{\mathbf{u}}_{i,j}}+a{\mathbf{u}}+\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{2q_{i}}{b_{i,j}{\mathbf{v}}_{i,j}}+b{\mathbf{v}}+\sum\limits_{i=1}^{k}\sum\limits_{j=1}^{2r_{i}}{c_{i,j}{\mathbf{w}}_{i,j}}+c{\mathbf{w}}\mid$
$a_{i,j},a,b_{i,j},b,c_{i,j},c\in{\mathbb{R}}\mbox{ for all }i,j\}.$
According to [25, Section 6], we have the following result.
Lemma 5.4.
The minimal vectors of $\mathrm{relint}({\mathbb{R}}_{+}(C))\cap\mathcal{S}(C)$ are following $m+n+k$ vectors: $\alpha_{\ell}$ for $\ell=1,\ldots,m,$ $\beta_{\ell}$ for ${\ell}=1,\ldots,n$ and $\gamma_{\ell}$ for ${\ell}=1,\ldots,k$. Here, For $\ell=1,\ldots,m$,
$\alpha_{\ell}:=\sum\limits_{i=1}^{m}\sum\limits_{j=1}^{2p_{i}}{\mathbf{u}}_{i,j}+\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{2q_{i}}{\mathbf{v}}_{i,j}+\sum\limits_{i=1}^{k}\sum\limits_{j=1}^{2r_{i}}{\mathbf{w}}_{i,j}+{\mathbf{v}}+{\mathbf{w}}+2{\ell}{\mathbf{u}}.$
For ${\ell}=1,\ldots,n$,
$\beta_{\ell}=\sum\limits_{i=1}^{m}\sum\limits_{j=1}^{2p_{i}}{\mathbf{u}}_{i,j}+\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{2q_{i}}{\mathbf{v}}_{i,j}+\sum\limits_{i=1}^{k}\sum\limits_{j=1}^{2r_{i}}{\mathbf{w}}_{i,j}+{\mathbf{w}}+{\mathbf{u}}+2{\ell}{\mathbf{v}}.$
For ${\ell}=1,\ldots,k$,
$\gamma_{\ell}=\sum\limits_{i=1}^{m}\sum\limits_{j=1}^{2p_{i}}{\mathbf{u}}_{i,j}+\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{2q_{i}}{\mathbf{v}}_{i,j}+\sum\limits_{i=1}^{k}\sum\limits_{j=1}^{2r_{i}}{\mathbf{w}}_{i,j}+{\mathbf{u}}+{\mathbf{v}}+2{\ell}{\mathbf{w}}.$
In other words, the canonical module $\omega_{\mathbb{K}[C]}$ of $\mathbb{K}[C]$ is a monomial ideal of $\mathbb{K}[C]$ generated by
$\Theta_{C}u^{2}vw,\Theta_{C}u^{4}vw,\ldots,\Theta_{C}u^{2m}vw,\quad\Theta_{C}uv^{2}w,\Theta_{C}uv^{4}w,\ldots,\Theta_{C}uv^{2n}w$
and
$\quad\Theta_{C}uvw^{2},\Theta_{C}uvw^{4},\ldots,\Theta_{C}uvw^{2k}.$
The computation of $D_{C}$ is straightforward, resulting in $D_{C}=\Theta_{C}u^{2m+2}v^{2n+2}w^{2k+2}$. From this equality and applying Formula (1) in Subsection 2.5, we obtain the following result immediately.
Lemma 5.5.
If $H=\mathbf{C}_{\underline{p}:\underline{q}:\underline{r}}$, then
$\mathcal{N}_{H}=$$\{\Theta_{H}u^{2\ell}v^{2n+1}w^{2k+1}\mid\ell=1,\ldots,m\}\sqcup\{\Theta_{H}u^{2m+1}v^{2\ell}w^{2k+1}\mid\ell=1,\ldots,n\}$
$\sqcup\{\Theta_{H}u^{2m+1}v^{2n+1}w^{2\ell}\mid\ell=1,\ldots,k\}.$
Till now we have computed the top supports of all compact graphs. Considering the composition of these sets, we can conclude that, intriguingly, for any compact graph $H$, $\mathcal{N}_{H}$ contains precisely $\mathfrak{t}(H)-1$ elements, where $\mathfrak{t}(H)$ denotes the number of minimal cycles in graph $H$. Moreover, we can conclude that for any compact graph $H$, $\Theta_{H}$ divides $\alpha$ for all $\alpha\in\mathcal{N}_{H}$. This provides another perspective to confirm that $H$ is a top-Betti graph.
6. Betti numbers of compact graphs
In this section, we employ the induced-subgraph method to derive the multi-graded Betti numbers of the edge rings associated with compact graphs. Our objective is achieved by proving that the Betti numbers of the toric ideals of compact graphs coincide with those of a particular square-free initial ideal.
Let $\mathfrak{t}(G)$ represent the count of minimal cycles in graph $G$
Lemma 6.1.
Let $G$ be a compact graph. Then
$\mathrm{(1)}$ The projective dimension of $\mathbb{K}[G]$ is equal to $\mathfrak{t}(H)-1$;
$\mathrm{(2)}$ there is a monomial order $<$ such that $\beta_{i}(\mathrm{in}_{<}(I_{G}))=(i+1)\binom{\mathfrak{t}(G)}{i+2}$ for all $i\geq 0$. Thus, for all $i\geq 1$, one has
$\beta_{i}(\mathbb{K}[G])\leq\beta_{i}(\mathbb{K}[E(G)]/\mathrm{in}_{<}(I_{G}))=i\binom{\mathfrak{t}(G)}{i+1}.$
Proof.
The conclusion (1) follows from [25, Corollary 5.1] and the conclusion of (2) is immediate from [25, Theorem 5.1] as well as Lemma 2.1.
Proposition 6.2.
Let $G$ be a compact graph of either type 1 or type 2. Then, for any $10\leq i\leq\mathfrak{t}(G)-$ 1, we have
$\beta_{i}(\mathbb{K}[G])=\beta_{i}(\mathbb{K}[E(G)]/\mathrm{in}_{<}(I_{G})).$
Proof.
If $i=0$, then $\beta_{i}(\mathbb{K}[G])=\beta_{i}(\mathbb{K}[E(G)]/\mathrm{in}_{<}(I_{G}))=1.$ Fix $11\leq i\leq\mathfrak{t}(G)-$ 1. Since there are $\mathfrak{t}(G)$ minimal cycles, there are $\binom{\mathfrak{t}(G)}{i+1}$ compact induced subgraphs of $G$ that contain exactly $i+1$ minimal cycles of $G$. These induced subgraphs are of either type one or type two, and are denoted by $H_{1},\ldots,H_{r}$, respectively. Here, $r=\binom{\mathfrak{t}(G)}{i+1}$. Note that $\beta_{i,\alpha}(\mathbb{K}[H_{\ell}])=1$ for any $\alpha\in\mathcal{N}_{H_{\ell}}$ by Lemmas 5.2 and 5.3, it follows that $\beta_{i,\alpha}(\mathbb{K}[G])=1$ for any $\alpha\in\bigcup_{\ell=1}^{r}\mathcal{N}_{H_{\ell}}$ in view of Lemma 2.8. Put $\mathcal{N}_{i}(G)=\bigcup_{\ell=1}^{r}\mathcal{N}_{H_{\ell}}$.
Note that $\mathrm{supp}(\alpha)=V(H_{k})$ for any $\alpha\in\mathcal{N}_{H_{k}}$, we have $\mathcal{N}_{H_{1}},\ldots,\mathcal{N}_{H_{r}}$ are pairwise disjoint. Hence,
$|\mathcal{N}_{i}(G)|=\sum_{\ell=1}^{r}|\mathcal{N}_{H_{\ell}}|=ir=i\binom{\mathfrak{t}(G)}{i+1}.$
This implies
$i\binom{\mathfrak{t}(G)}{i+1}=\sum_{\alpha\in\mathcal{N}_{i}(G)}\beta_{i,\alpha}(\mathbb{K}[G])\leq\beta_{i}(\mathbb{K}[G])\leq\beta_{i}(\mathbb{K}[E(G)]/\mathrm{in}_{<}(I_{G}))=i\binom{\mathfrak{t}(G)}{i+1},$
completing the proof.
From the proof of Proposition 6.2, it is evident that
$\beta_{i}(\mathbb{K}[G])=\sum_{\alpha\in\mathcal{N}_{i}(G)}\beta_{i,\alpha}(\mathbb{K}[G]).$
In particular, if $\beta_{i,\alpha}(\mathbb{K}[G])\neq 0$, then $\alpha\in\mathcal{N}_{i}(G)$. Therefore, we can derive the multi-graded and graded Betti numbers of $\mathbb{K}[G]$ as follows.
Corollary 6.3.
Let $G$ be a compact graph of either type 1 or type 2. For any monomial $\alpha$ in $\mathbb{K}[G]$, we have
$\beta_{i,\alpha}(\mathbb{K}[G])=\left\{\begin{array}[]{ll}1,&\hbox{$\alpha\in\mathcal{N}_{i}(G)$;}\\ 0,&\hbox{otherwise.}\end{array}\right.$
In particular, $\beta_{i,j}(\mathbb{K}[G])$ is equal to the number of the set $\{\alpha\in\mathcal{N}_{i}(G)\mid|\mathrm{deg}(\alpha)|=2j\}.$
Corollary 6.3 implies that if $G$ is a compact graph of type 1 or type 2 then $G$ is both a top-Betti graph and a second-Betti graph, but not a strict second-Betti graph.
We now turn to compact graphs of type 3. In the case when $G$ is of type one or type two, there are exactly $\binom{\mathfrak{t}(G)}{i+1}$ distinct compact induced subgraphs of $G$ which contains exactly $i+1$ minimal cycles for all $11\leq i\leq\mathfrak{t}(G)-$ 1. This is the reason why the proof of Theorem 6.2 holds. But this is not the case when considering compact graphs of type 3. Let us see the following example.
Example 6.4.
Assume that $G=\mathbf{C}_{(1):(1):(1,1)}$, as shown in Figure 2. Then $G$ contains 5 induced cycles. But there are only 4 distinct compact induced subgraphs of $G$ which contains 4 induced cycles. This is because the 4 cycles around the triangle formed three big vertices $\{u,v,w\}$ does not form an induced subgraph of $G$. This shows that $G$ is a strict second-Betti graph.
$x_{1}$$x_{2}$$w_{1}$$w_{2}$$u_{2}$$u_{1}$$v_{1}$$v_{2}$$u$$v$$w$
Figure 2. The graph $\mathbf{C}_{(1):(1):(1,1)}$
Since type 3 compact graphs are strict second-Betti graphs, we need to compute their second top graded Betti numbers, leveraging Lemma 2.2. To apply Lemma 2.2, we make an observation related to the computation of simplicial complexes $\Gamma_{h}(I)$. Notably, we adhere to the notation established in Lemma 2.2 throughout our analysis. Let $\Omega$ be a set comprised of a number of monomials from $T$, and assume that the ideal $I$ is generated by all the monomials in $\Omega$. For each $\alpha\in\Omega$, we introduce a simplicial complex $\Gamma_{h}^{\alpha}(I)$ as follows. If $\alpha$ divides $Y^{h}$ and $Y^{h}/\alpha=y_{1}^{a_{1}}y_{2}^{a_{2}}\cdots y_{n}^{a_{n}}$ for some $a_{i}\geq 0$, then $\Gamma_{h}^{\alpha}(I)$ is specified as the complex: $F$ is a facet of $\Gamma_{h}^{\alpha}(I)$Longleftrightarrow there are $a_{i}\geq 0$ for $i\in[n]$ such that $Y^{h}=\alpha y_{1}^{a_{1}}y_{2}^{a_{2}}\cdots y_{n}^{a_{n}}$ and $F=\{i\in[n]\mid a_{i}\neq 0\}.$ Otherwise, $\Gamma_{h}^{\alpha}(I)$ is simply an empty set. We can easily check that
$\Gamma_{h}(I)=\bigcup_{\alpha\in\Omega}\Gamma_{h}^{\alpha}(I).$
Example 6.5.
Let $\mathbb{K}[G]$ be the edge ring of $G=\mathbf{C}_{(1):(1):(1,1)}$. Note that all the vertices of $G$ are labeled, as shown in Figure 2. Thus, $\Theta_{G}=x_{1}x_{2}w_{1}w_{2}u_{1}u_{2}v_{2}v_{2}$. In the following, we denote $\Theta_{G}$ by $\Theta$ for short. Moreover, in view of Lemma 5.4, the canonical module $\omega_{\mathbb{K}[G]}$, as an ideal of $\mathbb{K}[G]$, is generated by monomials
$\Theta uvw^{2},\Theta uvw^{4},\Theta u^{2}vw,\Theta uv^{2}w.$
The four monomials are denoted by $\beta_{i}$ for $i=1,\cdots,4$. Put $\alpha_{1}=\Theta u^{2}v^{2}w^{2}$ and $\alpha_{2}=\Theta u^{2}v^{2}w^{4}.$ Denote $uv,vw,wu$ by $y_{1},y_{2},y_{3}$ respectively.
Since $\Theta u^{2}v^{2}w^{2}=\beta_{1}y_{1}$, it follows that $\Gamma_{\alpha_{1}}^{\beta_{1}}(\omega_{\mathbb{K}[G]})=\langle\{1\}\rangle$;
Since $\beta_{2}$ does not divide $\Theta u^{2}v^{2}w^{2}$, it follows that $\Gamma_{\alpha_{1}}^{\beta_{2}}(\omega_{\mathbb{K}[G]})=\emptyset$;
Since $\Theta u^{2}v^{2}w^{2}=\beta_{3}y_{2}$, it follows that $\Gamma_{\alpha_{1}}^{\beta_{3}}(\omega_{\mathbb{K}[G]})=\langle\{2\}\rangle$;
Since $\Theta u^{2}v^{2}w^{2}=\beta_{4}y_{3}$, it follows that $\Gamma_{\alpha_{1}}^{\beta_{4}}(\omega_{\mathbb{K}[G]})=\langle\{3\}\rangle$.
Thus, $\Gamma_{\alpha_{1}}(\omega_{\mathbb{K}[G]})=\langle\{1\},\{2\},\{3\}\rangle$, and so $\dim_{\mathbb{K}}\widetilde{H}_{0}(\Gamma_{\alpha_{1}}(\omega_{\mathbb{K}[G]}),\mathbb{K})=2$.
Similarly, we have $\Gamma_{\alpha_{2}}^{\beta_{1}}(\omega_{\mathbb{K}[G]})=\langle\{2,3\}\rangle$, $\Gamma_{\alpha_{2}}^{\beta_{2}}(\omega_{\mathbb{K}[G]})=\langle\{1\}\rangle$ and $\Gamma_{\alpha_{2}}^{\beta_{3}}(\omega_{\mathbb{K}[G]})=\Gamma_{\alpha_{2}}^{\beta_{4}}=\emptyset$. Hence, $\Gamma_{\alpha_{2}}(\omega_{\mathbb{K}[G]})=\langle\{2,3\},\{1\}\rangle$ and it follows $\dim_{\mathbb{K}}\widetilde{H}_{0}(\Gamma_{\alpha_{2}}(\omega_{\mathbb{K}[G]}),\mathbb{K})=1$.
Note that $D_{G}=\Theta^{2}u^{4}v^{4}w^{6}$. Since $D_{G}/\alpha_{1}=\Theta u^{2}v^{2}w^{4}$ and $D_{G}/\alpha_{2}=\Theta u^{2}v^{2}w^{2}$, it follows from Lemma 2.2 as well as Lemma 2.7 that
$\beta_{3,\Theta u^{2}v^{2}w^{4}}(\mathbb{K}[G])=\beta_{1,\alpha_{1}}(\omega_{\mathbb{K}[G]})=2$
and
$\beta_{3,\Theta u^{2}v^{2}w^{2}}(\mathbb{K}[G])=\beta_{1,\alpha_{2}}(\omega_{\mathbb{K}[G]})=1.$
There are precisely four induced subgraphs of $G$, that consist of four minimal cycles of $G$. These four induced subgraphs are induced on $V(G)\setminus\{x_{1},x_{2}\}$ and $V(G)\setminus\{w_{1},w_{2}\}$, $V(G)\setminus\{u_{1},u_{2}\}$ and $V(G)\setminus\{v_{1},v_{2}\}$, denoted by $G_{1}$ and $G_{2}$, $G_{3}$ and $G_{4}$, respectively. According to Lemmas 5.4, we have
$\mathcal{N}_{G_{1}}=\Theta_{G_{1}}\{u^{2}v^{3}w^{3},u^{3}v^{2}w^{3},u^{3}v^{3}w^{2}\},\quad\mathcal{N}_{G_{2}}=\Theta_{G_{2}}\{u^{2}v^{3}w^{3},u^{3}v^{2}w^{3},u^{3}v^{3}w^{2}\}$
and
$\mathcal{N}_{G_{3}}=\Theta_{G_{3}}\{v^{3}w^{2},v^{3}w^{4},v^{2}w^{5}\},\quad\mathcal{N}_{G_{4}}=\Theta_{G_{4}}\{u^{3}w^{2},u^{3}w^{4},u^{2}w^{5}\}.$
Here, for a set $A$ of monomials and for a monomial $\alpha$, the notation $\alpha A$ denotes the set $\{\alpha a\mid a\in A\}$.
Combining the above facts, the sum of Betti numbers $\beta_{3,\alpha}(\mathbb{K}[G])$ with $\mbox{supp}(\alpha)=V(G)$ is greater than or equal to 3, and the sum of Betti numbers $\beta_{3,\alpha}(\mathbb{K}[G])$ with $\mbox{supp}(\alpha)=V(G_{i})$ is also greater than or equal to $\geq$ 3 for $i=1,2,3,4$. Therefore, it follows that
$1515\leq\beta_{3}(\mathbb{K}[G])\leq\beta_{3}(\mathbb{K}[E(G)]/\mathrm{in}_{<}(I_{G}))=(2+1)\binom{5}{4}=$ 15 .
Hence, $\beta_{3}(\mathbb{K}[G])=\beta_{3}(\mathbb{K}[E(G)]/\mathrm{in}_{<}(I_{G}))$, and $\beta_{3,\alpha}(\mathbb{K}[G])$ for $\alpha\in\mathbb{Z}^{11}$ could also be written down.
The approach outlined in Example 6.5 can be applied to the general case.
Definition 6.6.
Let $H$ be the graph $\mathbf{C}_{\underline{p}:\underline{q}:\underline{r}}$ with its vertices of $H$ labeled as in Subsection 3.3. We define $\mathcal{M}_{H}^{1}$, $\mathcal{M}_{H}^{2}$ and $\mathcal{M}_{H}$ to be the following sets:
$\mathcal{M}^{1}_{H}:=$$\{\Theta_{H}u^{2\ell}v^{2n}w^{2k}\mid\ell=1,\ldots,m-1\}\sqcup\{\Theta_{H}u^{2m}v^{2\ell}w^{2k}\mid\ell=1,\ldots,n-1\}$
$\sqcup\{\Theta_{H}u^{2m}v^{2n}w^{2\ell}\mid\ell=1,\ldots,k-1\},$
$\mathcal{M}_{H}^{2}:=$$\{\Theta_{H}u^{2m}v^{2n}w^{2k}\}\mbox{ and }\mathcal{M}_{H}:=\mathcal{M}_{H}^{1}\sqcup\mathcal{M}_{H}^{2}$
Lemma 6.7.
Let $H=\mathbf{C}_{\underline{p}:\underline{q}:\underline{r}}$ denote a compact graph of type 3 and let $p=m+n+k$ represent the projective dimension of $\mathbb{K}[H]$. Then,
$\beta_{p-1,\alpha}(\mathbb{K}[H])=\left\{\begin{array}[]{ll}1,&\hbox{$\alpha\in\mathcal{M}^{1}_{H}$;}\\ 2,&\hbox{$\alpha\in\mathcal{M}^{2}_{H}$.}\end{array}\right.$
In particular, we have $\sum\limits_{\alpha}\beta_{p-1,\alpha}(\mathbb{K}[H])\geq p-1$, where the summation is over all $\alpha\in\mathrm{Mon}(\mathbb{K}[H])$ with $\mathrm{supp}(\alpha)=V(H)$.
Proof.
Denote $uv,vw$ and $wu$ by $y_{1},y_{2}$ and $y_{3}$ respectively. Let $\omega$ denote the canonical module $\omega_{\mathbb{K}[H]}$. Note that $\omega$ is a monomial ideal of $\mathbb{K}[H]$, whose generators are given in Lemma 5.4. Put $\alpha=\Theta u^{2}v^{2}w^{2}$, where $\Theta$ is used as a shorthand for $\Theta_{H}$. Then
$\Gamma_{\alpha}^{\Theta u^{2}vw}(\omega)=\langle\{2\}\rangle,\Gamma_{\alpha}^{\Theta uv^{2}w}(\omega)=\langle\{3\}\rangle,\Gamma_{\alpha}^{\Theta uvw^{2}}(\omega)=\langle\{1\}\rangle.$
For other generators $\beta$ of $\omega$, one easily sees that $\Gamma_{\alpha}^{\beta}(\omega)=\emptyset$. This implies $\Gamma_{\alpha}(\omega)=\langle\{2\},\{3\},\{1\}\rangle$ and so $\beta_{1,\alpha}(\omega)=2$ by Lemma 2.2. In the same vein, we can prove
$\beta_{1,\Theta u^{2\ell}v^{2}w^{2}}(\omega)=1\mbox{ for }\ell=2,3,\ldots,m,$
$\beta_{1,\Theta u^{2}v^{2\ell}w^{2}}(\omega)=1\mbox{ for }\ell=2,3,\ldots,n,$
and
$\beta_{1,\Theta u^{2}v^{2}w^{2\ell}}(\omega)=1\mbox{ for }\ell=2,3,\ldots,k.$
Note that $D_{H}=\Theta^{2}u^{2m+2}v^{2n+2}w^{2k+2},$ the first assertion follows from Lemma 2.7. For the second assertion, it follows from the facts that $|\mathcal{M}^{1}_{H}|=m+n+k-3$, $|\mathcal{M}^{2}_{H}|=1$, and $\mathrm{supp}(\alpha)=V(H)$ for each $\alpha\in\mathcal{M}^{1}_{H}\sqcup\mathcal{M}^{2}_{H}$.
Proposition 6.8.
Let $G=\mathbf{C}_{\underline{p}:\underline{q}:\underline{r}}$ denote a compact graph of type 3. Then
$\beta_{i}(\mathbb{K}[G])=\beta_{i}(\mathbb{K}[E(G)]/\mathrm{in}_{<}(I_{G}))$
for all $i\geq 0$.
Proof.
We categorize the minimal cycles of $G$ into four distinct classes. The cycle encompassing the three big vertices is designated as the big cycle. A minimal cycle that is distinct from the big cycle but includes the vertex $u$ (or $v$, or $w$ again) is referred to as a cycle belonging to $u$ (or $v$, or $w$ respectively).
Fix $i\geq 1$. Then the number of subgraphs of $G$ that contain exactly $i+1$ minimal cycles is given by $\binom{m+n+k+1}{i+1}$. But not all of these subgraphs are induced subgraphs. Consider a subgraph $H$ of $G$ and let $C_{1},\ldots,C_{i+1}$ represent all its minimal cycles. Then one easily sees that $H$ is not an induced subgraph if and only if
  • None of $C_{j}$ for $j=1,\ldots,i+1$ is the big cycle;
  • For each $x\in\{u,v,w\}$, there is at least one index $j\in[i+1]$ such that $C_{j}$ is a minimal cycle belonging to $x$.
Hence, there are $\sum\limits_{a+b+c=i+1}^{a,b,c\geq 1}\binom{m}{a}\binom{n}{b}\binom{k}{c}$ such subgraphs and it follows that
$\sum_{\alpha}\beta_{i,\alpha}(\mathbb{K}[G])\geq\left(\binom{m+n+k+1}{i+1}-\sum\limits_{a+b+c=i+1}^{a,b,c\geq 1}\binom{m}{a}\binom{n}{b}\binom{k}{c}\right)i,$
where the summation is over all $\alpha\in\mathrm{Mon}(\mathbb{K}[G])$ such that the induced subgraph on $\mathrm{supp}(\alpha)$ consisting of $i+1$ minimal cycles. On the other hand, we notice the following fact: if $H$ is an induced subgraph of $G$ containing exactly $i+2$ minimal cycles, say $C_{1},\ldots,C_{i+2}$, then $H$ belongs to type 3 if and only if
  • One of $C_{j}$ for $j=1,\ldots,i+2$ is the big cycle. For simplicity, let us assume that $C_{i+2}$ is the big cycle;
  • For each $x\in\{u,v,w\}$, there is at least one index $j\in[i+1]$ such that $C_{j}$ is a minimal cycle belonging to $x$.
Thus, there are $\sum\limits_{a+b+c=i+1}^{a,b,c\geq 1}\binom{m}{a}\binom{n}{b}\binom{k}{c}$ compact induced subgraphs of $G$ that belong to type 3 and contain exactly $i+2$ minimal cycles. It follows from Lemma 6.7 that
$\sum_{\alpha}\beta_{i,\alpha}(\mathbb{K}[G])\geq\sum\limits_{a+b+c=i+1}^{a,b,c\geq 1}\binom{m}{a}\binom{n}{b}\binom{k}{c}i,$
where the summation is over all $\alpha\in\mathrm{Mon}(\mathbb{K}[G])$ such that the induced subgraph on $\mathrm{supp}(\alpha)$ is a compact graph of type 3 consisting of $i+2$ minimal cycles. Hence, we have
$\beta_{i}(\mathbb{K}[G])\geq\binom{m+n+k+1}{i+1}i.$
Now, the result follows from Lemma 6.1.
From the proof of Proposition 6.8, we see that the inequality in Lemma 6.7 is actually an equality. By combining all the findings we have made so far, we can immediately obtain the main result of this section.
Theorem 6.9.
Let $G$ be a compact graph, and denote by $p=\mathfrak{t}(G)-1$ the projective dimension of $\mathbb{K}[G]$. Consider an integer $i$ satisfying $1\leq i\leq p$. Let $H_{j}$ for $j=1,\ldots,s$ represent all the compact induced subgraphs of $G$ having exactly $i+1$ minimal cycles, and let $L_{j}$ for $j=1,\ldots,t$ be all the compact induced subgraphs of $G$ belonging to type 3 with $i+2$ minimal cycles. Define the following sets:
$\mathcal{N}_{i}(G)=\bigcup_{j=1}^{s}\mathcal{N}_{H_{j}},\quad\mathcal{M}_{i}^{1}(G)=\bigcup_{j=1}^{t}\mathcal{M}_{L_{j}}^{1},\quad\text{and}\quad\mathcal{M}_{i}^{2}(G)=\bigcup_{j=1}^{t}M_{L_{j}}^{2}.$
Then, the $i$-th multi-graded Betti numbers of $\mathbb{K}[G]$ can be expressed as follows:
$\beta_{i,\alpha}(\mathbb{K}[G])=\begin{cases}2,&\text{if }\alpha\in\mathcal{M}_{i}^{2}(G),\\ 1,&\text{if }\alpha\in\mathcal{M}_{i}^{1}(G)\cup\mathcal{N}_{i}(G),\\ 0,&\text{otherwise}.\end{cases}$
Note that $t$ is always equal to 0 if $G$ is a compact graph of type 1 or 2.
For all possible integers $i,j$, we set $O_{i,j}=\{\alpha\in\mathcal{M}_{i+1}^{1}(G)\cup\mathcal{N}_{i+1}(G)\mid|\mathrm{deg}(\alpha)|=2j\}$ and $T_{i,j}=\{\alpha\in\mathcal{M}_{i+1}^{2}(G)\mid|\mathrm{deg}(\alpha)|=2j\}$.
Corollary 6.10.
Let $G$ be a compact graph, and let $I_{G}$ denote its toric ideal. There exists a square-free initial ideal $\mathrm{in}_{<}(I_{G})$ of $I_{G}$ such that $\beta_{i,j}(\mathrm{in}_{<}(I_{G}))=\beta_{i,j}(I_{G})=|O_{i.j}|+2|T_{i,j}|$.
Acknowledgment: This project is supported by NSFC (No. 11971338). The authors are grateful to the software systems [7] for providing us with a large number of examples that help us develop ideas and test the results.