### v3

# Betti numbers of normal edge rings (I)

Abstract.
We introduce a novel approach named the induced-subgraph approach for investigating the Betti numbers of normal edge rings. Utilizing this approach, we compute all the multi-graded Betti numbers of the edge rings associated with two-ear graphs (Definition 4.1) and compact graphs (Definition 5.1). In particular, we show that for two-ear graphs and compact graphs of type 1 or 2, their multi-graded Betti numbers are always equal to the top multi-graded Betti numbers of some of their induced subgraphs. In contrast, some of the multi-graded Betti numbers of compact graphs of type 3 are not the top multi-graded Betti numbers of any of their induced subgraph. We speculate that our approach can be applicable to many other normal edge rings.
Key words and phrases:
Induced-subgraph approach, Normal edge ring, Multi-graded Betti number, Top-Betti graph, Compact graph, Canonical module
2010 Mathematics Subject Classification:
Primary 13A02 Secondary 05E40.
1. Introduction
[…]
According to [16, Chapter IV], both $I_{G}$ and $\mathbb{K}[G]$ admit both minimal $\mathbb{Z}_{\geq 0}^{n}$-graded and minimal $\mathbb{Z}_{\geq 0}$-graded free resolutions over the polynomial ring $\mathbb{K}[E(G)]$. This fact allows us to establish the following relationships between their Betti numbers. It is clear that $\beta_{i,h}(I_{G})=\beta_{i+1,h}(\mathbb{K}[G])$ for all $i\geq 0$ and $h\in\mathbb{Z}_{\geq 0}^{n}$, and $\beta_{i,j}(I_{G})=\beta_{i+1,j}(\mathbb{K}[G])$ for all $i\geq 0$ and $j\in\mathbb{Z}_{\geq 0}$. Define $|h|=\sum_{k=1}^{n}h_{k}$ for $h=(h_{1},\ldots,h_{n})\in\mathbb{Z}_{\geq 0}^{n}$. Then, we have for $i\geq 0$ and $j\geq 0$, the following equality holds:
$\beta_{i,j}(I_{G})=\sum_{\begin{subarray}{c}h\in\mathbb{Z}_{\geq 0}^{n}\\ |h|=2j\end{subarray}}\beta_{i,h}(I_{G}).$
Due to the difficulty of calculating multi-graded Betti numbers or even graded Betti numbers of $I_{G}$ for arbitrary graphs, researchers are inclined to investigate restricted families. In [3, Theorem 5.1], the graded Betti numbers of $I_{\mathbf{K}_{2,d}}$ are obtained, where $\mathbf{K}_{2,d}$ is a complete bipartite graph with a bipartition $\{x_{1},x_{2}\}$ and $\{y_{1},\ldots,y_{d}\}$. In [6], Galetto et al. computed the graded Betti numbers of the toric ideal of $G_{r,d}$, where $G_{r,d}$ is the graph formed by adding a path of length $22r-$ 2 between the two vertices $x_{1},x_{2}$ of degree $d$ in $\mathbf{K}_{2,d}$. This was primarily achieved by observing that the toric ideal of $G_{r,d}$ possesses an initial ideal which has linear quotients. Lately, Nandi and Nanduri [13] added a path of length $22s-$ 2 between the two vertices of degree $d+1$ in $G_{r,d}$ to obtain the graph $G_{r,s,d}$. They proved that there exists an initial ideal $\mathrm{in}_{<}(I_{G_{r,s,d}})$ of $I_{G_{r,s,d}}$ which has regular quotients. Here we say a monomial ideal $I$ has regular quotients if there exists an order $u_{1},\ldots,u_{t}$ of minimal monomial generators of $I$ such that the colon ideal $(u_{1},\ldots,u_{k}):u_{k+1}$ is a complete intersection, i.e., it is generated by a regular sequence for $k=1,\ldots,t-1$. By comparing the graded Betti numbers of $I_{G_{r,s,d}}$ with the graded Betti numbers of $I_{G_{r,d}}$ and of $\mathrm{in}_{<}(I_{G_{r,s,d}})$, they computed all the graded Betti numbers of $I_{G_{r,s,d}}$. In those computations, the Hilbert series also play an important role, as a bridge connecting the graded Betti numbers of $I_{G}$ and those of an initial ideal of $I_{G}$. One may see that computing the Betti numbers of $I_{G_{r,s,d}}$ is much more complicated than computing the Betti numbers of $I_{G_{r,d}}$. So far, there are no other examples $G$ where all graded Betti numbers $\beta_{i,j}(I_{G})$ have been computed in the literatures.
In this paper, we will introduce a unified approach to evaluate the graded Betti numbers of the toric ideal of a simple graph satisfying the odd-cycle condition. By definition, a simple graph satisfies the odd-cycle condition whenever any two minimal cycles $C_{1}$ and $C_{2}$ either share at least one vertex or are connected by an edge (which is) bridging a vertex from $C_{1}$ to a vertex from $C_{2}$. Recall that Simis et al. in [18] and Ohsugi-Hibi in [14] independently established that the edge ring $\mathbb{K}[G]$ is a normal algebra precisely when the graph $G$ fulfills the odd-cycle condition. It is worth noting that $\mathbf{K}_{2,d}$, $G_{r,d}$ and $G_{r,s,d}$ all fall into this class since they are all bipartite graphs.
As we will see, compared with previous methods, the new approach offers several advantages. First, it can be applied to a wide range of classes of graphs that satisfy the odd-cycle condition. For instance, in this paper we compute the multi-graded Betti numbers for two-ear graphs and compact graphs, while the multi-path graphs are examined in a subsequent paper [22]. By definition, a multi-path graph is a simple graph composed of two distinct vertices connected by multiple paths, including families such as $\mathbf{K}_{2,d}$, $G_{r,d}$, and $G_{r,s,d}$. It is worth noting that the previous computations of Betti numbers for $G_{r,d}$ and $G_{r,s,d}$ differed in several respects. Second, while we still need to identify an initial ideal for $I_{G}$ that has regular quotients, we only require approciate upper bounds for the total Betti numbers of this initial ideal rather than precise graded Betti numbers. Third, our approach enables us to obtain multi-graded Betti numbers instead of merely graded Betti numbers.
We first provide some definitions and notions to facilitate the introduction of our approach and the main results of this paper. Assuming the normality of $\mathbb{K}[G]$, [1, Theorem 6.31] reveals that the multi-graded canonical module of the multi-graded $\mathbb{K}[E(G)]$-algebra $\mathbb{K}[G]$, denoted as $\omega_{\mathbb{K}[G]}$, is isomorphic to an ideal of $\mathbb{K}[G]$ with a generating set that possesses a geometric interpretation, thereby simplifying its computation. Leveraging this understanding, we can determine the top multi-graded Betti numbers of $\mathbb{K}[G]$ by invoking Formula 6.6 from [1]. Note that the top multi-graded Betti numbers of a finitely generated $\mathbb{Z}^{n}$-graded module $M$ refer to the numbers $\beta_{p,h}(M)$ for $h\in\mathbb{Z}^{n}$ if $M$ has a projective dimension $p$. We will also utilize the terminology the second top multi-graded Betti numbers of $M$ to refer to the numbers $\beta_{p-1,h}(M)$ for $h\in\mathbb{Z}^{n}$.
Besides [1, Formula 6.6], our approach for computing the Betti numbers of normal edge rings also relies on the following results. The first result is as follows.
Lemma 1.1.
(Lemma 2.8) Let $H$ be an induced subgraph of a simple graph $G$. Then for $h\in\mathbb{Z}_{\geq 0}^{V(G)}$ with $\mathrm{supp}(h)\subseteq V(H)$, we have
$\beta_{i,h}(\mathbb{K}[H])=\beta_{i,h}(\mathbb{K}[G])\qquad\forall i\geq 0.$
The second result we need is a generalization of Hochster’s formula. Let $T$ represent the subalgebra of the polynomial ring $\mathbb{K}[Y_{1},\ldots,Y_{m}]$ generated by monomials $y_{1},\ldots,y_{n}$. Within $T$, all the monomials form a monoid under multiplication, denoted by $\mathcal{U}$. Let $R$ be the polynomial ring $\mathbb{K}[X_{1},\ldots,X_{n}]$. We endow it with a multi-grading in $\mathbb{Z}^{m}$ by assigning the multi-degree of $y_{i}$ to $X_{i}$, where $i=1,\ldots,n$. There is a multi-graded homomorphism $\phi$ from $R$ to $T$ sending $X_{i}$ to $y_{i}$ for $i=1,\ldots,n$. Consider an ideal $I$ of $T$ generated by certain monomials from $\mathcal{U}$. Note that $I$ is a finitely generated multi-graded $T$-module, so it is a a finitely generated multi-graded $R$-module as well via the map $\phi$. The formula for expressing $\beta_{i,h}^{R}(T/I)$ has been provided in [2]. However, to compute the Betti numbers of the canonical module of a normal edge ring, we require the following lemma.
Lemma 1.2.
(Lemma 2.2) Let $I$ be a monomial ideal of $T$. To each multi-degree $h\in\mathbb{Z}_{\geq 0}^{m}$, we associate a simplicial complex
$\Gamma_{h}(I):=\{F\subseteq[n]\mid Y^{h}=y^{F}\alpha\mbox{ for some monomial }\alpha\in I\}.$
Then, for all $i\geq 0$, the multi-graded Betti numbers of $I$ are expressed as
$\beta_{i,h}^{R}(I)=\dim_{\mathbb{K}}\widetilde{H}_{i-1}(\Gamma_{h}(I),\mathbb{K}).$
We are ready to introduce our approach. Briefly speaking, our approach expresses every multi-graded Betti number of a normal edge ring $\mathbb{K}[G]$ as the top or second-top multi-graded Betti number of the edge ring associated with certain induced subgraph of $G$, and the latter can be derived from the data of its canonical module.
The detailed steps of our approach is as follows. Let $G$ be a connected graph that satisfies the odd-cycle condition.
  1. (1)
    Initially, we compute the projective dimension $p$ of $\mathbb{K}[G]$, and then identify a term order $<$ such that, for each $i$ with $1\leq i\leq p$, we can derive a suitable upper bound for the total Betti number $\beta_{i}(\mathbb{K}[E(G)]/\mathrm{in}_{<}(I_{G}))$. In practice, the upper bounds are achieved by showing that $\mathrm{in}_{<}(I_{G})$ has regular quotients.
  2. (2)
    Secondly, we compute the top multi-graded Betti numbers $\beta_{p,h}(\mathbb{K}[G])$ of $\mathbb{K}[G]$ by determining the minimal generators of $\omega_{\mathbb{K}[G]}$.
  3. (3)
    Next, for indices $i$ with $0<i<p$, we identify all of the induced subgraphs $H$ of the graph $G$ such that (i) the projective dimension of $\mathbb{K}[H]$ is equal to $i$; (ii) if $\beta_{i,h}(\mathbb{K}[H])\neq 0$ for some $h\in\mathbb{Z}_{\geq 0}^{V(H)}$, then $\mbox{supp}(h)=V(H)$. Subsequently, by leveraging [1, Formula 6.6], we compute the top multi-graded Betti numbers of the edge rings associated with these induced subgraphs, and then sum them up.
  4. (4)
    If the sum obtained in the third step matches the upper bound for the total Betti number of $\mathbb{K}[E(G)]/\mathrm{in}_{<}(I_{G})$ obtained the first step, then we have achieved our goal of computing the multi-graded Betti numbers $\beta_{i,h}(\mathbb{K}[G])$ for $h\in\mathbb{Z}_{\geq 0}^{V(G)}$. In this case, $\beta_{i,h}(\mathbb{K}[G])$ is equal to the top Betti number $\beta_{i,h}(\mathbb{K}[H])$, where $H$ is the induced graph of $G$ on $\mathrm{supp}(h)$. Moreover, if $\beta_{i,h}(\mathbb{K}[G])\neq 0$, then the projective dimension of $\mathbb{K}[H]$ is equal to $i$. In addition, the given upper bound for $\beta_{i}(\mathbb{K}[E(G)]/\mathrm{in}_{<}(I_{G}))$ is actually the exact value of $\beta_{i}(\mathbb{K}[E(G)]/\mathrm{in}_{<}(I_{G}))$.
  5. (5)
    Otherwise, we have to consider induced subgraphs with edge rings of projective dimension $i+1$ and to compute the second top Betti numbers of the edge rings of these graphs. This is conducted by [1, Formula 6.6] together with Lemma 1.2.
We refer to this approach as the induced-subgraph approach. In view of the second step of this approach, we introduce the following definition.
Definition 1.3.
Let $G$ be a connected graph that fulfills the odd-cycle condition. Denote by $p$ the projective dimension of $\mathbb{K}[G]$. We call $G$ to be a top-Betti graph provided that if $\beta_{p,h}(\mathbb{K}[G])\neq 0$ for some $h\in\mathbb{Z}_{\geq 0}^{V(G)}$, then $\mathrm{supp}(h)=V(G)$. Here and after, for $h\in\mathbb{Z}_{\geq 0}^{V(G)}$, $\mathrm{supp}(h)$ represents the set of $x_{i}\in V(G)$ with $h(x_{i})\neq 0$.
A combinatorial characterization for top-Betti graphs is presented in Proposition 3.1. Based on this result, it is not difficult to determine all the top-Betti induced subgraphs of any given graph that meets the odd-cycle condition. This fact eases our study.
In the fifth step, we need to discern strict second-Betti induced subgraphs of a given graph, which is defined below.
Definition 1.4.
Let $G$ be a connected graph that fulfills the odd-cycle condition. Denote by $p$ the projective dimension of $\mathbb{K}[G]$. For a subset $S\subseteq V(G)$, let $G[S]$ stand for the induced graph of $G$ on $S$. We call $G$ to be a second-Betti graph provided that if $\beta_{p-1,h}(\mathbb{K}[G])\neq 0$ for some multi-degree $h$, then either $\mathrm{supp}(h)=V(G)$ or the projective dimension of $\mathbb{K}[G[\mathrm{supp}(h)]]$ is equal to $p-1.$ A second-Betti graph is termed strict if there is a multi-degree $h$ such that $\mathrm{supp}(h)=V(G)$ and $\beta_{p-1,h}(\mathbb{K}[G])\neq 0$.
A sufficient condition for $G$ to be a second-Betti subgraph are given in Proposition 3.5, which enables us to discern all strict second-Betti induced subgraphs of the graphs that we are studying.
In this paper, we successfully utilize this approach to compute the multi-graded Betti numbers of edge rings associated with two distinct classes of simple graphs that adhere to the odd-cycle condition. To introduce our main results, we need an additional notation.
Definition 1.5.
Let $H$ be a simple graph satisfying the odd-cycle condition. We define $\mathcal{N}_{H}$, which we call the top support of $H$, as the set of monomials $X^{h}$ for which $\beta_{p,h}(\mathbb{K}[H])$ is non-zero, where $p$ is the projective dimension of $\mathbb{K}[H]$. It is evident that the top multi-graded Betti numbers of $\mathbb{K}[H]$ are determined by the set $\mathcal{N}_{H}$.
The first class of graphs under consideration is $\mathbf{G}_{m}$, which was introduced in [9] for all $m\geq 1$. Its precise definition could be found in Definition 4.1. Owing to its shape, as depicted in Figure 1, we refer to the graph $\mathbf{G}_{m}$ as a two-ear graph. In this paper, we show that the multi-graded Betti numbers of the edge ring of $\mathbf{G}_{m}$ could be computed using the initial four steps of the induced-subgraph approach.
For convenience, when describing the multi-graded Betti numbers of an edge ring, we identify a monomial with its multi-degree.
Theorem 1.6.
(Theorems 4.12) Let $G$ denote a two-ear graph $\mathbf{G}_{m}$ for some $m\geq 1$. For $i\geq 1$, let $\mathcal{N}_{i}(G)$ denote the disjoint union of all $\mathcal{N}_{H}$ such that $H$ is a top-Betti induced subgraph of $G$ whose edge ring has a projective dimension of $i$. Then for any multi-degree $h$ with $X^{h}$ belonging to $\mathbb{K}[G]$ and for any integer $i\geq 1$, we have
$\beta_{i,h}(\mathbb{K}[G])=\left\{\begin{array}[]{ll}1,&\hbox{$X^{h}\in\mathcal{N}_{i}(G)$;}\\ 0,&\hbox{otherwise.}\end{array}\right.$
In particular, there exists a square-free initial ideal $\mathrm{in}_{<}(I_{G})$ of $I_{G}$ such that
$\beta_{i,j}(\mathrm{in}_{<}(I_{G}))=\beta_{i,j}(I_{G})$ for all $i,j$.
We note that there are exactly three types of top-Betti induced subgraphs in a two-ear graph, namely, two-ear graphs, one-ear graphs and complete bipartite graphs of the form $\mathbf{K}_{2,d}$. The top supports for these subgraphs are provided in Lemma 4.5, and Corollaries 4.10 and 4.11, respectively. As immediate applications, we can derive the regularity of $\mathbb{K}[G]$ when $G$ is a two-ear graph or an one-ear graph.
The second class of graphs under consideration is compact graphs. Let $G$ be a connected simple graph where every vertex has a degree at least two. Following [21], we call $G$ compact if it lacks even cycles and satisfies the odd-cycle condition. A vertex of a compact graph is said to be a big vertex if it has a degree greater than 2. To a large extent, the structure of a compact graph is determined by the quantity of its big vertices. In view of [21, Proposition 3.3], there are at most three big vertices in a compact graph. In the scenario where $G$ possesses three big vertices, i.e., $G$ is a compact graph of type 3, the fifth step of the induced-subgraph approach becomes necessary. The following theorem presents the multi-graded Betti numbers of the edge rings of compact graphs. Henceforth, $\sqcup$” denotes the disjoint union throughout this paper.
Theorem 1.7.
(Theorem 6.9 and Corollary 6.10) Let $G$ be a compact graph, and denote by $p$ the projective dimension of $\mathbb{K}[G]$. Consider an integer $i$ satisfying $1\leq i\leq p$. Let $H_{j}$ for $j=1,\ldots,s$ represent all the compact induced subgraphs of $G$ having exactly $i+1$ minimal cycles, and let $L_{j}$ for $j=1,\ldots,t$ be all the compact induced subgraphs of $G$ belonging to type 3 with $i+2$ minimal cycles. Define the following sets:
$\mathcal{N}_{i}(G)=\bigsqcup_{j=1}^{s}\mathcal{N}_{H_{j}},\quad\mathcal{M}_{i}^{1}(G)=\bigsqcup_{j=1}^{t}\mathcal{M}_{L_{j}}^{1},\quad\text{and}\quad\mathcal{M}_{i}^{2}(G)=\bigsqcup_{j=1}^{t}M_{L_{j}}^{2}.$
Then, the $i$-th multi-graded Betti numbers of $\mathbb{K}[G]$ can be expressed as follows:
$\beta_{i,h}(\mathbb{K}[G])=\begin{cases}2,&\text{if }X^{h}\in\mathcal{M}_{i}^{2}(G),\\ 1,&\text{if }X^{h}\in\mathcal{M}_{i}^{1}(G)\sqcup\mathcal{N}_{i}(G),\\ 0,&\text{otherwise}.\end{cases}$
Here, if $H$ is a labelled compact graph, then the monomial sets $\mathcal{N}_{H}$, $\mathcal{M}_{H}^{1}$ and $\mathcal{M}_{H}^{2}$ can be derived from $H$ through Lemma 5.2, Lemma 5.3, Lemma 5.5, and Definition 6.6. Note that if $G$ is of type one or two, then $t$ is always zero.
Furthermore, there exists a square-free initial ideal $\mathrm{in}_{<}(I_{G})$ of $I_{G}$ such that
$\beta_{i,j}(\mathrm{in}_{<}(I_{G}))=\beta_{i,j}(I_{G})$ for all $i,j$.
From the perspective of homological algebra, combining the above theorems reveals that compact graphs of type 1 or 2 are similar to two-ear graphs, yet they are significantly different from compact graphs of type 3. The reason is that the multi-graded Betti numbers of the former are always the top multi-graded Betti numbers of some of their induced subgraphs. In contrast, and this is a key differentiating point, some of the multi-graded Betti numbers of the latter are not the top multi-graded Betti numbers of any of their induced subgraph; instead, they are the second multi-graded Betti numbers of certain induced subgraphs. Specifically, at least one of the multi-Betti numbers of the latter is equal to 2, while those of the former are always either 0 or 1.
In light of the above theorems, when $G$ is either a two-ear graph or a compact graph, all the graded Betti numbers of $I_{G}$ are equal to those of a square-free initial ideal of $I_{G}$. Thus, this work aligns with the spirit of Conca and Varbaro’s paper [5], which established that the extremal Betti numbers of an ideal are identical to the extremal Betti numbers of its initial ideal if the initial ideal is square-free.
The paper is structured as follows. In Section 2, we lay the groundwork by introducing fundamental concepts and preliminary results. In Section 3, we present basic properties of top-Betti graphs and second-Betti graphs. Subsequently, in section 4 we compute the multi-graded Betti numbers of edge rings associated with two-ear graphs. As induced subgraphs of two-ear graphs, the multi-graded Betti numbers of edge rings associated with one-ear graphs and complete bipartite graphs with one part consisting of two elements are also given. In Sections 5 and 6, we explore the Betti-numbers of the edge rings of compact graphs, culminating in the presentation of Theorem 6.9.
2. Preliminaries
In this section, we provide a brief review of the notation and fundamental facts that will be utilized later on.
2.1. Betti numbers and Monomial ideals with regular quotients
Let $\mathbb{K}$ be a field, and let $R:=\mathbb{K}[X_{1},\ldots,X_{n}]$ be the polynomial ring in variables $X_{1},\ldots,X_{n}$ that are $\mathbb{Z}^{m}$-graded. Here, we do not require that $m=n$. For a finitely generated $\mathbb{Z}^{m}$-graded $R-$module $M$, there exists the minimal multi-graded free resolution of $M$ that has the form:
$0\rightarrow\underset{h\in\mathbb{Z}^{m}}{\bigoplus}R[-h]^{\beta_{p,h}(M)}\rightarrow\cdots\rightarrow\underset{h\in\mathbb{Z}^{m}}{\bigoplus}R[-h]^{\beta_{1,h}(M)}\rightarrow\underset{h\in\mathbb{Z}^{m}}{\bigoplus}R[-h]^{\beta_{0,h}(M)}\rightarrow M\rightarrow 0$
Note that $R[-h]$ is the cyclic free $R$-module generated in multi-degree $h$. The projective dimension of $M$ are defined to be $\mathrm{pdim}\,(M):={\mbox{max}}\,\{i\mid\beta_{i,\,h}(M)\neq 0\mbox{ for some}h\}.$ The number $\beta_{i,h}(M):={\rm{dim}}_{\mathbb{K}}\mathrm{Tor}_{i}^{R}(M,\mathbb{K})_{h}$ is called the $(i,h)$-th multi-graded Betti number of $M$ and $\beta_{i}(M):=\sum_{h\in\mathbb{Z}^{m}}\beta_{i,h}(M)$ is called the $i$-th total Betti number of $M$.
Denote by $p$ the projective dimension of the module $M$. Subsequently, $\beta_{p}(M)$ is referred to as the top total Betti number of $M$, while $\beta_{p,h}(M)$ for $h\in\mathbb{Z}^{m}$ are designated as the top multi-graded Betti numbers of $M$. On the other hand, $\beta_{p-1}(M)$ is labeled as the second top total Betti number of $M$, and $\beta_{p-1,h}(M)$ for $h\in\mathbb{Z}^{m}$ are designated as the second top multi-graded Betti numbers of $M$.
To derive upper bounds for Betti numbers of toric ideals under our investigation, we need a notion that was originally introduced in [11] and subsequently explored in [10]. Recall that a monomial ideal $I$ in $R$ is said to have regular quotients if there exists a linear ordering $\alpha_{1}\prec\alpha_{2}\prec\ldots\prec\alpha_{r}$ among the minimal generators $\{\alpha_{1},\alpha_{2},\ldots,\alpha_{r}\}$ of $I$ such that, for each $s=2,\ldots,r$, the quotient ideal $(\alpha_{1},\ldots,\alpha_{s-1}):\alpha_{s}$ is generated by a regular sequence. Notably, a sequence of monomials $\beta_{1},\ldots,\beta_{r}$ is regular precisely when the supports of distinct monomials are disjoint, i.e., $\mathrm{supp}(\beta_{i})\cap\mathrm{supp}(\beta_{j})=\emptyset$ for all $1\leq i<j\leq r$. Here, the support of a monomial $\alpha$, denoted by $\mathrm{supp}(\alpha)$, refers to the set of variables that divide $\alpha$.
Assuming that the colon ideal $(\alpha_{1},\ldots,\alpha_{s-1}):\alpha_{s}$, denoted as $L_{s}$, is generated by a regular sequence of length $k_{s}$ for each $s$ ranging from $2$ to $r$, we can apply Koszul theory to deduce that the total Betti numbers $\beta_{i}(R/L_{s})$ are given by $\beta_{i}(R/L_{s})=\binom{k_{s}}{i}$ for all non-negative integers $i$. Let $I_{s}$ represent the monomial ideal generated by $\alpha_{1},\ldots,\alpha_{s}$, and let $d_{s}$ represent the degree of $\alpha_{s}$ for each $s$ from $1$ to $r$.
Now, consider the following short exact sequences:
$0\rightarrow R/L_{s}[-d_{s}]\rightarrow R/I_{s-1}\rightarrow R/I_{s}\rightarrow 0,\quad\text{for all}\ s=2,\ldots,r.$
Utilizing the mapping cone construction (see e.g. [16, section 27] for the details), we can deduce that
$\beta_{i}(R/I_{s})\leq\beta_{i-1}(R/L_{s})+\beta_{i}(R/I_{s-1}).$
Repeatedly applying the above inequality, we arrive at
$\beta_{i}(R/I)\leq\beta_{i-1}(R/L_{r})+\cdots+\beta_{i-1}(R/L_{2})+\beta_{i}(R/(\alpha_{1}))$
and
$\beta_{i}(I)\leq\beta_{i}(R/L_{r})+\cdots+\beta_{i}(R/L_{2})+\beta_{i}(R).$
This implies that, for all non-negative integers $i$, we have the following inequality, which will be crucial in our subsequent discussions:
(1) $\beta_{i}(I)\leq\sum_{j=2}^{r}\binom{k_{j}}{i}+\binom{0}{i}.$
Apart from this formula, the following result, as mentioned, for instance, in [8, Corollary 3.3.3], is necessary for us to derive the upper bound for Betti numbers of edge rings.
Lemma 2.1.
Let $I$ be a graded ideal of the polynomial ring $R=\mathbb{K}[X_{1},\ldots,X_{n}]$, and let $<$ be any monomial order on $R$. Then for all $i,j\geq 0$, we have
$\beta_{i,j}(R/I)\leq\beta_{i,j}(R/\mathrm{in}_{<}(I)).$
In particular,
$\beta_{i}(R/I)\leq\beta_{i}(R/\mathrm{in}_{<}(I)).$
2.2. Betti numbers of monomial ideals in a monomial algebras
Let $T$ represent the subalgebra of the polynomial ring $\mathbb{K}[Y_{1},\ldots,Y_{m}]$ generated by monomials $y_{1},\ldots,y_{n}$. In [19], $T$ is called a monomial algebra. Within $T$, all the monomials form a monoid under multiplication, denoted by $\mathcal{U}$, and $T$ is isomorphic to the positive affine monoid ring $\mathbb{K}[\mathcal{U}]$. We define $R$ as the polynomial ring $\mathbb{K}[X_{1},\ldots,X_{n}]$ and endow it with a multi-grading in $\mathbb{Z}^{m}$ by assigning the multi-degree of $y_{i}$ to $X_{i}$. Consider an ideal $I$ of $T$ generated by certain monomials from $\mathcal{U}$.
By the existence of a surjective homomorphism from $R$ to $T$ mapping $X_{i}$ to $y_{i}$, it follows that $T$, $I$ and $T/I$ are all finitely generated $\mathbb{Z}^{m}$-graded $R$-modules.
A formula for computing of the multi-graded Betti numbers of $T/I$ was previously established in [2] using relative simplicial homology. However, in this paper, we require to calculate the multi-graded Betti numbers of $I$. For this, we associate every multi-degree $h\in\mathbb{Z}_{\geq 0}^{m}$ with the following simplicial complex:
$\Gamma_{h}(I):=\{F\subseteq[n]\mid Y^{h}=y^{F}\alpha\mbox{ for some monomial }\alpha\in I\}.$
Here, $Y^{h}$ is defined as $Y_{1}^{h_{1}}\cdots Y_{m}^{h_{m}}$ for $h=(h_{1},\ldots,h_{m})\in\mathbb{Z}_{\geq 0}^{m}$, and $y^{F}$ represents the product of all monomials $y_{i}$ with $i\in F$.
Lemma 2.2.
Let $I$ be a monomial ideal of $T$, and let $h\in\mathbb{Z}_{\geq 0}^{m}$ be a multi-degree. Then, for all $i\geq 0$, one has
$\beta_{i,h}^{R}(I)=\dim_{\mathbb{K}}\widetilde{H}_{i-1}(\Gamma_{h}(I),\mathbb{K}).$
Proof.
Let $K_{\bullet}$ denote the Koszul complex of $X_{1},\ldots,X_{n}$. Namely, $K_{\bullet}$ is the following complex
$0\rightarrow K_{n}\rightarrow\cdots\rightarrow K_{i}\rightarrow\cdots\rightarrow K_{0}\rightarrow 0$
Here, $K_{i}$ is a multi-graded free $R$-module that possesses a base composed of elements $e_{F}$ such that $|F|=i$ and $F\subseteq[n]$. Moreover,
$\deg(e_{F})=\sum_{i\in F}\deg(X_{i})=\sum_{i\in F}\deg(y_{i})\in\mathbb{Z}_{\geq 0}^{m}.$
In particular, $K_{0}=Re_{\emptyset}=R$. Since $K_{\bullet}$ is a multi-graded minimal free resolution of $\mathbb{K}$, we have
$\beta_{i,h}^{R}(I)=\dim_{\mathbb{K}}\mathrm{Tor}_{i}^{R}(\mathbb{K},I)_{h}=\dim_{\mathbb{K}}H_{i}(K_{\bullet}\otimes_{R}I)_{h}.$
On the other hand,
$(K_{i}\otimes I)_{h}=\bigoplus_{|F|=i}(Re_{F}\otimes_{R}I)_{h}=\bigoplus_{|F|=i}(I(\deg(e_{F}))_{h}.$
Note that
$(I(\deg(e_{F}))_{h}=I_{h-\deg(e_{F})}=\left\{\begin{array}[]{ll}\mathbb{K},&\hbox{$F\in\Gamma_{h}(I)$;}\\ 0,&\hbox{otherwise,}\end{array}\right.$
it follows that there is an isomorphism between the complex $K_{\bullet}\otimes_{R}I$ and the complex $\widetilde{C}_{\bullet}(\Gamma_{h}(I),\mathbb{K})[-1]$, yielding the desired formula.
We may look up this lemma as a generalization of [12, Theorem 1.34], which address the same question on monomial ideals in a polynomial ring.
2.3. Simple graphs and minimal cycles
By definition, a simple graph is a finite graph devoid of loops and multiple edges. Let $G$ be a simple graph and denote the vertex set and edge set of $G$ by $V(G)$ and $E(G)$ respectively. Recall that a walk of $G$ of length $q$ is a subgraph $W$ of $G$ such that $E(W)=\{\{i_{0},i_{1}\},\{i_{1},i_{2}\},\ldots,\{i_{q-1},i_{q}\}\}$, where $i_{0},i_{1},\ldots,i_{q}$ are vertices of $G$. The graph $G$ is connected if, for any two vertices $i_{j}$ and $i_{k}$ of $G$, there is a walk between $i_{j}$ and $i_{k}$.
A cycle of $G$ of length $q$ is a subgraph $C$ of $G$ such that
$E(C)=\{\{i_{1},i_{2}\},\{i_{2},i_{3}\},\ldots,\{i_{q-1},i_{q}\},\{i_{q},i_{1}\}\}$
where $i_{1},i_{2},\ldots,i_{q}$ are vertices of $G$ and where $i_{j}\neq i_{k}$ if $j\neq k$. A cycle is called even (resp. odd) if $q$ is even (resp. odd). For a subset $W$ of $V(G)$, the induced subgraph $G[W]$ is the graph with vertex set $W$ and for every pair $j,k\in W$, they are adjacent in $G[W]$ if and only if they are adjacent in $G$. For a vertex $j\in V(G)$, let $G\setminus j$ be the induced subgraph $G[V(G)\setminus\{j\}]$. A cycle of $G$ is called an induced cycle if it is an induced subgraph of $G$. We also call an induced cycle of $G$ to be a minimal cycle of $G$, and use $\mathfrak{t}(G)$ to denote the number of the minimal cycles of $G$.
2.4. Conventions and notions
We set conventions and fix notions that play an important role in the rest of this paper.
Conventions 2.3.
Let $G$ be a simple graph, with its vertices labelled by letters or subscripted letters, such as $u,u_{i,j},v,v_{i,j},w,w_{i,j}$, and so forth. We establish a correspondence between these labels and unit vectors in ${\mathbb{R}}^{V(G)}$ as follows:
$u,u_{i,j},v,v_{i,j},w,w_{i,j},x,y$
correspond to
${\mathbf{u}},{\mathbf{u}}_{i,j},{\mathbf{v}},{\mathbf{v}}_{i,j},{\mathbf{w}},{\mathbf{w}}_{i,j},{\mathbf{x}},{\mathbf{y}}$
respectively. With this notation, we can express ${\mathbb{R}}^{V(G)}$ as the set
$\left\{\sum_{i,j}a_{i,j}{\mathbf{u}}_{i,j}+a{\mathbf{u}}+\sum_{i,j}b_{i,j}{\mathbf{v}}_{i,j}+b{\mathbf{v}}+\sum_{i,j}c_{i,j}{\mathbf{w}}_{i,j}+c{\mathbf{w}}\mid a_{i,j},a,b_{i,j},b,c_{i,j},c\in{\mathbb{R}}\right\}.$
By way of example, $\deg(uv)={\mathbf{u}}_{i}+{\mathbf{v}}_{j}$, and $\deg(u^{2}wu_{1,2})=2{\mathbf{u}}+{\mathbf{w}}+{\mathbf{u}}_{1,2}$, and so on.
We identify a monomial with its multi-degree. For example, we will write $\beta_{i,uv}(\mathbb{K}[G])$ for $\beta_{i,{\mathbf{u}}+{\mathbf{v}}}(\mathbb{K}[G])$. Under this identification, we look up $\mathbb{Z}^{V(H)}$ as a subset of $\mathbb{Z}^{V(G)}$ if $H$ is an induced subgraph of $G$.
Notation 2.4.
Let $G$ be a simple graph satisfying the odd-cycle condition. For every edge $e=\{u,v\}$ of $G$, let $\mathbf{v}_{e}=\mathbf{u}+\mathbf{v}$. We define $D_{H}$ to be the multi-degree $\sum_{e\in E(H)}\mathbf{v}_{e}\in\mathbb{Z}_{\geq 0}^{V(G)}$. Since we identify a monomial with its multi-degree, $D_{H}$ can alternatively be interpreted as the product of all edges of $H$. Here, an edge is considered as the product of its two vertices. Furthermore, we employ $\Theta_{H}$ to represent the product of all the vertices of $H$.
It is straightforward to observe that $\Theta_{H}^{2}$ divides $D_{H}$. Moreover, a simple graph $H$ is a top-Betti graph if and only if $\Theta_{H}$ divides every monomial in $\mathcal{N}_{H}$.
Let $G$ be a simple graph, and we denote the set of all monomials in the edge ring $\mathbb{K}[G]$ of $G$ by $\mathrm{Mon}(\mathbb{K}[G])$.
2.5. Edge cones and canonical modules
Let $G$ be a simple graph with vertex set $V(G)=\{1,\ldots,n\}$ and edge set $E(G)$. For any edge $f=\{i,j\}\in E(G)$, we define $\mathbf{v}_{f}=\mathbf{e}_{i}+\mathbf{e}_{j}$, where $\mathbf{e}_{i}$ represents the $i$th unit vector in $\mathbb{R}^{n}$.
The edge cone of $G$, denoted as $\mathbb{R}_{+}(G)$, is the cone in $\mathbb{R}^{n}$ spanned by $\{\mathbf{v}_{f}|f\in E(G)\}$. Specifically, it can be expressed as:
$\mathbb{R}_{+}(G):=\left\{\sum_{f\in E(G)}{a_{f}\mathbf{v}_{f}}\bigg{|}a_{f}\in\mathbb{R}_{\geq 0}\text{ for all }f\in E(G)\right\}.$
Let us express the edge cone ${\mathbb{R}}_{+}(G)$ in terms of linear inequalities, or equivalently, through its facets. To achieve this, additional notation is required. A non-empty subset $T\subseteq V(G)$ is called an independent set of $G$ if $\{j,k\}\not\in E(G)$ for any $j,k\in T$. For an independent set $T\subseteq V(G)$, we use $N_{G}(T)$ (or simply $N(T)$ for brevity) to denote the neighborhood set of $T$, namely,
$N_{G}(T):=\{j\in V(G)\mid\{j,k\}\in E(G)\mbox{ for some }k\in T\}.$
Given a vertex $i$ of $G$, let $H_{i},H_{i}^{+}$ and $H_{i}^{>}$ denote the following equalities or inequalities respectively:
$H_{i}:x_{i}=0,\qquad H_{i}^{+}:x_{i}\geq 0,\qquad H_{i}^{>}:x_{i}>0.$
Given an independent set $T$ of $G$, let $H_{T},H_{T}^{-}$ and $H_{T}^{<}$ denote the following equalities or inequalities respectively:
$H_{T}:\sum_{i\in T}x_{i}=\sum_{i\in N(T)}x_{i},\qquad H_{T}^{-}:\sum_{i\in T}x_{i}\leq\sum_{i\in N(T)}x_{i},\qquad H_{T}^{<}:\sum_{i\in T}x_{i}<\sum_{i\in N(T)}x_{i}.$
Note that $H_{i}\neq H_{\{i\}}$, since the latter represents the equality $x_{i}=\sum_{j\in N(\{i\})}x_{j}$. If $G$ is bipartite, the facets of ${\mathbb{R}}_{+}(G)$ are given in [20, Proposition 3.6], that we record in the following result for convenience.
Lemma 2.5.
If $G$ is a bipartite connected graph with bipartition $(V_{1},V_{2})$, then $F$ is a facet of ${\mathbb{R}}_{+}(G)$ if and only if either
  1. $(a)$
    $F=H_{i}\cap{\mathbb{R}}_{+}(G)$ for some $i\in V(G)$ with $G\setminus i$ connected; or
  2. $(b)$
    $F=H_{T}\cap{\mathbb{R}}_{+}(G)$ for some $T\varsubsetneq V_{1}$, such that $G[T\sqcup N(T)]$ and $G[V(G)\setminus(T\sqcup N(T))]$ connected.
Let $\mathcal{I}$ denote the set of all vertices satisfying condition (a) and $\mathcal{T}$ denote the set of all independent sets satisfying (b). Then, since $\mathrm{aff}(\mathbb{R}_{+}(G))=H_{V_{1}}=H_{V_{2}}$, we have
($\Delta_{1}$) $\mathbb{R}_{+}(G)=H_{V_{1}}\cap\bigcap_{i\in\mathcal{I}}H_{i}^{+}\cap\bigcap_{T\in\mathcal{T}}H_{T}^{-}$
and
$\mathrm{relint}(\mathbb{R}_{+}(G))=H_{V_{1}}\cap\bigcap_{i\in\mathcal{I}}H_{i}^{>}\cap\bigcap_{T\in\mathcal{T}}H_{T}^{<}.$
For the case when $G$ is non-bipartite, we need to introduce the following notions on graphs.
  • We call a vertex $j$ of $G$ regular if each connected component of $G\setminus j$ contains an odd cycle.
  • We say that an independent set $T$ of $V(G)$ is a fundamental set if
    • the bipartite graph on the vertex set $T\sqcup N(T)$ with the edge set $E(G)\cap\{\{j,k\}\mid j\in T,k\in N(T)\}$ is connected, and
    • either $T\sqcup N(T)=V(G)$ or each of the connected components of the graph $G[V(G)\setminus(T\sqcup N(T))]$ contains an odd cycle.
We herby record a result from [19, Theorem 10.7.6] as follows:
Lemma 2.6.
Let $G$ be a non-bipartite connected simple graph on $V(G)=\{1,\ldots,n\}$. Then ${\mathbb{R}}_{+}(G)$ consists of the elements $(x_{1},\ldots,x_{n})\in{\mathbb{R}}^{n}$ satisfying all the following inequalities:
($\Delta_{2}$) $\begin{split}x_{i}&\geq 0\;\;\text{ for any regular vertex }i;\\ \sum\limits_{i\in N(T)}x_{i}&\geq\sum\limits_{i\in T}x_{i}\;\;\text{ for any fundamental set }T.\end{split}$
It is established by [14, Corollary 2.3] that if $G$ satisfies the odd-cycle condition, then the edge ring $\mathbb{K}[G]$ is normal. Let $\mathcal{S}(G)$ represent the semigroup composed of all $h\in\mathbb{Z}^{n}$ with $X^{h}\in\mathbb{K}[G].$ According to [1, Theorem 6.31], the ideal of $\mathbb{K}[G]$ generated by all monomials $X^{h}$ with $h\in\text{relint}(\mathbb{R}{+}(G))\cap\mathcal{S}(G)$ constitutes the $\mathbb{Z}^{V(G)}$-graded canonical module of the $\mathbb{Z}^{V(G)}$-graded local algebra $\mathbb{K}[G]$, which we denote by $\omega_{\mathbb{K}[G]}$. Then, based on [1, Formula 6.6], we are able to derive the following result.
Lemma 2.7.
Let $G$ be a simple graph satisfying the odd-cycle condition and denote by $p$ the projective dimension of $\mathbb{K}[G]$. Then
$\beta_{p-i,h}(\mathbb{K}[G])=\beta_{i,D_{G}-h}(\omega_{\mathbb{K}[G]})\qquad\mbox{ for all }\quad 0\leq i\leq p\mbox{ and for all }h\in\mathbb{Z}_{\geq 0}^{V(G)}.$
The following two lemmas are important, especially the former, which serves as the cornerstone of the induced-subgraph approach. Although the former is a special case of [15, Corollary 2.5] and the latter is also known, for completeness, we still provide proofs.
Lemma 2.8.
Let $H$ be an induced subgraph of a simple graph $G$. Then for $h\in\mathbb{Z}_{\geq 0}^{V(G)}$ with $\mathrm{supp}(h)\subseteq V(H)$, we have
$\beta_{i,h}(\mathbb{K}[H])=\beta_{i,h}(\mathbb{K}[G])\qquad\mbox{ for all }i\geq 0.$
Proof.
Since $\mathrm{supp}(h)\subseteq V(H)$ we have $\Gamma_{h}(\mathbb{K}[H])=\Gamma_{h}(\mathbb{K}[G])$ by definition. From this fact, the result follows directly in view of Lemma 2.2.
Let $c_{0}(G)$ denote the number of bipartite connected components of a simple $G$.
Lemma 2.9.
Let $G$ be a simple graph with vertex set $V(G)$ and edge set $E(G)$. If $G$ fulfills the odd-condition, then $\mathrm{pdim}(\mathbb{K}[G])=|E(G)|-|V(G)|+c_{0}(G).$
Proof.
First consider the case when $G$ is connected. By [17, Proposition 13.15] and [2, Theorem 6.3.5], $\mathbb{K}[G]$ is Cohen-Macaulay. It follows from [19, Corollary 10.1.21] that
$\mathrm{depth}(\mathbb{K}[G])=\mathrm{dim}(\mathbb{K}[G])=\left\{\begin{array}[]{ll}|V(G)|-1,&\text{if $G$ is bipartite;}\\ |V(G)|,&\text{otherwise.}\end{array}\right.$
According to the Auslander-Buchsbaum formula, as stated, for instance, in [8, Corollary A.4.3], we have
$\mathrm{pdim}(\mathbb{K}[G])=\left\{\begin{array}[]{ll}|E(G)|-|V(G)|+1,&\text{if $G$ bipartite;}\\ |E(G)|-|V(G)|,&\text{otherwise.}\end{array}\right.$
Next, assume $G$ has connected components $G_{1},\ldots,G_{k}$. Then, by [7, Lemma 3.8], we have
$\mathrm{pdim}(\mathbb{K}[G])=\mathrm{pdim}(\mathbb{K}[G_{1}])+\cdots+\mathrm{pdim}(\mathbb{K}[G_{k}]).$
Now, the result follows by combining the above two equalities.
3. top-Betti graphs and second-Betti graphs
In the present section, we provide a combinatorial characterization of a top-Betti graph, and establish a sufficient condition for a graph to be a second-Betti graph. Regarding the definitions of a top-Betti graph and a (strict) second-Betti graph, we refer to Definitions 1.3 and 1.4.
Proposition 3.1.
Let $G$ be a connected graph that fulfills the odd-cycle condition.
$\mathrm{(1)}$ If $G$ is bipartite, then $G$ is a top-Betti graph if and only if for every $x\in V(G)$, one has $c_{0}(G\setminus x)<\deg x.$
$\mathrm{(2)}$ If $G$ is non-bipartite, then $G$ is a top-Betti graph if and only if for every $x\in V(G)$, one has $c_{0}(G\setminus x)<\deg x-1.$
Proof.
We only prove (1) since (2) can be proved similarly. Let $p$ be the projective dimension of $\mathbb{K}[G]$. Assume that there is $x\in V(G)$ such that $c_{0}(G\setminus x)\geq\deg x.$ Then $c_{0}(G\setminus x)=\deg x$. Note that $|E(G\setminus x)|=|E(G)|-\deg x$ and $|V(G\setminus x)|=|V(G)|-1$. Then, we have
$\mathrm{pdim}(\mathbb{K}[G\setminus x])=|E(G)|-\deg x-(|V(G)|-1)+c_{0}(G\setminus x)=p.$
It follows that there is a multi-degree $h\in\mathbb{Z}_{\geq 0}^{V(G)}$ with $\mathrm{supp}(h)\subseteq V(G)\setminus\{x\}$ such that $\beta_{p,h}(\mathbb{K}[G\setminus x]))\neq 0$. Hence, we have $\beta_{p,h}(\mathbb{K}[G])\neq 0$ by Lemma 2.8, which implies $G$ is not a Betti-graph.
Conversely, if $G$ is not a top-Betti graph, then there is a multi-degree $h\in\mathbb{Z}_{\geq 0}^{V(G)}$ with $\mathrm{supp}(h)\subsetneq V(G)$ such that $\beta_{p,h}(\mathbb{K}[G])\neq 0$. Let $x\in V(G)\setminus\mathrm{supp}(h)$. Since $\mathrm{supp}(h)\subseteq V(G)\setminus\{x\}$, we have $\beta_{p,h}(\mathbb{K}[G\setminus x])\neq 0$ by Lemma 2.8. It follows that $\mathrm{pdim}(\mathbb{K}[G\setminus x])=p$, and thus $c_{0}(G\setminus x)=\deg x$, completing the proof.
Corollary 3.2.
Every vertex of a top-Betti graph $G$ has a degree at least two.
Proof.
Assume on the contrary that $\deg x=1$ for some $x\in V(G)$. Then, it follows that $c_{0}(G\setminus x)=1$ if $G$ is bipartite and $c_{0}(G\setminus\{x\})=0$ if $G$ is non-bipartite. This implies $G$ is not a top-Betti graph by Proposition 3.1, a contradiction.
Example 3.3.
Let $\mathbf{K}_{m,n}$ denote the complete bipartite graph with two parts consisting of $m$ and $n$ vertices respectively. Then $\mathbf{K}_{m,n}$ is top-Betti graph if and only if $m\geq 2$ and $n\geq 2$.
Example 3.4.
A 2-connected graph is a graph that remains connected even after the removal of any single vertex and its incident edges. Let $G$ be a 2-connected. Then $G$ is not a top-Betti graph if and only if $G$ is a non-bipartite and there is $x\in G$ such that $G\setminus x$ is a bipartite graph. One such graph is the one consisting of a triangle and a rectangle sharing a common edge.
In general, we can use the following results to find all the strict second-Betti induced subgraphs of a given graph.
Proposition 3.5.
If $G$ is a top-Betti graph, then $G$ is a second-Betti graph
Proof.
From the proof of Proposition 3.1, we see that if $G$ fulfills the odd-cycle condition, then $G$ is a top-Betti graph if and only if $\mathrm{pdim}(\mathbb{K}[H])<\mathrm{pdim}(\mathbb{K}[G])$ for every induced subgraph $H$ of $G$ with $H\neq G$. Let $h$ be a multi-degree such that $\beta_{p-1,h}(\mathbb{K}[G])\neq 0$. We are done if $\mathrm{supp}(h)=V(G)$. Suppose that $\mathrm{supp}(h)\subsetneq V(G)$. Let $H$ be the subgraph of $G$ induced on $\mathrm{supp}(h)$. Then we have $\mathrm{pdim}(\mathbb{K}[H])<\mathrm{pdim}(\mathbb{K}[G])$, and thus $\mathrm{pdim}(\mathbb{K}[H])\leq p-1$. However, we have $\beta_{p-1,h}(\mathbb{K}[H])=\beta_{p-1,h}(\mathbb{K}[G])\neq 0$. Hence, $\mathrm{pdim}(\mathbb{K}[H])=p-1$, as desired.
If $G$ has a vertex $x$ of degree 1, then the edge rings of $G$ and $G\setminus x$ have the same multi-graded Betti numbers. See [21, Lemma 2.1] for the details. This observation yields the following result.
Lemma 3.6.
If $G$ is a strict top-Betti graph, then every vertex of $G$ has a degree at least 2.
4. Betti numbers of $\mathbf{G}_{m}$
In [9], a specific class of simple graphs, designated as $\mathbf{G}_{m}$, was introduced for all $m\geq 1$. It was proved there that the toric ideal of $\mathbb{K}[\mathbf{G}_{m}]$ admits two different initial ideals. One of them is square-free, which implies that it is Cohen-Macaulay, but the other fails to be Cohen-Macaulay. We refer to such a graph as a two-ear graph due to its shape. In this section, we employ the induced-subgraph approach to demonstrate that the toric ideal of $\mathbf{G}_{m}$ possesses exactly the same Betti numbers as the square-free initial ideal of the two initial ideals, as previously mentioned. Furthermore, we present all of its multi-graded Betti numbers.
Definition 4.1.
Let $m\geq 1$ be an integer. Following [9], we define $\mathbf{G}_{m}$ as a simple graph with vertex set
$V(\mathbf{G}_{m})=\{v_{1},v_{2},x_{1},x_{2},y_{1},y_{2}\}\sqcup\{u_{i}~{}|~{}1\leq i\leq m-1\}$
and with edge set
[…]
$\quad\sqcup\{\{v_{1},u_{i}~{}|~{}1\leq i\leq m-1\}\}\sqcup\{\{v_{2},u_{i}~{}|~{}1\leq i\leq m-1\}\}.$
Let $\mathbf{P}_{m}$ denote the induced subgraph $\mathbf{G}_{m}\setminus x_{2}$ of $\mathbf{G}_{m}$. We call a graph $G$ a two-ear graph of type $m$ or an one-ear graph of type $m$ if it is isomorphic to the graph $\mathbf{G}_{m}$ or $\mathbf{P}_{m}$, respectively.
$v_{1}$$v_{2}$$x_{1}$$u_{m-1}$$y_{1}$$u_{2}$$u_{1}$$x_{2}$$y_{2}$$e_{m-1,1}$$e_{m-1,2}$$e_{2,1}$$e_{2,2}$$e_{1,1}$$e_{1,2}$$f_{1}$$g_{1}$$f_{2}$$g_{2}$$f_{3}$$g_{3}$$w$
Figure 1. The two-ear graph $\mathbf{G}_{m}$
Let us denote $\mathbf{G}_{m}$ as $G$. To write out the minimal generators of $\mathrm{in}_{<}(I_{G})$, we label the edges of $G$ as follows: for $i\in[m-1]$, $e_{i,1}=\{v_{1},u_{i}\},e_{i,2}=\{v_{2},u_{i}\}$, and $f_{1}=\{v_{1},x_{1}\},f_{2}=\{x_{1},x_{2}\},f_{3}=\{v_{1},x_{2}\},g_{1}=\{v_{2},y_{1}\},g_{2}=\{y_{1},y_{2}\},g_{3}=\{v_{2},y_{2}\},w=\{x_{1},y_{1}\}$. In [9, Corollary 4.3 ], the minimal set of monomial generators of $\mathrm{in}_{<}(I_{G})$ is specified for a particular monomial order $<$. Note that the monomial order $<$ in this section refers to this order. We record this result in the following lemma for later use.
Lemma 4.2.
There is a monomial order $<$ such that the minimal set of monomial generators of $\mathrm{in}_{<}(I_{G})$ is given by $\mathcal{M}=\mathcal{M}_{1}\sqcup\mathcal{M}_{2}\sqcup\mathcal{M}_{3}\sqcup\mathcal{M}_{4}$, where
  1. $(i)$
[…]
  2. $(ii)$
    $\mathcal{M}_{2}=\left\{f_{1}f_{2}g_{1}g_{2}\right\}$,
  3. $(iii)$
[…]
  4. $(iv)$
[…]
We sort the set of minimal generators of $\mathrm{in}_{<}(I_{G})$ as follows:
$\begin{split}&e_{1,1}e_{2,2}\prec e_{1,1}e_{3,2}\prec\cdots\prec e_{1,1}e_{m-1,2}\prec e_{2,1}e_{3,2}\prec e_{2,1}e_{4,2}\prec\cdots\prec e_{2,1}e_{m-1,2}\prec\cdots\\ &\prec e_{m-2,1}e_{m-1,2}\prec g_{1}f_{2}e_{1,1}\prec g_{1}f_{2}e_{2,1}\prec\cdots\prec g_{1}f_{2}e_{m-1,1}\prec f_{1}f_{2}g_{1}g_{2}\prec g_{2}f_{1}e_{m-1,2}\\ &\prec g_{2}f_{1}e_{m-2,2}\prec\cdots\prec g_{2}f_{1}e_{1,2}.\end{split}$
For brevity, we label each term in this sequence consecutively: $e_{1,1}e_{2,2}$ as $u_{1}$, $e_{1,1}e_{3,2}$ as $u_{2}$, and so forth, culminating with $g_{2}f_{1}e_{1,2}$ designated as $u_{r}$. Here, $r=\frac{m(m+1)}{2}$.
Lemma 4.3.
The monomial ideal $I=\mathrm{in}_{<}(I_{G})$ has regular quotients with respect to the above order. More precisely, the colon ideal $(u_{1},\ldots,u_{j-1}):u_{j}$ is generated by a regular sequence of length $k_{j}$ for $j=2,\ldots,\frac{m(m+1)}{2}$, where
$(k_{2},\ldots,k_{\frac{m(m+1)}{2}+1})=(1,2,\ldots,m-3,1,2,\ldots,m-3,2,\ldots,m-3,\ldots,m-3),$
$(k_{\frac{m(m+1)}{2}+2},\ldots,k_{\frac{m(m+1)}{2}})=(\underbrace{m-2,m-2,\ldots,m-2}_{m-1},\ldots,\underbrace{m-1,m-1,\ldots,m-1}_{m}).$
Proof.
Let $I_{j-1}=(u_{1},\ldots,u_{j-1})$ be the ideal generated by the first $j-1$ elements of the sequence. Now, we consider several cases based on the form of $u_{j}$:
1. If $u_{j}=e_{a,1}e_{b,2}$ for some $11\leq a<b\leq m-$ 1, then the colon ideal $I_{j-1}:u_{j}$ is given by
$I_{j-1}:u_{j}=(e_{1,1},e_{2,1},\ldots,e_{a-1,1},e_{a+1,2},e_{a+2,2},\ldots,e_{b-1,2}).$
2. If $u_{j}=g_{1}f_{2}e_{a,1}$ for some $11\leq a\leq m-$ 1, the colon ideal becomes
$I_{j-1}:u_{j}=(e_{1,1},e_{2,1},\ldots,e_{a-1,1},e_{a+1,2},\ldots,e_{m-1,2}).$
3. If $u_{j}=f_{1}f_{2}g_{1}g_{2}$, the colon ideal simplifies to
$I_{j-1}:u_{j}=(e_{1,1},e_{2,1},\ldots,e_{m-1,1}).$
4. Finally, if $u_{j}=g_{2}f_{1}e_{a,2}$ for some $11\leq a\leq m-$ 1, the colon ideal is
$I_{j-1}:u_{j}=(g_{1}f_{2},e_{1,1},e_{2,1},\ldots,e_{a-2,1},e_{a,2},e_{a+1,2},\ldots,e_{m-1,2}).$
By enumerating the minimal generators of these colon ideals, the desired result follows.
Lemma 4.4.
Let $G$ be the graph as described above. Then there is a monomial order $<$ such that $\beta_{i}(\mathrm{in}_{<}(I_{G}))\leq(i+1)\binom{m+1}{i+2}$ for all $i\geq 0$. Thus,
$\beta_{i}(\mathbb{K}[G])\leq\beta_{i}(\mathbb{K}[E(G)]/\mathrm{in}_{<}(I_{G}))\leq i\binom{m+1}{i+1}.$
Proof.
Based on Lemma 4.3 and considering Formula (1), we obtain,
$\beta_{i}(\mathrm{in}_{<}(I_{G}))$ $\leq\sum_{k=1}^{m-1}(k+1)\binom{k}{i}+\binom{0}{i}=(i+1)\sum_{k=1}^{m-1}\binom{k+1}{i+1}+\binom{0}{i}$
$=(i+1)\sum_{k=0}^{m-1}\binom{k+1}{i+1}=(i+1)\binom{m+1}{i+2}.$
Here, the third-to-last equality follows from the identity $k\binom{n}{k}=n\binom{n-1}{k-1}$, and the last equality follows from the identity $\sum\limits_{j=0}^{n}\binom{j}{k}=\binom{n+1}{k+1}$ for $k\geq 0$. From this bound and Lemma 2.1, the second assertion is immediate.
Notice that a two-ear graph can have top-Betti induced subgraphs of three types: two-ear graphs, one-ear graphs, and complete bipartite graphs of the form $\mathbf{K}_{2,d}$ for some $d\geq 2$. Next, we will determine the elements in the top support for the three types of graphs by considering the canonical module of their respective edge rings. The most complex scenario is when $H$ is a two-ear graph, for which we will provide a detailed proof. When $H$ is an one-ear graph or a complete bipartite graph $\mathbf{K}_{2,d}$, the proofs are similar with only minor differences, so we will only outline the proofs for those cases.
4.1. The composition of $\mathcal{N}_{\mathbf{G}_{m}}$
Notice that when $m=1$, the two-ear graph $G_{1}$ is the compact graph $B^{0}_{(1):(1)}$, which we discuss in the next section. Its top support is as shown in Lemma 5.3. Therefore, in this subsection, we assume $m\geq 2$ and, for brevity, use $G$ to represent $\mathbf{G}_{m}$.
It is easy to see that $|V(G)|=m+5$. It follows from Conventions 2.3 that:
${\mathbb{R}}^{V(G)}=\{\sum\limits_{i=1}^{m-1}{a_{i}{\mathbf{u}}_{i}}+\sum\limits_{i=1}^{2}({b_{i}{\mathbf{v}}_{i}+c_{i}{\mathbf{x}}_{i}+d_{i}{\mathbf{y}}_{i}})\mid a_{i},b_{i},c_{i},d_{i}\in{\mathbb{R}}\mbox{ for all }i\},$
where ${\mathbf{u}}_{i},{\mathbf{v}}_{i},{\mathbf{x}}_{i},{\mathbf{y}}_{i}$ are the unit vectors of ${\mathbb{R}}^{V(G)}$, each ${\mathbf{u}}_{i}$ (resp. ${\mathbf{v}}_{i},{\mathbf{x}}_{i},{\mathbf{y}}_{i}$) corresponds to $u_{i}$ (resp. $v_{i},x_{i},y_{i}$) for all $i$.
Recall that an integral vector in $\mathrm{relint}({\mathbb{R}}_{+}(G))$ is minimal if it cannot be written as the sum of a vector in $\mathrm{relint}({\mathbb{R}}_{+}(G))\cap\mathcal{S}(G)$ and a nonzero vector of ${\mathbb{R}}_{+}(G)\cap\mathcal{S}(G)$. It is clear that a minimal vector of $\mathrm{relint}({\mathbb{R}}_{+}(G))$ corresponds to a minimal generator of the canonical module of $\mathbb{K}[G]$. In what follows, we will construct $m$ integral vectors in ${\mathbb{R}}^{V(G)}$ and then show that they are minimal vectors of $\mathrm{relint}({\mathbb{R}}_{+}(G))\cap\mathcal{S}(G)$. The construction is as follows:
For $\ell=1,\ldots,m$, let
$\alpha_{\ell}:=\sum\limits_{i=1}^{m-1}{\mathbf{u}}_{i}+\sum\limits_{i=1}^{2}({\mathbf{x}}_{i}+{\mathbf{y}}_{i})+\ell{\mathbf{v}}_{1}+(m+1-\ell){\mathbf{v}}_{2}.$
We now verify that $\alpha_{\ell}\in\mathrm{relint}({\mathbb{R}}_{+}(G))$ for all $\ell=1,\ldots,m$. First, we have the following easy observations.
  • The vertices of $G$ are all regular vertices.
  • An independent subset $T$ of $V(G)$ is fundamental if and only if $T$ is one of the following sets:
    1. $(i)$
      $\{u_{1},u_{2},\cdots,u_{m-1},x_{1},y_{2}\};$
    2. $(ii)$
      $\{u_{1},u_{2},\cdots,u_{m-1},x_{2},y_{1}\};$
    3. $(iii)$
      $\{v_{1},v_{2}\};$
    4. $(iv)$
      $\{v_{1},y_{1}\};$
    5. $(v)$
      $\{v_{2},x_{1}\}.$
Hence, it follows from ($\Delta_{2}$) that a vector of ${\mathbb{R}}^{V(G)}$ of the form:
$\sum\limits_{i=1}^{m-1}{a_{i}{\mathbf{u}}_{i}}+\sum\limits_{i=1}^{2}({b_{i}{\mathbf{v}}_{i}+c_{i}{\mathbf{x}}_{i}+d_{i}{\mathbf{y}}_{i}})$
belongs to ${\mathbb{R}}_{+}(G)$ if and only if the following inequalities are satisfied:
  1. (1)
    $a_{i}\geq 0$ for any $11\leq i\leq m-$ 1 ;
  2. (2)
    $b_{i},c_{i},d_{i}\geq 0$ for any $1\leq i\leq 2$ ;
  3. (3)
    $b_{1}+b_{2}+c_{2}+d_{1}\geq\sum\limits_{i=1}^{m-1}a_{i}+c_{1}+d_{2}$ ;
  4. (4)
    $b_{1}+b_{2}+c_{1}+d_{2}\geq\sum\limits_{i=1}^{m-1}a_{i}+c_{2}+d_{1}$ ;
  5. (5)
    $\sum\limits_{i=1}^{m-1}a_{i}+c_{1}+c_{2}+d_{1}+d_{2}\geq b_{1}+b_{2}$ ;
  6. (6)
    $\sum\limits_{i=1}^{m-1}a_{i}+b_{2}+c_{1}+c_{2}+d_{2}\geq b_{1}+d_{1}$ ;
  7. (7)
    $\sum\limits_{i=1}^{m-1}a_{i}+b_{1}+c_{2}+d_{1}+d_{2}\geq b_{2}+c_{1}$
It is straightforward to check that $\alpha_{\ell}$ satisfies these inequalities, with strict inequalities holding for each $\alpha_{\ell}$. It is obvious that the sum of all components of $\alpha_{\ell}$ is even, and according to [19, Proposition 3.4], we have $\alpha_{\ell}\in\mathrm{relint}({\mathbb{R}}_{+}(G))\cap\mathcal{S}(G)$.
Next, we show that $\alpha_{\ell}$ is a minimal vector in $\mathrm{relint}({\mathbb{R}}_{+}(G))\cap\mathcal{S}(G)$.
Suppose on the contrary that $\alpha_{\ell}=\alpha^{\prime}+\alpha^{\prime\prime}$ for some $\alpha^{\prime}\in\mathrm{relint}({\mathbb{R}}_{+}(G))\cap\mathcal{S}(G)$ and $\alpha^{\prime\prime}\in{\mathbb{R}}_{+}(G)\cap\mathcal{S}(G)\setminus\{{\bf 0}\}$. Write
$\alpha^{\prime}=\sum\limits_{i=1}^{m-1}{a^{\prime}_{i}{\mathbf{u}}_{i}}+\sum\limits_{i=1}^{2}({b^{\prime}_{i}{\mathbf{v}}_{i}+c^{\prime}_{i}{\mathbf{x}}_{i}+d^{\prime}_{i}{\mathbf{y}}_{i}}),$
$\alpha^{\prime\prime}=\sum\limits_{i=1}^{m-1}{a^{\prime\prime}_{i}{\mathbf{u}}_{i}}+\sum\limits_{i=1}^{2}({b^{\prime\prime}_{i}{\mathbf{v}}_{i}+c^{\prime\prime}_{i}{\mathbf{x}}_{i}+d^{\prime\prime}_{i}{\mathbf{y}}_{i}}).$
Note that $\mathrm{relint}({\mathbb{R}}_{+}(G))\cap\mathcal{S}(G)\subseteq{\mathbb{R}}_{+}(G)\cap\mathcal{S}(G)\subseteq\mathbb{Z}^{V(G)}$. In view of the inequalities (1) and (2), we see that $a^{\prime}_{i},c^{\prime}_{i},d^{\prime}_{i}\geq 1$ and $a^{\prime\prime}_{i},c^{\prime\prime}_{i},d^{\prime\prime}_{i}\geq 0$ for all $i,j$. Hence, $a^{\prime}_{i}=c^{\prime}_{i}=d^{\prime}_{i}=1$ and $a^{\prime\prime}_{i}=c^{\prime\prime}_{i}=d^{\prime\prime}_{i}=0$ for all possible $i,j$. Because of (3) ,we have $b^{\prime}_{1}+b^{\prime}_{2}\geq m$. Hence $b^{\prime\prime}_{1}+b^{\prime\prime}_{2}\leq 1$. Because of (6) and (7), we have $b^{\prime\prime}_{1}=b^{\prime\prime}_{2}$. But from (2) we know $b^{\prime\prime}_{1},b^{\prime\prime}_{2}\geq 0$, so we have $b^{\prime\prime}_{1}=b^{\prime\prime}_{2}=0$.
Hence, $\alpha^{\prime\prime}=0.$ This is a contradiction, which shows that $\alpha_{\ell}$ is a minimal vector in $\mathrm{relint}({\mathbb{R}}_{+}(G))\cap\mathcal{S}(G)$ for $\ell=1,\ldots,m$.
Actually, $\alpha_{\ell}$ for $\ell=1,\ldots,m$ are all minimal vector in $\mathrm{relint}({\mathbb{R}}_{+}(G))\cap\mathcal{S}(G)$, as shown below.
Lemma 4.5.
Let $G$ represent the graph $\mathbf{G}_{m}$ with the labeling as shown in Figure 1. Then, we have
$\mathrm{pdim}(\mathbb{K}[G])=\mathrm{pdim}(\mathbb{K}[E(G)]/\mathrm{in}_{<}(I_{G}))=m.$
Furthermore,
$\mathcal{N}_{G}=\{\Theta_{G}x_{1}y_{1}v_{1}^{m-\ell}v_{2}^{\ell-1}\mid\ell=1,\ldots,m\}.$
Proof.
The first statement is a consequence of Lemma 2.9 combined with [5, Corollary 2.7]. For the second one, note that since $\mathrm{pdim}(\mathbb{K}[G])=m$, the top total Betti number of $\mathbb{K}[G]$ is at most $m$ by Lemma 4.4. Thus, the cardinality of the set $\mathcal{N}_{G}$ is bounded above by $m$, i.e., $|\mathcal{N}_{G}|\leq m$. On the other hand, from previous discussions, we know that $|\mathcal{N}_{G}|\geq m$. Now, considering that $D_{G}=\Theta_{G}^{2}x_{1}y_{1}(v_{1}v_{2})^{m-1}$ and applying Lemma 2.7, we can conclude that $\mathcal{N}_{G}=\{\Theta_{G}x_{1}y_{1}v_{1}^{m-\ell}v_{2}^{\ell-1}\mid\ell=1,\ldots,m\}$, as desired.
Since the regularity of Cohen-Macaulay module is determined by its top Betti numbers, the following result is immediate.
Corollary 4.6.
the regularity of the edge ring $\mathbb{K}[G_{m}]$ is equal to $3$.
4.2. The composition of $\mathcal{N}_{\mathbf{P}_{m}}$.
Let $m\geq 2$ be an integer. Let $G^{\prime}$ (resp. $G^{\prime\prime}$) denote the induced subgraph $\mathbf{G}_{m}\setminus x_{2}$ (resp. $\mathbf{G}_{m}\setminus y_{2}$) of $\mathbf{G}_{m}$. According to Definition 4.1, both $G^{\prime}$ and $G^{\prime\prime}$ are isomorphic to the one-ear graph $\mathbf{P}_{m}$.
It follows from Conventions 2.3 that:
${\mathbb{R}}^{V(G^{\prime})}=\{\sum\limits_{i=1}^{m-1}{a_{i}{\mathbf{u}}_{i}}+\sum\limits_{i=1}^{2}({b_{i}{\mathbf{v}}_{i}+d_{i}{\mathbf{y}}_{i}})+c_{1}{\mathbf{x}}_{1}\mid a_{i},b_{i},c_{1},d_{i}\in{\mathbb{R}}\mbox{for all }i\},$
here, ${\mathbf{u}}_{i},{\mathbf{v}}_{i},{\mathbf{x}}_{1},{\mathbf{y}}_{i}$ correspond the vertices of $G^{\prime}$ in a natural way. We could show the following vectors employing the same process as in the case of a two-ear graph:
$\alpha_{\ell}:=\sum\limits_{i=1}^{m-1}{\mathbf{u}}_{i}+{\mathbf{x}}_{1}+2{\mathbf{y}}_{1}+{\mathbf{y}}_{2}+\ell{\mathbf{v}}_{1}+(m-\ell){\mathbf{v}}_{2},\ \ell=1,\ldots,m-1$
are the minimal vectors of $\mathrm{relint}({\mathbb{R}}_{+}(G^{\prime}))\cap\mathcal{S}(G^{\prime})$.
It is clear that $D_{G^{\prime}}=\Theta_{G^{\prime}}^{2}y_{1}v_{1}^{m-2}v_{2}^{m-1}$. The following results are now immediate from Lemma 2.9 and Lemma 2.7.
Lemma 4.7.
Assume that $G^{\prime}$ is labelled as in Figure 1. Then
$\mathrm{pdim}(\mathbb{K}[G^{\prime}])=m-1\mbox{ and }\{\Theta_{G^{\prime}}v_{1}^{m-\ell-1}v_{2}^{\ell}\mid\ell=1,\ldots,m-1\}\subseteq\mathcal{N}_{G^{\prime}}.$
4.3. The composition of $\mathcal{N}_{\mathbf{K}_{2,d}}$.
For certain values of $d\geq 2$, a complete bipartite graph of the form $\mathbf{K}_{2,d}$ may also be an induced subgraph of graph $G$.
It is easy to see that $|V(\mathbf{K}_{2,d})|=d+2$. Assume that $\mathbf{K}_{2,d}$ has two parts $\{v_{1},v_{2}\}$ and $\{u_{1},\ldots,u_{d}\}$. It follows from Conventions 2.3 that:
${\mathbb{R}}^{V(\mathbf{K}_{2,d})}=\{\sum\limits_{i=1}^{d}{a_{i}{\mathbf{u}}_{i}}+{b_{1}{\mathbf{v}}_{1}}+{b_{2}{\mathbf{v}}_{2}}\mid a_{i},b_{i}\in{\mathbb{R}}\mbox{ for all }i\},$
where ${\mathbf{u}}_{i},{\mathbf{v}}_{i}$ are the unit vectors of ${\mathbb{R}}^{V(\mathbf{K}_{2,d})}$, each ${\mathbf{u}}_{i}$ (resp. ${\mathbf{v}}_{i}$) corresponds to $u_{i}$ (resp. $v_{i}$) for all $i$.
In what follows, we will construct $(d-1)$ integral vectors in ${\mathbb{R}}^{V(\mathbf{K}_{2,d})}$ and then show that they are minimal vectors of $\mathrm{relint}({\mathbb{R}}_{+}(\mathbf{K}_{2,d}))\cap\mathcal{S}(\mathbf{K}_{2,d})$. The construction is as follows:
For $\ell=1,\ldots,d-1$, let
$\alpha_{\ell}:=\sum\limits_{i=1}^{d}{\mathbf{u}}_{i}+\ell{\mathbf{v}}_{1}+(d-\ell){\mathbf{v}}_{2}.$
According to Lemma 2.5, all the facets of ${\mathbb{R}}_{+}(\mathbf{K}_{2,d})$ are exactly $H_{\{v_{1}\}}\cap{\mathbb{R}}_{+}(\mathbf{K}_{2,d})$ and $H_{\{v_{2}\}}\cap{\mathbb{R}}_{+}(\mathbf{K}_{2,d})$.
Hence, it follows from ($\Delta_{1}$) that a vector of ${\mathbb{R}}^{V(\mathbf{K}_{2,d})}$ of the form:
$\sum\limits_{i=1}^{d}{a_{i}{\mathbf{u}}_{i}}+b_{1}{\mathbf{v}}_{1}+b_{2}{\mathbf{v}}_{2}$
belongs to ${\mathbb{R}}_{+}(\mathbf{K}_{2,d})$ if and only if the following formulas are satisfied:
  1. (1)
    $a_{i}\geq 0$ for any $1\leq i\leq d$ ;
  2. (2)
    $b_{i}\geq 0$ for any $i=1,2$ ;
  3. (3)
    $\sum\limits_{i=1}^{d}a_{i}\geq b_{1}$ ;
  4. (4)
    $\sum\limits_{i=1}^{d}a_{i}\geq b_{2}$
It is routine to prove that $\alpha_{\ell}$ is a minimal vector in $\mathrm{relint}({\mathbb{R}}_{+}(G))\cap\mathcal{S}(G)$ for $\ell=1,\ldots,d-1$. Denote $H=\mathbf{K}_{2,d}$. Then it is clear that $D_{H}=\Theta_{H}^{2}(v_{1}v_{2})^{d-2}$. Therefore, the subsequent result follows directly from Lemma 2.9 and Lemma 2.7.
Lemma 4.8.
Denote $H=\mathbf{K}_{2,d}$, then
$\mathrm{pdim}(\mathbb{K}[H])=d-1\mbox{ and }\{\Theta_{H}v_{1}^{d-\ell-1}v_{2}^{\ell-1}\mid\ell=1,\ldots,d-1\}\subseteq\mathcal{N}_{H}.$
Thus far, we have computed the projective dimension of the edge rings for the three types of graphs under consideration. Moreover, we have determined the top support of a two-ear graph. As for the one-ear graph and the complete bipartite graph $\mathbf{K}_{2,d}$, we have respectively obtained two particular subsets of their top supports. We are now ready to present the first main result of this section.
Theorem 4.9.
Let $G$ be a two-ear graph of type $m$. Then, for any $0\leq i\leq m$, we have
$\beta_{i}(\mathbb{K}[G])=\beta_{i}(\mathbb{K}[E(G)]/\mathrm{in}_{<}(I_{G})).$
Proof.
From Lemma 4.5, it can be seen that $\mathrm{pdim}(\mathbb{K}[G])=\mathrm{pdim}(\mathbb{K}[E(G)]/\mathrm{in}_{<}(I_{G}))=m$. We only need to prove the two Betti numbers are identical.
If $i=0$, then $\beta_{i}(\mathbb{K}[G])=\beta_{i}(\mathbb{K}[E(G)]/\mathrm{in}_{<}(I_{G}))=1.$ Suppose now that $1\leq i\leq m$. We now identify all top-Betti induced subgraphs of $G$ whose edge rings have a projective dimension of $i$. Note that $\mathrm{pdim}(\mathbb{K}[\mathbf{G}_{i}])=\mathrm{pdim}(\mathbb{K}[\mathbf{P}_{i+1}])=\mathrm{pdim}(\mathbb{K}[\mathbf{K}_{2,i+1}])=i$(Notice that $G_{1}$ is the compact graph $B^{0}_{(1);(1)}$, and its projective dimension is 1, as stated in Lemma 6.1). Observe that $G$ has $\binom{m-1}{i-1}$ induced subgraphs that are isomorphic to $\mathbf{G}_{i}$, 2$\binom{m-1}{i}$ induced subgraphs that are isomorphic to $\mathbf{P}_{i+1}$, and $\binom{m-1}{i+1}$ induced subgraphs which are isomorphic to $\mathbf{K}_{2,i+1}$. Hence, there are at least $r$ induced subgraphs $H$ of $G$ such that $\mathrm{pdim}(\mathbb{K}[H])=i$. Here, $r$ is given by
$r=\binom{m-1}{i-1}+2\binom{m-1}{i}+\binom{m-1}{i+1}=\binom{m+1}{i+1}.$
Let $\mathcal{N}_{i}(G)$ denote the union of all $\mathcal{N}_{H}$ such that $H$ is a top-Betti induced subgraph of $G$ satisfying $\mathrm{pdim}(\mathbb{K}[H])=i$.
Observe that if $H$ is isomorphic to $\mathbf{G}_{i}$, then according to Lemma 4.5 and Lemma 5.3, $\mathcal{N}_{H}$ comprises $i$ monomials. If $H$ is isomorphic to either $\mathbf{P}_{i+1}$ or $\mathbf{K}_{2,i+1}$, then the corresponding lemma (Lemma 4.7 for $\mathbf{P}_{i+1}$ and Lemma 4.8 for $\mathbf{K}_{2,i+1}$) ensures that $\mathcal{N}_{H}$ contains at least $i$ monomials.
Hence,
$|\mathcal{N}_{i}(G)|\geq ir=i\binom{m+1}{i+1}.$
This implies
$i\binom{m+1}{i+1}\leq\sum_{\alpha\in\mathcal{N}_{i}(G)}\beta_{i,\alpha}(\mathbb{K}[G])\leq\beta_{i}(\mathbb{K}[G])\leq\beta_{i}(\mathbb{K}[E(G)]/\mathrm{in}_{<}(I_{G}))\leq i\binom{m+1}{i+1},$
as desired.
From the proof of Proposition 4.9, we can immediately draw the following two corollaries. These represent an improvement over Lemmas 4.7 and 4.8, enabling us to enumerate all elements of top support for all one-ear graphs and complete bipartite graphs $\mathbf{K}_{2,d}$.
Corollary 4.10.
Denote by $H$ the graph $G^{\prime}$ described above, then
$\mathrm{(1)}$ $\mathcal{N}_{H}=\{\Theta_{H}v_{1}^{m-\ell-1}v_{2}^{\ell}\mid\ell=1,\ldots,m-1\};$
$\mathrm{(2)}$ the regularity of the edge ring $\mathbb{K}[H]$ is equal to $3$.
Corollary 4.11.
Denote by $H$ the graph $\mathbf{K}_{2,d}$ described above, then
$\mathrm{(1)}$ $\mathcal{N}_{H}=\{\Theta_{H}v_{1}^{d-\ell-1}v_{2}^{\ell-1}\mid\ell=1,\ldots,d-1\};$
$\mathrm{(2)}$ the regularity of the edge ring $\mathbb{K}[H]$ is equal to $1$.
We now present the second main result of this section. As in the proof of Theorem 4.9, we use $\mathcal{N}_{i}(G)$ to denote the disjoint union of all $\mathcal{N}_{H}$ such that $H$ is a top-Betti induced subgraph of $G$ satisfying $\mathrm{pdim}(\mathbb{K}[H])=i$.
Theorem 4.12.
Let $G$ be a two-ear graph of type $m$. Then for any monomial $\alpha$ in $\mathbb{K}[G]$ and for any integer $i\geq 1$, we have
$\beta_{i,\alpha}(\mathbb{K}[G])=\left\{\begin{array}[]{ll}1,&\hbox{$\alpha\in\mathcal{N}_{i}(G)$;}\\ 0,&\hbox{otherwise.}\end{array}\right.$
In particular, there exists a square-free initial ideal $\mathrm{in}_{<}(I_{G})$ of $I_{G}$ such that
$\beta_{i,j}(\mathrm{in}_{<}(I_{G}))=\beta_{i,j}(I_{G})=|\{\alpha\in\mathcal{N}_{i+1}(G)\mid\mathrm{deg}(\alpha)=2j\}|$ for all $i,j$.
5. Compact graphs revisited
The primary objective of this section is to determine the elements of the top support $\mathcal{N}_{H}$ for every compact graph $H$, thereby facilitating the application of the induced-subgraph approach for computing the Betti numbers of their edge rings in the subsequent section. To accomplish this, we revisit the precise definition and classification of compact graphs as presented in [21].
Definition 5.1.
Let $G$ be a connected simple graph where every vertex has a degree greater than one. We call $G$ to be a compact graph if it does not contain any even cycles and satisfies the odd-cycle condition.
We call a vertex of a compact graph to be a big vertex if it has a degree greater than 2. According to [21, Proposition 3.3], there are at most three big vertices in a compact graph. Therefore, compact graphs containing more than one cycle can be categorized into three classes, each determined by the number of big vertices. Specifically, we say a compact graph falls into type $i$ if it possesses $i$ big vertices for $i=1,2,3.$
We present a complete classification of compact graphs as follows, which is given in [21, Theorem 2.12]. A compact graph of type 1 is a finite collection of odd cycles that share a vertex.
A compact graph of type 2 consists of two disjoint compact graphs of type 1, where the two big vertices are connected either by an edge or by an edge as well as a path of even length.
A compact graph of type 3 consists of three disjoint compact graphs of type 1, where every pair of big vertices is connected by an edge.
Suppose $\underline{p}=(p_{1},\ldots,p_{m})$, $\underline{q}=(q_{1},\ldots,q_{n})$ and $\underline{r}=(r_{1},\ldots,r_{k})$ are positive integral vectors with dimensions $m,n$ and $k$ respectively. We denote a compact graph of type 1, where the odd cycles have lengths $12p_{1}+1,\ldots,2p_{m}+$ 1 respectively, as $\mathbf{A}_{\underline{p}}$ or $\mathbf{A}_{(p_{1},\ldots,p_{m})}$.
By $\mathbf{B}_{\underline{p}:\underline{q}}^{0}$ we mean a compact graph of type 2 where the two disjoint compact graphs of type 1 that compose it are $\mathbf{A}_{\underline{p}}$ and $\mathbf{A}_{\underline{q}}$ and where the two big vertices are connected by an edge. Furthermore, if $s>0$ is an even number, then $\mathbf{B}_{\underline{p};\underline{q}}^{s}$ represents the graph obtained by appending a path of length $s$ connecting two big vertices to $\mathbf{B}_{\underline{p}:\underline{q}}^{0}$.
A compact graph of type 3 is denoted by $\mathbf{C}_{\underline{p}:\underline{q}:\underline{r}}$ if the three disjoint compact graphs of type 1 that make up it are $\mathbf{A}_{\underline{p}}$, $\mathbf{A}_{\underline{q}}$ and $\mathbf{A}_{\underline{r}}$ respectively.
By Proposition 3.1, it is easy to see that all compact graphs are top-Betti graphs, which, according to Proposition 3.5, implies that they are also second-Betti graphs. We shall revisit the findings from [21] pertaining to the minimal generators of the edge cone of compact graphs, and subsequently establish the compositions of top supports of three types of compact graphs, respectively.
5.1. Type one
Given positive integers $m\geq 2$, as well as the positive integral tuple $\underline{p}=(p_{1},\ldots,p_{m})$, we denote the graph $\mathbf{A}_{\underline{p}}$ as $A$ for brevity. Here, $p_{i}$ are all positive integers. We may assume that $A$ has vertex set
$V(A)=\{u\}\sqcup\{u_{i,j}\mid 1\leq i\leq m,1\leq j\leq 2p_{i}\}$
and edge set
[…]
$\sqcup\{\{u,u_{i,1}\},\{u,u_{i,2p_{i}}\}\mid 1\leq i\leq m\}.$
Recall again that an integral vector in $\mathrm{relint}({\mathbb{R}}_{+}(G))$ is minimal if it cannot be written as the sum of a vector in $\mathrm{relint}({\mathbb{R}}_{+}(G))\cap\mathcal{S}(G)$ and a nonzero vector of ${\mathbb{R}}_{+}(G)\cap\mathcal{S}(G)$. We showed in [21, Section 6] that the following vectors
$\alpha_{\ell}:=\sum\limits_{i=1}^{m}\sum\limits_{j=1}^{2p_{i}}{\mathbf{u}}_{i,j}+2\ell{\mathbf{u}},\ \ell=1,\ldots,m-1$
are all the minimal vectors of $\mathrm{relint}({\mathbb{R}}_{+}(A))\cap\mathcal{S}(A)$.
It is clear that $D_{A}=\Theta_{A}^{2}u^{2m-2}$. Therefore, the subsequent result follows directly from Lemma 2.7.
Lemma 5.2.
Let $H$ be a compact graph of type $1$ labelled as previously. Then
$\mathcal{N}_{H}=\{\Theta_{H}u^{2(m-\ell)-1}\mid\ell=1,\ldots,m-1\}.$
5.2. type two
Given positive integers $m,n$ and an even integer $s\geq 0$, as well as the positive integral tuples $\underline{p}=(p_{1},\ldots,p_{m})$ and $\underline{q}=(q_{1},\ldots,q_{n})$, we denote the graph $\mathbf{B}^{s}_{\underline{p}:\underline{q}}$ as $B$ for brevity. Here, $p_{i},q_{i}$ are all positive integers. We may assume that $B$ has vertex set
$V(B)=$$\{u,v\}\sqcup\{w_{1},\ldots,w_{s-1}\}$
$\sqcup\{u_{i,j}\mid 1\leq i\leq m,1\leq j\leq 2p_{i}\}\sqcup\{v_{i,j}\mid 1\leq i\leq n,1\leq j\leq 2q_{i}\}$
and edge set
[…]
$\sqcup\{\{u,u_{i,1}\},\{u,u_{i,2p_{i}}\}\mid 1\leq i\leq m\}$
$\sqcup\{\{u,v\},\{u,w_{1}\},\{v,w_{s-1}\}\}\cup\{\{w_{i},w_{i+1}\}\mid 1\leq i\leq s-2\}$
$\sqcup\{\{v_{i,j},v_{i,j+1}\}\mid 1\leq i\leq n,1\leq j\leq 2q_{i}-1\}$
$\sqcup\{\{v,v_{i,1}\},\{v,v_{i,2q_{i}}\}\mid 1\leq i\leq n\}.$
Let $B^{0}$ and $B^{s}$ denote the compact graph $\mathbf{B}_{\underline{p}:\underline{q}}^{0}$ and $\mathbf{B}_{\underline{p}:\underline{q}}^{s}$ respectively, where $s\geq 2$ is an even number. We may write by Conventions 2.3
${\mathbb{R}}^{V(B^{0})}=\{\sum\limits_{i=1}^{m}\sum\limits_{j=1}^{2p_{i}}a_{i,j}{\mathbf{u}}_{i,j}+\sum\limits_{i=1}^{m}\sum\limits_{j=1}^{2p_{i}}b_{i,j}{\mathbf{v}}_{i,j}+a{\mathbf{u}}+b{\mathbf{v}}\mid\mbox{ all }a_{i,j},b_{i,j},a,b\in{\mathbb{R}}\}.$
We showed in [21, Section 5] that the following vectors
$\alpha_{\ell}:=\sum\limits_{i=1}^{m}\sum\limits_{j=1}^{2p_{i}}{\mathbf{u}}_{i,j}+\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{2q_{i}}{\mathbf{v}}_{i,j}+{\mathbf{v}}+(2\ell+1){\mathbf{u}},\quad\ell=0,\ldots,m-1$
and
$\beta_{\ell}:=\sum\limits_{i=1}^{m}\sum\limits_{j=1}^{2p_{i}}{\mathbf{u}}_{i,j}+\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{2q_{i}}{\mathbf{v}}_{i,j}+{\mathbf{u}}+(2\ell+1){\mathbf{v}},\quad\ell=1,\ldots,n-1$
are all the minimal vectors of $\mathrm{relint}({\mathbb{R}}_{+}(B^{0}))\cap\mathcal{S}(B^{0})$.
Similarly we may write
${\mathbb{R}}^{V(B^{s})}=\{\sum\limits_{i=1}^{m}\sum\limits_{j=1}^{2p_{i}}a_{i,j}{\mathbf{u}}_{i,j}+\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{2q_{i}}b_{i,j}{\mathbf{v}}_{i,j}+\sum\limits_{i=1}^{s-1}c_{i}{\mathbf{w}}_{i}+a{\mathbf{u}}+b{\mathbf{v}}\mid\mbox{ all }a_{i,j},b_{i,j},c_{i},a,b\in{\mathbb{R}}\}.$
and showed in [21, Section 6] that the following vectors
$\alpha_{\ell}:=\sum\limits_{i=1}^{m}\sum\limits_{j=1}^{2p_{i}}{\mathbf{u}}_{i,j}+\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{2q_{i}}{\mathbf{v}}_{i,j}+\sum\limits_{i=1}^{s-1}{\mathbf{w}}_{i}+{\mathbf{v}}+2\ell{\mathbf{u}},\ell=1,\ldots,m$
and
$\beta_{\ell}:=\sum\limits_{i=1}^{m}\sum\limits_{j=1}^{2p_{i}}{\mathbf{u}}_{i,j}+\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{2q_{i}}{\mathbf{v}}_{i,j}+\sum\limits_{i=1}^{s-1}{\mathbf{w}}_{i}+{\mathbf{u}}+2\ell{\mathbf{v}},\ell=1,\ldots,n$
are all the minimal vectors of $\mathrm{relint}({\mathbb{R}}_{+}(B^{s}))\cap\mathcal{S}(B^{s})$.
Lemma 5.3.
For each compact graph $H$ of type 2,
  1. $(1)$
    If $H=\mathbf{B}_{\underline{p}:\underline{q}}^{0}$, then
    $\mathcal{N}_{H}=\{\Theta_{H}u^{2(m-\ell)-1}v^{2n-1}\mid\ell=0,\ldots,m-1\}\sqcup\{\Theta_{H}u^{2m-1}v^{2(n-\ell)-1}\mid\ell=1,\ldots,n-1\}.$
  2. $(2)$
    If $H=\mathbf{B}_{\underline{p}:\underline{q}}^{s}$, then
    $\mathcal{N}_{H}=\{\Theta_{H}u^{2(m-\ell)+1}v^{2n}\mid\ell=1,\ldots,m\}\sqcup\{\Theta_{H}u^{2m}v^{2(n-\ell)+1}\mid\ell=1,\ldots,n\}.$
Proof.
If $H=\mathbf{B}_{\underline{p}:\underline{q}}^{0}$, then $D_{H}=\Theta_{H}^{2}u^{2m-1}v^{2n-1}$, and if $H=\mathbf{B}_{\underline{p}:\underline{q}}^{s}$, the expression for $D_{H}$ becomes $\Theta_{H}^{2}u^{2m}v^{2n}$. Given the minimal generators within the relative interior of the edge cones of $H$ as enumerated previously, the result is straightforwardly deduced.
Analogously in Lemma 5.3, the notation $u,v$ utilized in the expression of $\mathcal{N}_{H}$ represents the two big vertices of the graph $H$. Nevertheless, upon considering $H$ as an induced subgraph of a compact graph of type 3, there arises a possibility that the two big vertices of $H$ might alternately be denoted as $w,u$ or $w,v$. In such cases, it becomes necessary to substitute $w,u$ with either $v$ or $w,v$ accordingly.
5.3. type three
Given positive integers $m,n,k$, as well as the positive integral tuples $\underline{p}=(p_{1},\ldots,p_{m})$, $\underline{q}=(q_{1},\ldots,q_{n})$ and $\underline{r}=(r_{1},\ldots,r_{k})$, we denote the graph $\mathbf{C}_{\underline{p}:\underline{q}:\underline{r}}$ as $C$ for brevity. Here, $p_{i},q_{i},r_{i}$ are all positive integers. By definition, we may assume that $C$ has vertex set
$V(C)=$$\{\{u,v,w\}\sqcup\{u_{i,j}\mid 1\leq i\leq m,1\leq j\leq 2p_{i}\}$
$\sqcup\{v_{i,j}\mid 1\leq i\leq n,1\leq j\leq 2q_{i}\}\sqcup\{w_{i,j}\mid 1\leq i\leq k,1\leq j\leq 2r_{i}\},$
and edge set
$E(C)=$$\{\{u_{i,j},u_{i,j+1}\}\mid 1\leq i\leq m,1\leq j\leq 2p_{i}-1\}$
$\sqcup\{\{u,u_{i,1}\},\{u,u_{i,2p_{i}}\}\mid 1\leq i\leq m\}$
$\sqcup\{\{v_{i,j},v_{i,j+1}\}\mid 1\leq i\leq n,1\leq j\leq 2q_{i}-1\}$
$\sqcup\{\{v,v_{i,1}\},\{v,v_{i,2q_{i}}\}\mid 1\leq i\leq n\}$
$\sqcup\{\{w_{i,j},w_{i,j+1}\}\mid 1\leq i\leq k,1\leq j\leq 2r_{i}-1\}$
$\sqcup\{\{w,w_{i,1}\},\{w,w_{i,2r_{i}}\}\mid 1\leq i\leq k\}$
$\sqcup\{\{u,v\},\{v,w\},\{w,u\}\}.$
It follows from Conventions 2.3 that
${\mathbb{R}}^{V(C)}=\{\sum\limits_{i=1}^{m}\sum\limits_{j=1}^{2p_{i}}{a_{i,j}{\mathbf{u}}_{i,j}}+a{\mathbf{u}}+\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{2q_{i}}{b_{i,j}{\mathbf{v}}_{i,j}}+b{\mathbf{v}}+\sum\limits_{i=1}^{k}\sum\limits_{j=1}^{2r_{i}}{c_{i,j}{\mathbf{w}}_{i,j}}+c{\mathbf{w}}\mid$
$a_{i,j},a,b_{i,j},b,c_{i,j},c\in{\mathbb{R}}\mbox{ for all }i,j\}.$
According to [21, Section 6], we have the following result.
Lemma 5.4.
The minimal vectors of $\mathrm{relint}({\mathbb{R}}_{+}(C))\cap\mathcal{S}(C)$ are following $m+n+k$ vectors: $\alpha_{\ell}$ for $\ell=1,\ldots,m,$ $\beta_{\ell}$ for ${\ell}=1,\ldots,n$ and $\gamma_{\ell}$ for ${\ell}=1,\ldots,k$. Here, For $\ell=1,\ldots,m$,
$\alpha_{\ell}:=\sum\limits_{i=1}^{m}\sum\limits_{j=1}^{2p_{i}}{\mathbf{u}}_{i,j}+\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{2q_{i}}{\mathbf{v}}_{i,j}+\sum\limits_{i=1}^{k}\sum\limits_{j=1}^{2r_{i}}{\mathbf{w}}_{i,j}+{\mathbf{v}}+{\mathbf{w}}+2{\ell}{\mathbf{u}}.$
For ${\ell}=1,\ldots,n$,
$\beta_{\ell}=\sum\limits_{i=1}^{m}\sum\limits_{j=1}^{2p_{i}}{\mathbf{u}}_{i,j}+\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{2q_{i}}{\mathbf{v}}_{i,j}+\sum\limits_{i=1}^{k}\sum\limits_{j=1}^{2r_{i}}{\mathbf{w}}_{i,j}+{\mathbf{w}}+{\mathbf{u}}+2{\ell}{\mathbf{v}}.$
For ${\ell}=1,\ldots,k$,
$\gamma_{\ell}=\sum\limits_{i=1}^{m}\sum\limits_{j=1}^{2p_{i}}{\mathbf{u}}_{i,j}+\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{2q_{i}}{\mathbf{v}}_{i,j}+\sum\limits_{i=1}^{k}\sum\limits_{j=1}^{2r_{i}}{\mathbf{w}}_{i,j}+{\mathbf{u}}+{\mathbf{v}}+2{\ell}{\mathbf{w}}.$
In other words, the canonical module $\omega_{\mathbb{K}[C]}$ of $\mathbb{K}[C]$ is a monomial ideal of $\mathbb{K}[C]$ generated by
$\Theta_{C}u,\Theta_{C}u^{3},\ldots,\Theta_{C}u^{2m-1},\quad\Theta_{C}v,\Theta_{C}v^{3},\ldots,\Theta_{C}v^{2n-1}$
and
$\quad\Theta_{C}w,\Theta_{C}w^{3},\ldots,\Theta_{C}w^{2k-1}.$
The computation of $D_{C}$ is straightforward, resulting in $D_{C}=\Theta_{C}^{2}u^{2m}v^{2n}w^{2k}$. From this equality and applying Formula (1) in Subsection 2.5, we obtain the following result immediately.
Lemma 5.5.
If $H=\mathbf{C}_{\underline{p}:\underline{q}:\underline{r}}$, then
$\mathcal{N}_{H}=$$\{\Theta_{H}u^{2(m-\ell)+1}v^{2n}w^{2k}\mid\ell=1,\ldots,m\}\sqcup\{\Theta_{H}u^{2m}v^{2(n-\ell)+1}w^{2k}\mid\ell=1,\ldots,n\}$
$\sqcup\{\Theta_{H}u^{2m}v^{2n}w^{2(k-\ell)+1}\mid\ell=1,\ldots,k\}.$
So far, we have computed the top supports of all compact graphs. Considering the composition of these sets, we can conclude that, intriguingly, for each compact graph $H$, its top support $\mathcal{N}_{H}$ contains precisely $\mathfrak{t}(H)-1$ elements, where $\mathfrak{t}(H)$ denotes the number of minimal cycles in graph $H$. Moreover, we can conclude that for any compact graph $H$, $\Theta_{H}$ divides $\alpha$ for all $\alpha\in\mathcal{N}_{H}$. This provides another perspective to confirm that $H$ is a top-Betti graph.
6. Betti numbers of compact graphs
In this section, we employ the induced-subgraph approach to derive the multi-graded Betti numbers of the edge rings associated with compact graphs. Our objective is achieved by proving that the Betti numbers of the toric ideals of compact graphs coincide with those of a particular square-free initial ideal.
Lemma 6.1.
Let $G$ be a compact graph. Then
$\mathrm{(1)}$ The projective dimension of $\mathbb{K}[G]$ is equal to $\mathfrak{t}(H)-1$;
$\mathrm{(2)}$ There is a monomial order $<$ such that $\beta_{i}(\mathrm{in}_{<}(I_{G}))=(i+1)\binom{\mathfrak{t}(G)}{i+2}$ for all $i\geq 0$. Thus, for all $i\geq 1$, one has
$\beta_{i}(\mathbb{K}[G])\leq\beta_{i}(\mathbb{K}[E(G)]/\mathrm{in}_{<}(I_{G}))=i\binom{\mathfrak{t}(G)}{i+1}.$
Proof.
The conclusion (1) follows from [21, Corollary 5.2], and the conclusion (2) follows from [21, Theorem 5.1] as well as Lemma 2.1.
In the subsequent discussions, the monomial order denoted by “$<$” always refers to the order in Lemma 6.1, which was defined in [21].
Proposition 6.2.
Let $G$ be a compact graph of either type 1 or type 2. Then, for any $10\leq i\leq\mathfrak{t}(G)-$ 1, we have
$\beta_{i}(\mathbb{K}[G])=\beta_{i}(\mathbb{K}[E(G)]/\mathrm{in}_{<}(I_{G})).$
Proof.
If $i=0$, then $\beta_{i}(\mathbb{K}[G])=\beta_{i}(\mathbb{K}[E(G)]/\mathrm{in}_{<}(I_{G}))=1.$ Fix $11\leq i\leq\mathfrak{t}(G)-$ 1. Since there are $\mathfrak{t}(G)$ minimal cycles, there are $\binom{\mathfrak{t}(G)}{i+1}$ compact induced subgraphs of $G$ that contain exactly $i+1$ minimal cycles of $G$. These induced subgraphs are of either type one or type two, and are denoted by $H_{1},\ldots,H_{r}$, respectively. Here, $r=\binom{\mathfrak{t}(G)}{i+1}$. Note that $\beta_{i,\alpha}(\mathbb{K}[H_{\ell}])=1$ for any $\alpha\in\mathcal{N}_{H_{\ell}}$ by Lemmas 5.2 and 5.3, it follows that $\beta_{i,\alpha}(\mathbb{K}[G])=1$ for any $\alpha\in\bigsqcup_{\ell=1}^{r}\mathcal{N}_{H_{\ell}}$ in view of Lemma 2.8. Put $\mathcal{N}_{i}(G)=\bigsqcup_{\ell=1}^{r}\mathcal{N}_{H_{\ell}}$.
Note that $\mathrm{supp}(\alpha)=V(H_{k})$ for any $\alpha\in\mathcal{N}_{H_{k}}$, we have $\mathcal{N}_{H_{1}},\ldots,\mathcal{N}_{H_{r}}$ are pairwise disjoint. Hence,
$|\mathcal{N}_{i}(G)|=\sum_{\ell=1}^{r}|\mathcal{N}_{H_{\ell}}|=ir=i\binom{\mathfrak{t}(G)}{i+1}.$
This implies
$i\binom{\mathfrak{t}(G)}{i+1}=\sum_{\alpha\in\mathcal{N}_{i}(G)}\beta_{i,\alpha}(\mathbb{K}[G])\leq\beta_{i}(\mathbb{K}[G])\leq\beta_{i}(\mathbb{K}[E(G)]/\mathrm{in}_{<}(I_{G}))=i\binom{\mathfrak{t}(G)}{i+1},$
completing the proof.
From the proof of Proposition 6.2, it is evident that
$\beta_{i}(\mathbb{K}[G])=\sum_{\alpha\in\mathcal{N}_{i}(G)}\beta_{i,\alpha}(\mathbb{K}[G]).$
In particular, if $\beta_{i,\alpha}(\mathbb{K}[G])\neq 0$, then $\alpha\in\mathcal{N}_{i}(G)$. Therefore, we can derive the multi-graded and graded Betti numbers of $\mathbb{K}[G]$ as follows.
Corollary 6.3.
Let $G$ be a compact graph of either type 1 or type 2. For any monomial $\alpha$ in $\mathbb{K}[G]$, we have
$\beta_{i,\alpha}(\mathbb{K}[G])=\left\{\begin{array}[]{ll}1,&\hbox{$\alpha\in\mathcal{N}_{i}(G)$;}\\ 0,&\hbox{otherwise.}\end{array}\right.$
In particular, $\beta_{i,j}(\mathbb{K}[G])$ is equal to the number of the set $\{\alpha\in\mathcal{N}_{i}(G)\mid\mathrm{deg}(\alpha)=2j\}.$
Corollary 6.3 implies that if $G$ is a compact graph of type 1 or type 2 then $G$ is both a top-Betti graph and a second-Betti graph, but not a strict second-Betti graph.
We now turn to compact graphs of type 3. In the case when $G$ is of type one or type two, there are exactly $\binom{\mathfrak{t}(G)}{i+1}$ distinct compact induced subgraphs of $G$ which contains exactly $i+1$ minimal cycles for all $11\leq i\leq\mathfrak{t}(G)-$ 1. This is the reason why the proof of Theorem 6.2 holds. But this is not the case when considering compact graphs of type 3. Let us see the following example.
Example 6.4.
Assume that $G=\mathbf{C}_{(1):(1):(1,1)}$, as shown in Figure 2. Then $G$ contains 5 induced cycles. But there are only 4 distinct compact induced subgraphs of $G$ which contains 4 induced cycles. This is because the 4 cycles around the triangle formed three big vertices $\{u,v,w\}$ does not form an induced subgraph of $G$. This shows that $G$ is a strict second-Betti graph.
$x_{1}$$x_{2}$$w_{1}$$w_{2}$$u_{2}$$u_{1}$$v_{1}$$v_{2}$$u$$v$$w$
Figure 2. The graph $\mathbf{C}_{(1):(1):(1,1)}$
Since type 3 compact graphs are strict second-Betti graphs, we need to compute their second top multi-graded Betti numbers, leveraging Lemma 2.2. To apply Lemma 2.2, we make an observation related to the computation of simplicial complexes $\Gamma_{h}(I)$. Notably, we adhere to the notation established in Lemma 2.2 throughout our analysis. Let $\Omega$ be a set comprised of a number of monomials from $T$, and assume that the ideal $I$ is generated by all the monomials in $\Omega$. For each $\alpha\in\Omega$, we introduce a simplicial complex $\Gamma_{h}^{\alpha}(I)$ as follows. If $\alpha$ divides $Y^{h}$ and $Y^{h}/\alpha=y_{1}^{a_{1}}y_{2}^{a_{2}}\cdots y_{n}^{a_{n}}$ for some $a_{i}\geq 0$, then $\Gamma_{h}^{\alpha}(I)$ is specified as the complex: $F$ is a facet of $\Gamma_{h}^{\alpha}(I)$Longleftrightarrow there are $a_{i}\geq 0$ for $i\in[n]$ such that $Y^{h}=\alpha y_{1}^{a_{1}}y_{2}^{a_{2}}\cdots y_{n}^{a_{n}}$ and $F=\{i\in[n]\mid a_{i}\neq 0\}.$ Otherwise, $\Gamma_{h}^{\alpha}(I)$ is simply an empty set. We can easily check that
$\Gamma_{h}(I)=\bigcup_{\alpha\in\Omega}\Gamma_{h}^{\alpha}(I).$
Example 6.5.
Let $\mathbb{K}[G]$ be the edge ring of $G=\mathbf{C}_{(1):(1):(1,1)}$. Note that all the vertices of $G$ are labeled, as shown in Figure 2. Thus, $\Theta_{G}=uvwx_{1}x_{2}w_{1}w_{2}u_{1}u_{2}v_{1}v_{2}$. In the following, we denote $\Theta_{G}$ by $\Theta$ for short. Moreover, in view of Lemma 5.4, the canonical module $\omega_{\mathbb{K}[G]}$, as an ideal of $\mathbb{K}[G]$, is generated by monomials
$\Theta w,\Theta w^{3},\Theta u,\Theta v.$
The four monomials are denoted by $\beta_{i}$ for $i=1,\cdots,4$. Put $\alpha_{1}=\Theta uvw$ and $\alpha_{2}=\Theta uvw^{3}.$ Denote $uv,vw,wu$ by $y_{1},y_{2},y_{3}$ respectively.
Since $\Theta uvw=\beta_{1}y_{1}$, it follows that $\Gamma_{\alpha_{1}}^{\beta_{1}}(\omega_{\mathbb{K}[G]})=\langle\{1\}\rangle$;
Since $\beta_{2}$ does not divide $\Theta uvw$, it follows that $\Gamma_{\alpha_{1}}^{\beta_{2}}(\omega_{\mathbb{K}[G]})=\emptyset$;
Since $\Theta uvw=\beta_{3}y_{2}$, it follows that $\Gamma_{\alpha_{1}}^{\beta_{3}}(\omega_{\mathbb{K}[G]})=\langle\{2\}\rangle$;
Since $\Theta uvw=\beta_{4}y_{3}$, it follows that $\Gamma_{\alpha_{1}}^{\beta_{4}}(\omega_{\mathbb{K}[G]})=\langle\{3\}\rangle$.
Thus, $\Gamma_{\alpha_{1}}(\omega_{\mathbb{K}[G]})=\langle\{1\},\{2\},\{3\}\rangle$, and so $\dim_{\mathbb{K}}\widetilde{H}_{0}(\Gamma_{\alpha_{1}}(\omega_{\mathbb{K}[G]}),\mathbb{K})=2$.
Similarly, we have $\Gamma_{\alpha_{2}}^{\beta_{1}}(\omega_{\mathbb{K}[G]})=\langle\{2,3\}\rangle$, $\Gamma_{\alpha_{2}}^{\beta_{2}}(\omega_{\mathbb{K}[G]})=\langle\{1\}\rangle$ and $\Gamma_{\alpha_{2}}^{\beta_{3}}(\omega_{\mathbb{K}[G]})=\Gamma_{\alpha_{2}}^{\beta_{4}}=\emptyset$. Hence, $\Gamma_{\alpha_{2}}(\omega_{\mathbb{K}[G]})=\langle\{2,3\},\{1\}\rangle$ and it follows $\dim_{\mathbb{K}}\widetilde{H}_{0}(\Gamma_{\alpha_{2}}(\omega_{\mathbb{K}[G]}),\mathbb{K})=1$.
Note that $D_{G}=\Theta^{2}u^{2}v^{2}w^{4}$. Since $D_{G}/\alpha_{1}=\Theta uvw^{3}$ and $D_{G}/\alpha_{2}=\Theta uvw$, it follows from Lemma 2.2 as well as Lemma 2.7 that
$\beta_{3,\Theta uvw^{3}}(\mathbb{K}[G])=\beta_{1,\alpha_{1}}(\omega_{\mathbb{K}[G]})=2$
and
$\beta_{3,\Theta uvw}(\mathbb{K}[G])=\beta_{1,\alpha_{2}}(\omega_{\mathbb{K}[G]})=1.$
There are precisely four induced subgraphs of $G$, that consist of four minimal cycles of $G$. These four induced subgraphs are induced on $V(G)\setminus\{x_{1},x_{2}\}$ and $V(G)\setminus\{w_{1},w_{2}\}$, $V(G)\setminus\{u_{1},u_{2}\}$ and $V(G)\setminus\{v_{1},v_{2}\}$, denoted by $G_{1}$ and $G_{2}$, $G_{3}$ and $G_{4}$, respectively. According to Lemmas 5.4, we have
$\mathcal{N}_{G_{1}}=\Theta_{G_{1}}\{uv^{2}w^{2},u^{2}vw^{2},u^{2}v^{2}w\},\quad\mathcal{N}_{G_{2}}=\Theta_{G_{2}}\{uv^{2}w^{2},u^{2}vw^{2},u^{2}v^{2}w\}$
and
$\mathcal{N}_{G_{3}}=\Theta_{G_{3}}\{v^{2}w,v^{2}w^{3},vw^{4}\},\quad\mathcal{N}_{G_{4}}=\Theta_{G_{4}}\{u^{2}w,u^{2}w^{3},uw^{4}\}.$
Here, for a set $A$ of monomials and for a monomial $\alpha$, the notation $\alpha A$ denotes the set $\{\alpha a\mid a\in A\}$.
Combining the above facts, the sum of Betti numbers $\beta_{3,\alpha}(\mathbb{K}[G])$ with $\mbox{supp}(\alpha)=V(G)$ is greater than or equal to 3, and the sum of Betti numbers $\beta_{3,\alpha}(\mathbb{K}[G])$ with $\mbox{supp}(\alpha)=V(G_{i})$ is also greater than or equal to $\geq$ 3 for $i=1,2,3,4$. Therefore, it follows that
$1515\leq\beta_{3}(\mathbb{K}[G])\leq\beta_{3}(\mathbb{K}[E(G)]/\mathrm{in}_{<}(I_{G}))=(2+1)\binom{5}{4}=$ 15 .
Hence, $\beta_{3}(\mathbb{K}[G])=\beta_{3}(\mathbb{K}[E(G)]/\mathrm{in}_{<}(I_{G}))$, and $\beta_{3,\alpha}(\mathbb{K}[G])$ for $\alpha\in\mathbb{Z}^{V(G)}$ could also be written down.
The approach outlined in Example 6.5 can be applied to the general case.
Definition 6.6.
Let $H$ be the graph $\mathbf{C}_{\underline{p}:\underline{q}:\underline{r}}$ with its vertices of $H$ labeled as in Subsection 5.3. We define $\mathcal{M}_{H}^{1}$, $\mathcal{M}_{H}^{2}$ and $\mathcal{M}_{H}$ to be the following sets:
$\mathcal{M}^{1}_{H}:=$$\{\Theta_{H}u^{2\ell-1}v^{2n-1}w^{2k-1}\mid\ell=1,\ldots,m-1\}\sqcup\{\Theta_{H}u^{2m-1}v^{2\ell-1}w^{2k-1}\mid\ell=1,\ldots,n-1\}$
$\sqcup\{\Theta_{H}u^{2m-1}v^{2n-1}w^{2\ell-1}\mid\ell=1,\ldots,k-1\},$
$\mathcal{M}_{H}^{2}:=$$\{\Theta_{H}u^{2m-1}v^{2n-1}w^{2k-1}\}\mbox{ and }\mathcal{M}_{H}:=\mathcal{M}_{H}^{1}\sqcup\mathcal{M}_{H}^{2}.$
Lemma 6.7.
Let $H=\mathbf{C}_{\underline{p}:\underline{q}:\underline{r}}$ denote a compact graph of type 3 and let $p=m+n+k$ represent the projective dimension of $\mathbb{K}[H]$. Then,
$\beta_{p-1,\alpha}(\mathbb{K}[H])=\left\{\begin{array}[]{ll}1,&\hbox{$\alpha\in\mathcal{M}^{1}_{H}$;}\\ 2,&\hbox{$\alpha\in\mathcal{M}^{2}_{H}$.}\end{array}\right.$
In particular, we have $\sum\limits_{\alpha}\beta_{p-1,\alpha}(\mathbb{K}[H])\geq p-1$, where the summation is over all $\alpha\in\mathrm{Mon}(\mathbb{K}[H])$ with $\mathrm{supp}(\alpha)=V(H)$.
Proof.
Denote $uv,vw$ and $wu$ by $y_{1},y_{2}$ and $y_{3}$ respectively. Let $\omega$ denote the canonical module $\omega_{\mathbb{K}[H]}$. Note that $\omega$ is a monomial ideal of $\mathbb{K}[H]$, whose generators are given in Lemma 5.4. Put $\alpha=\Theta uvw$, where $\Theta$ is used as a shorthand for $\Theta_{H}$. Then
$\Gamma_{\alpha}^{\Theta u}(\omega)=\langle\{2\}\rangle,\Gamma_{\alpha}^{\Theta v}(\omega)=\langle\{3\}\rangle,\Gamma_{\alpha}^{\Theta w}(\omega)=\langle\{1\}\rangle.$
For other generators $\beta$ of $\omega$, one easily sees that $\Gamma_{\alpha}^{\beta}(\omega)=\emptyset$. This implies $\Gamma_{\alpha}(\omega)=\langle\{2\},\{3\},\{1\}\rangle$ and so $\beta_{1,\alpha}(\omega)=2$ by Lemma 2.2. In the same vein, we can prove
$\beta_{1,\Theta u^{2\ell-1}vw}(\omega)=1\mbox{ for }\ell=2,3,\ldots,m,$
$\beta_{1,\Theta uv^{2\ell-1}w}(\omega)=1\mbox{ for }\ell=2,3,\ldots,n,$
and
$\beta_{1,\Theta uvw^{2\ell-1}}(\omega)=1\mbox{ for }\ell=2,3,\ldots,k.$
Note that $D_{H}=\Theta^{2}u^{2m}v^{2n}w^{2k},$ the first assertion follows from Lemma 2.7. For the second assertion, it follows from the facts that $|\mathcal{M}^{1}_{H}|=m+n+k-3$, $|\mathcal{M}^{2}_{H}|=1$, and $\mathrm{supp}(\alpha)=V(H)$ for each $\alpha\in\mathcal{M}^{1}_{H}\sqcup\mathcal{M}^{2}_{H}$.
Proposition 6.8.
Let $G=\mathbf{C}_{\underline{p}:\underline{q}:\underline{r}}$ denote a compact graph of type 3. Then
$\beta_{i}(\mathbb{K}[G])=\beta_{i}(\mathbb{K}[E(G)]/\mathrm{in}_{<}(I_{G}))$
for all $i\geq 0$.
Proof.
We categorize the minimal cycles of $G$ into four distinct classes. The cycle encompassing the three big vertices is designated as the big cycle. A minimal cycle that is distinct from the big cycle but includes the vertex $u$ (or $v$, or $w$ again) is referred to as a cycle belonging to $u$ (or $v$, or $w$ respectively).
Fix $i\geq 1$. Then the number of subgraphs of $G$ that contain exactly $i+1$ minimal cycles is given by $\binom{m+n+k+1}{i+1}$. But not all of these subgraphs are induced subgraphs. Consider a subgraph $H$ of $G$ and let $C_{1},\ldots,C_{i+1}$ represent all its minimal cycles. Then one easily sees that $H$ is not an induced subgraph if and only if
  • None of $C_{j}$ for $j=1,\ldots,i+1$ is the big cycle;
  • For each $x\in\{u,v,w\}$, there is at least one index $j\in[i+1]$ such that $C_{j}$ is a minimal cycle belonging to $x$.
Hence, there are $\sum\limits_{a+b+c=i+1}^{a,b,c\geq 1}\binom{m}{a}\binom{n}{b}\binom{k}{c}$ such subgraphs and it follows that
$\sum_{\alpha}\beta_{i,\alpha}(\mathbb{K}[G])\geq\left(\binom{m+n+k+1}{i+1}-\sum\limits_{a+b+c=i+1}^{a,b,c\geq 1}\binom{m}{a}\binom{n}{b}\binom{k}{c}\right)i,$
where the summation is over all $\alpha\in\mathrm{Mon}(\mathbb{K}[G])$ such that the induced subgraph on $\mathrm{supp}(\alpha)$ consisting of $i+1$ minimal cycles. On the other hand, we notice the following fact: if $H$ is an induced subgraph of $G$ containing exactly $i+2$ minimal cycles, say $C_{1},\ldots,C_{i+2}$, then $H$ belongs to type 3 if and only if
  • One of $C_{j}$ for $j=1,\ldots,i+2$ is the big cycle. For simplicity, let us assume that $C_{i+2}$ is the big cycle;
  • For each $x\in\{u,v,w\}$, there is at least one index $j\in[i+1]$ such that $C_{j}$ is a minimal cycle belonging to $x$.
Thus, there are $\sum\limits_{a+b+c=i+1}^{a,b,c\geq 1}\binom{m}{a}\binom{n}{b}\binom{k}{c}$ compact induced subgraphs of $G$ that belong to type 3 and contain exactly $i+2$ minimal cycles. It follows from Lemma 6.7 that
$\sum_{\alpha}\beta_{i,\alpha}(\mathbb{K}[G])\geq\sum\limits_{a+b+c=i+1}^{a,b,c\geq 1}\binom{m}{a}\binom{n}{b}\binom{k}{c}i,$
where the summation is over all $\alpha\in\mathrm{Mon}(\mathbb{K}[G])$ such that the induced subgraph on $\mathrm{supp}(\alpha)$ is a compact graph of type 3 consisting of $i+2$ minimal cycles. Hence, we have
$\beta_{i}(\mathbb{K}[G])\geq\binom{m+n+k+1}{i+1}i.$
Now, the result follows from Lemma 6.1.
From the proof of Proposition 6.8, we see that the inequality in Lemma 6.7 is actually an equality. By combining all the findings we have made so far, we can immediately obtain the main result of this section.
Theorem 6.9.
Let $G$ be a compact graph, and denote by $p=\mathfrak{t}(G)-1$ the projective dimension of $\mathbb{K}[G]$. Consider an integer $i$ satisfying $1\leq i\leq p$. Let $H_{j}$ for $j=1,\ldots,s$ represent all the compact induced subgraphs of $G$ having exactly $i+1$ minimal cycles, and let $L_{j}$ for $j=1,\ldots,t$ be all the compact induced subgraphs of $G$ belonging to type 3 with $i+2$ minimal cycles. Define the following sets:
$\mathcal{N}_{i}(G)=\bigsqcup_{j=1}^{s}\mathcal{N}_{H_{j}},\quad\mathcal{M}_{i}^{1}(G)=\bigsqcup_{j=1}^{t}\mathcal{M}_{L_{j}}^{1},\quad\text{and}\quad\mathcal{M}_{i}^{2}(G)=\bigsqcup_{j=1}^{t}M_{L_{j}}^{2}.$
Then, the $i$-th multi-graded Betti numbers of $\mathbb{K}[G]$ can be expressed as follows:
$\beta_{i,\alpha}(\mathbb{K}[G])=\begin{cases}2,&\text{if }\alpha\in\mathcal{M}_{i}^{2}(G),\\ 1,&\text{if }\alpha\in\mathcal{M}_{i}^{1}(G)\sqcup\mathcal{N}_{i}(G),\\ 0,&\text{otherwise}.\end{cases}$
Note that $t$ is always equal to 0 if $G$ is a compact graph of type 1 or 2.
For all possible integers $i,j$, we set $O_{i,j}=\{\alpha\in\mathcal{M}_{i+1}^{1}(G)\sqcup\mathcal{N}_{i+1}(G)\mid\mathrm{deg}(\alpha)=2j\}$ and $T_{i,j}=\{\alpha\in\mathcal{M}_{i+1}^{2}(G)\mid\mathrm{deg}(\alpha)=2j\}$.
Corollary 6.10.
Let $G$ be a compact graph, and let $I_{G}$ denote its toric ideal. There exists a square-free initial ideal $\mathrm{in}_{<}(I_{G})$ of $I_{G}$ such that $\beta_{i,j}(\mathrm{in}_{<}(I_{G}))=\beta_{i,j}(I_{G})=|O_{i.j}|+2|T_{i,j}|$.
Acknowledgment: This project is supported by NSFC (No. 11971338). The authors are grateful to the software systems [4] for providing us with a large number of examples that help us develop ideas and test the results.